### v1

# Hierarchical filtrations of torsion-free sheaves and birational geometry

Abstract.
We introduce the notion of hierarchical filtrations of torsion-free sheaves on normal projective varieties and define the associated numerical invariant called hierarchical depth. This invariant measures the maximal length of filtrations by saturated subsheaves of equal rank whose successive quotients are torsion sheaves supported in codimension one.
We establish general bounds for hierarchical depth in terms of the divisor class of the determinant and give exact formulas in several basic geometric situations, including the case of smooth projective curves and varieties of Picard rank one. A key technical ingredient is the study of elementary transforms along effective divisors and their commutativity properties.
In dimension two, we analyze the behavior of hierarchical depth under birational morphisms and show that it admits a precise description along the minimal model program. In particular, we prove that hierarchical depth transforms additively with respect to exceptional divisors and is explicitly computable on minimal models.
As an application, we relate hierarchical depth to degeneracies in algebraic–geometric codes and show that birational simplification via the minimal model program leads to effective improvements of code parameters. These results demonstrate that hierarchical depth provides a new bridge between the birational geometry of vector bundles and arithmetic applications.
Key words and phrases:
Vector bundles, Hierarchical filtration, MMP, AG codes
2020 Mathematics Subject Classification:
primary; 14G50, Secondary; 14C20, 14H52
Introduction
Filtrations of vector bundles and coherent sheaves are among the most classical tools in algebraic geometry. Since the foundational work of Grothendieck on coherent sheaves and their cohomology [4], filtrations have provided a systematic way to decompose global objects into simpler pieces reflecting both geometric and numerical properties. Standard references such as [5, 3, 20, 21] illustrate the central role of torsion-free and reflexive sheaves in the study of divisors, intersection theory, and birational geometry.
One of the most influential filtration-theoretic constructions is the Harder–Narasimhan filtration, originally introduced for vector bundles on curves [7] and later extended to higher-dimensional varieties. This filtration underlies much of modern stability theory and plays a fundamental role in the geometry of moduli spaces of sheaves, as developed for instance in [12, 13, 8]. Closely related are Jordan–Hölder filtrations and refinements appearing in the theory of semistable sheaves. These constructions, however, depend essentially on the choice of a polarization and are therefore sensitive to auxiliary data.
In parallel, birational geometry has undergone a profound development through the minimal model program. Initiated by Mori and Shokurov [17, 18, 19], and culminating in the existence of minimal models for varieties of log general type [1], the MMP provides a structural framework for understanding higher-dimensional algebraic varieties via divisorial contractions, flips, and minimal models. Modern expositions may be found in [2, 10]. Despite this progress, comparatively little is known about how intrinsic invariants of vector bundles and torsion-free sheaves behave along the steps of the minimal model program.
The purpose of this paper is to introduce a new filtration-theoretic framework for torsion-free sheaves that is intrinsically adapted to divisorial and birational geometry. We define the notion of a hierarchical filtration, consisting of successive saturated subsheaves of equal rank whose quotients are torsion sheaves supported in codimension one. Associated to such filtrations is a discrete numerical invariant, called the hierarchical depth, measuring the maximal possible length of such a filtration. Hierarchical filtrations were first introduced and studied for line bundles in [14], where they were applied to the construction of optimal algebraic–geometric codes. The present paper develops a substantially broader theory for torsion-free sheaves of arbitrary rank and investigates its interaction with birational geometry and the minimal model program. Unlike stability-theoretic invariants, hierarchical depth is independent of any choice of polarization and is governed instead by the determinant of the sheaf and the divisor theory of the underlying variety.
The first part of the paper develops the basic properties of hierarchical filtrations. Using standard techniques from the theory of divisors and intersection theory [3, 9], we prove finiteness of hierarchical depth and derive effective bounds in terms of the divisor class of the determinant. In several basic geometric situations, including smooth projective curves and varieties of Picard rank one, we obtain explicit formulas. These results show that hierarchical depth is a computable invariant determined by classical numerical data.
A key technical aspect of the theory is the interpretation of hierarchical filtrations via iterated elementary transforms along effective divisors. Elementary transforms appear naturally in the study of vector bundles and reflexive sheaves, particularly in connection with degenerations and modifications along divisors; see for instance [6, 13]. We show that, under suitable conditions, such transforms commute and may be reordered without affecting the resulting hierarchical depth. This yields a concrete description of the invariant and clarifies its independence from auxiliary choices.
The second part of the paper is devoted to the interaction between hierarchical depth and birational geometry. After establishing general functorial properties under pullback and pushforward along birational morphisms, we specialize to the case of surfaces. In dimension two, we prove that hierarchical depth transforms additively along the steps of the minimal model program. In particular, the invariant admits an explicit description on minimal models, with its variation controlled by exceptional divisors and their numerical contributions. This situates hierarchical depth as a genuinely birational invariant in the surface case, in the spirit of the MMP as developed in [1, 2, 10].
Finally, we present applications to algebraic–geometric codes constructed from vector bundles. Vector bundle methods in coding theory have been studied by several authors; see for example [11, 16, 15]. Building on earlier work on hierarchical filtrations for line bundles [14], we show that hierarchical depth detects divisorial degeneracies affecting code parameters. By combining the filtration theory developed here with birational simplification via the minimal model program, we obtain effective improvements in code constructions.
The paper is organized as follows. Section 1 reviews preliminaries and fixes notation. Section 2 introduces hierarchical filtrations and hierarchical depth. Section 3 is devoted to elementary transforms and their commutativity properties. Determinantal bounds and finiteness results are established in Section 4. In Section 5 we analyze the behavior of hierarchical depth under birational morphisms and study its interaction with the surface minimal model program. Slope behavior under increments, birational stability under the MMP and comparisons with classical filtrations are given in Section 6. Applications to algebraic–geometric codes are discussed in Section 7.
Acknowledgments
The author is grateful to Professor Pierre Deligne for valuable comments that led to a clarification and correction of the definition of hierarchical filtration.
1. Preliminaries and conventions
Throughout this paper, we work over an algebraically closed field $k$ of arbitrary characteristic. All schemes are assumed to be separated and of finite type over $k$.
1.1. Varieties and divisors
A variety is an integral normal projective scheme over $k$. Unless explicitly stated otherwise, all varieties considered in this paper have dimension at most $2$.
For a variety $X$, we denote by $\mathrm{Pic}(X)$ its Picard group and by $\mathrm{Div}(X)$ the group of Weil divisors on $X$. A divisor $D\in\mathrm{Div}(X)$ is called effective, denoted $D\geq 0$, if it is a finite non-negative linear combination of irreducible prime divisors. Linear equivalence of divisors is denoted by $\sim$.
When $X$ is smooth, Weil and Cartier divisors are identified. In general, divisor classes will be considered in the divisor class group of $X$.
1.2. Coherent sheaves and saturation
All sheaves in this paper are coherent $\mathcal{O}_{X}$-modules. A coherent sheaf $\mathcal{E}$ on $X$ is called torsion-free if it has no nonzero subsheaves supported in codimension at least $1$.
Let $\mathcal{E}$ be a torsion-free sheaf on $X$. A subsheaf $\mathcal{F}\subseteq\mathcal{E}$ is said to be saturated if the quotient $\mathcal{E}/\mathcal{F}$ is torsion-free. Equivalently, $\mathcal{F}$ is saturated if it coincides with its image in the reflexive hull $\mathcal{E}^{\vee\vee}$.
Standing convention. Unless explicitly stated otherwise, all subsheaves of torsion-free sheaves are assumed to be saturated.
With this convention, if $\mathcal{F}\subseteq\mathcal{E}$ are torsion-free sheaves of the same rank, then the quotient $\mathcal{E}/\mathcal{F}$ is a torsion sheaf supported in codimension $1$.
1.3. Determinants
Let $\mathcal{E}$ be a torsion-free sheaf of rank $r$ on a normal variety $X$. The determinant of $\mathcal{E}$ is defined by
$\det(\mathcal{E}):=\bigl(\wedge^{r}\mathcal{E}\bigr)^{\vee\vee},$
which is a rank-one reflexive sheaf. We identify $\det(\mathcal{E})$ with the corresponding divisor class on $X$.
If
$0\longrightarrow\mathcal{F}\longrightarrow\mathcal{E}\longrightarrow\mathcal{Q}\longrightarrow 0$
is an exact sequence of coherent sheaves on $X$ with $\mathcal{F}$ and $\mathcal{E}$ torsion-free of the same rank, then $\mathcal{Q}$ is a torsion sheaf supported in codimension $1$, and there exists an effective divisor $D$ on $X$ such that
$\det(\mathcal{F})\simeq\det(\mathcal{E})(-D).$
We refer to $D$ as the determinantal divisor associated with the inclusion $\mathcal{F}\subseteq\mathcal{E}$.
1.4. Elementary transforms
Let $\mathcal{E}$ be a torsion-free sheaf on $X$, and let $D$ be an effective divisor on $X$. An elementary transform of $\mathcal{E}$ along $D$ is a torsion-free sheaf $\mathcal{E}^{\prime}$ fitting into an exact sequence
$0\longrightarrow\mathcal{E}^{\prime}\longrightarrow\mathcal{E}\longrightarrow\mathcal{T}\longrightarrow 0,$
where $\mathcal{T}$ is a torsion sheaf supported on $D$.
Elementary transforms preserve the rank of the sheaf and modify the determinant by an effective divisor supported on $D$. They will be used extensively in the construction of hierarchical filtrations.
2. Hierarchical filtration of vector bundles
Definition 2.1.
Let $X$ be a smooth projective variety over a field. Fix a reference line bundle
$\Lambda_{0}\in\mathrm{Pic}(X)$
(which most commonly one takes to be $\mathcal{O}_{X}$). Let $\mathcal{E}$ be a vector bundle of rank $r$ on $X$.
A hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$ is a finite chain of inclusions of locally free $\mathcal{O}_{X}$-subsheaves
$\mathfrak{F}:\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E},$
satisfying the following conditions for every $i=1,\dots,h$:
  1. (1)
    each $\mathcal{E}_{i}$ is a vector bundle of rank $r$, and the inclusion $\mathcal{E}_{i-1}\hookrightarrow\mathcal{E}_{i}$ is saturated, i.e. the quotient
    $\mathcal{Q}_{i}:=\mathcal{E}_{i}/\mathcal{E}_{i-1}$
    is a nonzero torsion sheaf (equivalently $\mathcal{E}_{i-1}$ is a saturated subsheaf of $\mathcal{E}_{i}$).
  2. (2)
    there exists a nonzero effective Cartier divisor $D_{i}\subset X$ such that the following equivalent formulation holds:
    $\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i}).$
  3. (3)
    the filtration is normalized by the requirement
    $\det(\mathcal{E}_{0})\simeq\Lambda_{0}.$
Note that each one of two following statements are equivalent to condition (2) of above definition:
(2-1) the torsion sheaf $\mathcal{Q}_{i}$ is supported (set-theoretically) on $D_{i}$, and
$c_{1}(\mathcal{E}_{i})=c_{1}(\mathcal{E}_{i-1})+[D_{i}]\quad\in\mathrm{NS}(X).$
(2-2) locally on $X$ the inclusion $\mathcal{E}_{i-1}\hookrightarrow\mathcal{E}_{i}$ is obtained from $\mathcal{E}_{i-1}$ by an elementary transform (Hecke modification) along the Cartier divisor $D_{i}$; equivalently there is an exact sequence
$0\longrightarrow\mathcal{E}_{i-1}\longrightarrow\mathcal{E}_{i}\longrightarrow\mathcal{T}_{i}\longrightarrow 0$
with $\mathcal{T}_{i}$ a torsion sheaf supported on $D_{i}$ whose determinant (in the Knudsen-Mumford sense, or via the Fitting ideal) is $\mathcal{O}_{X}(D_{i})$.
Definition 2.2 (Hierarchical depth).
Let $\mathcal{E}$ and $\Lambda_{0}$ be as above. The hierarchical depth of $\mathcal{E}$ with respect to $\Lambda_{0}$, denoted $h_{\Lambda_{0}}(\mathcal{E})$, is the maximal integer $h\geq 0$ for which there exists a hierarchical filtration of length $h$ normalized by $\Lambda_{0}$. If no such finite filtration exists we set $h_{\Lambda_{0}}(\mathcal{E})=$-$\infty$. We adopt the convention $h_{\Lambda_{0}}(\mathcal{E}_{0})=0$ when $\mathcal{E}_{0}$ already equals the chosen starting bundle (equivalently when the chain is trivial).
Remark 2.3.
  1. (i)
    Equivalence of (2),(2-1),(2-2). Under the saturation hypothesis the three formulations in (2)(a)–(c) are equivalent: (c) implies (b) by taking first Chern classes and (b) implies (a) because the determinant of a short exact sequence with torsion quotient acquires the line bundle associated to the divisor supporting the torsion; conversely (a) together with saturation forces the quotient to be torsion supported on the divisor appearing in the determinant relation.
  2. (ii)
    Normalization. If one does not normalize the filtration (for instance by fixing $\det(\mathcal{E}_{0})$), the notion of hierarchical filtration (and the numerical data $(D_{i})$) is only defined up to simultaneous twist of every term by a line bundle. In particular one can “renormalize” a filtration by tensoring the entire chain with $\mathcal{O}_{X}(-D)$ and thereby create non-canonical shifts. Requiring $\det(\mathcal{E}_{0})\simeq\Lambda_{0}$ (for example $\Lambda_{0}=\mathcal{O}_{X}$) pins down this ambiguity and prevents the filtration from sliding to -$\infty$.
  3. (iii)
    Twist–invariance property. The hierarchical depth is stable under tensor twists, provided the normalization is adjusted accordingly: for any $M\in\mathrm{Pic}(X)$ one has
    (1) $h_{\Lambda_{0}\otimes M^{\otimes r}}(\mathcal{E}\otimes M)=h_{\Lambda_{0}}(\mathcal{E}),$
    where $r=\mathrm{rank}(\mathcal{E})$. Indeed, tensoring a normalized filtration of $\mathcal{E}$ by $M$ yields a normalized filtration of $\mathcal{E}\otimes M$ with respect to $\Lambda_{0}\otimes M^{\otimes r}$, since $\det(\mathcal{E}_{i}\otimes M)\simeq\det(\mathcal{E}_{i})\otimes M^{\otimes r}$ at every step. Hence the numerical invariant $h_{\Lambda_{0}}(\mathcal{E})$ depends only on the pair $(\det(\mathcal{E}),\Lambda_{0})$ modulo this canonical renormalization.
  4. (iv)
    Intrinsic interpretation. Equation (1) shows that the hierarchical depth is an intrinsic invariant of the bundle $\mathcal{E}$ up to tensor equivalence, analogous to how the slope $\mu(\mathcal{E})=\deg(\mathcal{E})/\mathrm{rank}(\mathcal{E})$ is invariant under twists by degree-zero line bundles on curves. The normalization by $\Lambda_{0}$ thus provides a canonical reference frame for measuring divisorial increments in $\mathrm{Pic}(X)$.
  5. (v)
    Rank one case. When $r=1$ this definition recovers Definition 1.1 of [14] (with the same normalization choice $\Lambda_{0}=\mathcal{O}_{X}$ if one so desires): then $\mathcal{E}_{i}$ are line bundles and the determinant condition is the same as $\mathcal{E}_{i}\cong\mathcal{E}_{i-1}\otimes\mathcal{O}_{X}(D_{i})$.
Proposition 2.4.
Let $X$ be a smooth projective variety of dimension $n$ such that $\mathrm{Pic}(X)\cong\mathbb{Z}\cdot A$ for some ample effective divisor $A$. Fix a reference line bundle $\Lambda_{0}\in\mathrm{Pic}(X)$ and an ample polarization $\mathcal{O}_{X}(1)$. Let $\mathcal{E}$ be a vector bundle on $X$, and write
$\det(\mathcal{E})\simeq\mathcal{O}_{X}(dA)$
for a unique integer $d\in\mathbb{Z}$. Write $\Lambda_{0}\simeq\mathcal{O}_{X}(d_{0}A)$.
Then any hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$ has length bounded by
$h\leq d-d_{0}.$
In particular, when $\Lambda_{0}\cong\mathcal{O}_{X}$ (so $d_{0}=0$) one has the uniform upper bound
$h_{\Lambda_{0}}(\mathcal{E})\leq d.$
Moreover, equality $h_{\Lambda_{0}}(\mathcal{E})=d-d_{0}$ occurs whenever there exists a hierarchical filtration of $\mathcal{E}$ whose successive determinant increments are all linearly equivalent to $A$, i.e. there exist effective divisors $D_{1},\dots,D_{d-d_{0}}\sim A$ and a chain
$\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{d-d_{0}}=\mathcal{E}$
(normalized by $\det(\mathcal{E}_{0})\simeq\Lambda_{0}$) with $\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i})$ for every $i$.
Proof.
Let
$\mathfrak{F}:\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
be a hierarchical filtration normalized by $\Lambda_{0}$, so $\det(\mathcal{E}_{0})\simeq\Lambda_{0}$. For each $i$ there is a nonzero effective Cartier divisor $D_{i}$ with
$\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i}).$
Thus
$\det(\mathcal{E})\simeq\det(\mathcal{E}_{0})\otimes\mathcal{O}_{X}\Big(\sum_{i=1}^{h}D_{i}\Big),$
and in $\mathrm{NS}(X)$
$c_{1}(\det\mathcal{E})-c_{1}(\Lambda_{0})=\big[D_{1}+\cdots+D_{h}\big].$
Under the hypothesis $\mathrm{Pic}(X)\cong\mathbb{Z}\cdot A$, each effective Cartier divisor $D_{i}$ is linearly equivalent to a positive integer multiple of $A$. Hence there exist integers $d_{i}\geq 1$ such that
$D_{i}\sim d_{i}A\qquad(i=1,\dots,h).$
Therefore in $\mathrm{Pic}(X)$ we have
$\mathcal{O}_{X}(dA)\simeq\det(\mathcal{E})\simeq\Lambda_{0}\otimes\mathcal{O}_{X}\Big(\sum_{i=1}^{h}D_{i}\Big)\simeq\mathcal{O}_{X}\Big(d_{0}A+\sum_{i=1}^{h}d_{i}A\Big),$
so equality of divisor classes yields the integer identity
$d=d_{0}+\sum_{i=1}^{h}d_{i}.$
Since each $d_{i}\geq 1$ we obtain
$d-d_{0}=\sum_{i=1}^{h}d_{i}\geq h,$
which is exactly the desired inequality $h\leq d-d_{0}$.
The claim about equality is immediate: equality $h=d-d_{0}$ holds precisely when every $d_{i}=1$, i.e. when each determinant increment $D_{i}$ is linearly equivalent to $A$. This completes the proof. ∎
Remark 2.5.
(1) The upper bound $h_{\Lambda_{0}}(\mathcal{E})\leq d-d_{0}$ follows directly from the fact that, under the hypothesis $\mathrm{Pic}(X)\cong\mathbb{Z}\cdot A$, every effective increment $D_{i}$ satisfies $D_{i}\sim d_{i}A$ with $d_{i}\geq 1$. Thus
$d-d_{0}=\sum_{i=1}^{h}d_{i}\geq h,$
showing that the maximal possible hierarchical depth depends only on the difference of determinant classes $\det(\mathcal{E})$ and $\Lambda_{0}$, but not on the internal geometry of $\mathcal{E}$. This is the exact analogue of the line–bundle case.
(2) Unlike the line–bundle case, equality $h_{\Lambda_{0}}(\mathcal{E})=d-d_{0}$ is generally not automatic for vector bundles. Equality would require that each increment divisor satisfy $D_{i}\sim A$, i.e. that $d_{i}=1$ for all $i$ in the decomposition
$D_{i}\sim d_{i}A,\qquad d_{i}\geq 1.$
However, for a given bundle $\mathcal{E}$ there may be no hierarchical filtration whose determinant increments are all linearly equivalent to $A$. Geometrically, this means that although $\mathrm{Pic}(X)$ has rank one, the structure of $\mathcal{E}$ (e.g. its possible elementary transforms, base–locus behavior, or torsion constraints) may force certain increments to have degree $d_{i}>1$, reducing the achievable depth below the theoretical maximum $d-d_{0}$.
3. Elementary transforms and commutativity
Definition 3.1 (Elementary transformation).
[8, Definition 5.2.1] Let $X$ be a smooth projective surface and $C$ be an effective divisor on $X$. Let $i\colon C\hookrightarrow X$ denote the embedding $C\subset X$. Let $\mathcal{F}$ be a vector bundle on $X$ and let $\mathcal{G}$ be a vector bundle on $C$. A coherent sheaf $\mathcal{E}$ on $X$ is called an elementary transform of $\mathcal{F}$ along $\mathcal{G}$ if there exists a short exact sequence
(2) $0\longrightarrow\mathcal{E}\longrightarrow\mathcal{F}\longrightarrow i_{*}\mathcal{G}\longrightarrow 0.$
It follows from Proposition 5.2.2 from [8] that $\mathcal{E}$ is locally free and, moreover, it is a vector bundle since $X$ is smooth.
Lemma 3.2.
Let $X$ be a smooth projective variety and let $\mathcal{E}$ be a vector bundle on $X$. Let $D_{1},D_{2}\subset X$ be effective Cartier divisors with
$\mathrm{Supp}(D_{1})\cap\mathrm{Supp}(D_{2})=\varnothing.$
For $i=1,2$ let $\mathcal{Q}_{i}$ be a nonzero coherent torsion sheaf supported on $D_{i}$ and let $\varphi_{i}:\mathcal{E}\to\mathcal{Q}_{i}$ be surjective $\mathcal{O}_{X}$–linear maps. Define
$V_{1}:=\ker(\varphi_{1}),\qquad V_{2}:=\ker(\varphi_{2}),$
and the successive kernels
$V_{12}:=\ker\big(V_{1}\xrightarrow{\ \overline{\varphi}_{2}\ }\mathcal{Q}_{2}\big),\qquad V_{21}:=\ker\big(V_{2}\xrightarrow{\ \overline{\varphi}_{1}\ }\mathcal{Q}_{1}\big),$
where $\overline{\varphi}_{j}$ denotes the map induced by $\varphi_{j}$ on the relevant kernel. Finally set
$V:=\ker\big(\mathcal{E}\xrightarrow{\ (\varphi_{1},\varphi_{2})\ }\mathcal{Q}_{1}\oplus\mathcal{Q}_{2}\big).$
Then the following hold.
  1. (1)
    As subsheaves of $\mathcal{E}$ we have
    $V_{12}=V_{21}=V.$
    In particular, the order of taking kernels (elementary transforms) with respect to $\varphi_{1}$ and $\varphi_{2}$ does not affect the resulting subsheaf of $\mathcal{E}$.
  2. (2)
    If, in addition, the three kernels $V_{1},V_{2},V$ are locally free , then the equalities in (1) are equalities of subbundles of $\mathcal{E}$. In this case we say that the corresponding elementary transforms commute and the order of the transforms is immaterial in the category of vector bundles.
Proof.
(1) Since $\mathrm{Supp}(\mathcal{Q}_{1})$ and $\mathrm{Supp}(\mathcal{Q}_{2})$ are disjoint, the two surjective maps $\varphi_{1},\varphi_{2}$ impose independent conditions on local sections of $\mathcal{E}$. Concretely, for any open set $U\subset X$ we have
$V(U)$ $=\{s\in\mathcal{E}(U):\varphi_{1}(s)=0\text{ in }\mathcal{Q}_{1}(U),\ \varphi_{2}(s)=0\text{ in }\mathcal{Q}_{2}(U)\},$
$V_{12}(U)$ $=\{s\in\mathcal{E}(U):\varphi_{1}(s)=0\text{ and }\varphi_{2}(s)=0\},$
$V_{21}(U)$ $=\{s\in\mathcal{E}(U):\varphi_{2}(s)=0\text{ and }\varphi_{1}(s)=0\}.$
The equalities of the three sets above are immediate; hence the three kernels coincide as subsheaves of $\mathcal{E}$.
(2) If the kernels $V_{1},V_{2},V$ are locally free, then the equalities of subsheaves are equalities of locally free subsheaves (subbundles). In particular all constructions may be carried out inside the category of vector bundles and the elementary transforms commute as vector bundles. ∎
Remark 3.3.
(1) The disjointness hypothesis in Lemma 3.2 is essential. If $\mathrm{Supp}(D_{1})\cap\mathrm{Supp}(D_{2})\neq\varnothing$ then the two quotient maps may interact on the overlap and the iterated kernels need not coincide (indeed saturation may collapse intermediate steps). (2) The lemma is stated for arbitrary torsion quotients $\mathcal{Q}_{i}$ supported on the Cartier divisors. The additional local-freeness hypothesis in (2) is exactly the bridge to the Huybrechts–Lehn notion of elementary transform ([8]) where one typically takes $\mathcal{Q}_{i}=i_{*}(\mathcal{G}_{i})$ for a vector bundle $\mathcal{G}_{i}$ on $D_{i}$; in that case the kernel is usually a vector bundle.
Corollary 3.4.
Let $X$ be a smooth projective variety and let $\mathcal{E}$ be a vector bundle on $X$. Suppose a hierarchical filtration of $\mathcal{E}$ is given by a finite sequence of elementary transforms
$\mathcal{E}=V_{0}\supset V_{1}\supset\cdots\supset V_{h},\qquad V_{k+1}=\ker\bigl(V_{k}\xrightarrow{\ \varphi_{k}\ }\mathcal{Q}_{k}\bigr),$
where for each $k$ the sheaf $\mathcal{Q}_{k}$ is a nonzero torsion sheaf supported on an effective Cartier divisor $D_{k}\subset X$.
Assume that the set of divisors $\{D_{k}\}$ is partitioned into two disjoint collections
$\mathcal{A}=\{D_{a_{1}},\dots,D_{a_{r}}\},\qquad\mathcal{B}=\{D_{b_{1}},\dots,D_{b_{s}}\},$
such that
$\mathrm{Supp}(D_{a})\cap\mathrm{Supp}(D_{b})=\varnothing\qquad\text{for all }D_{a}\in\mathcal{A},\ D_{b}\in\mathcal{B}.$
Then:
  1. (1)
    For every pair $D_{a}\in\mathcal{A}$ and $D_{b}\in\mathcal{B}$, the elementary transforms along $D_{a}$ and $D_{b}$ commute. More precisely, for any surjective morphisms $\varphi_{a}:\mathcal{E}\to\mathcal{Q}_{a}$ and $\varphi_{b}:\mathcal{E}\to\mathcal{Q}_{b}$ with $\mathrm{Supp}(\mathcal{Q}_{a})\subset D_{a}$ and $\mathrm{Supp}(\mathcal{Q}_{b})\subset D_{b}$, one has isomorphisms of subsheaves of $\mathcal{E}$:
    $\ker\Bigl(\ker(\mathcal{E}\xrightarrow{\varphi_{a}}\mathcal{Q}_{a})\xrightarrow{\overline{\varphi}_{b}}\mathcal{Q}_{b}\Bigr)=\ker\Bigl(\ker(\mathcal{E}\xrightarrow{\varphi_{b}}\mathcal{Q}_{b})\xrightarrow{\overline{\varphi}_{a}}\mathcal{Q}_{a}\Bigr)=\ker\Bigl(\mathcal{E}\xrightarrow{\ (\varphi_{a},\varphi_{b})\ }\mathcal{Q}_{a}\oplus\mathcal{Q}_{b}\Bigr),$
    where $\overline{\varphi}_{j}$ denotes the map induced on the kernel of $\varphi_{i}$.
  2. (2)
    By repeatedly interchanging adjacent transforms involving one divisor from $\mathcal{A}$ and one from $\mathcal{B}$, the filtration can be reordered into another hierarchical filtration of the same length $h$, in which all transforms supported on divisors in $\mathcal{A}$ occur before all transforms supported on divisors in $\mathcal{B}$.
  3. (3)
    In particular, increments supported on disjoint divisors impose no obstruction to achieving the maximal possible hierarchical depth: their order can be rearranged arbitrarily without changing the total number of steps.
Proof.
We first prove (1). Let $D_{a}\in\mathcal{A}$, $D_{b}\in\mathcal{B}$ and suppose $\varphi_{a}:\mathcal{E}\to\mathcal{Q}_{a}$, $\varphi_{b}:\mathcal{E}\to\mathcal{Q}_{b}$ are surjective with $\mathrm{Supp}(\mathcal{Q}_{a})\subset D_{a}$, $\mathrm{Supp}(\mathcal{Q}_{b})\subset D_{b}$. Set
$V_{a}:=\ker(\varphi_{a}),\qquad V_{b}:=\ker(\varphi_{b}).$
Define the successive transforms
$V_{ab}:=\ker\bigl(V_{a}\xrightarrow{\overline{\varphi}_{b}}\mathcal{Q}_{b}\bigr),\qquad V_{ba}:=\ker\bigl(V_{b}\xrightarrow{\overline{\varphi}_{a}}\mathcal{Q}_{a}\bigr),$
and the simultaneous kernel
$V:=\ker\bigl(\mathcal{E}\xrightarrow{(\varphi_{a},\varphi_{b})}\mathcal{Q}_{a}\oplus\mathcal{Q}_{b}\bigr).$
Since $\mathrm{Supp}(D_{a})\cap\mathrm{Supp}(D_{b})=\varnothing$, the torsion sheaves $\mathcal{Q}_{a}$ and $\mathcal{Q}_{b}$ are supported on disjoint closed subsets. Therefore, for any section $s\in\mathcal{E}(U)$ on an open set $U\subset X$,
$\varphi_{a}(s)=0\quad\text{and}\quad\varphi_{b}(s)=0$
are independent conditions. Consequently,
$V(U)$ $=\{s\in\mathcal{E}(U):\varphi_{a}(s)=0,\ \varphi_{b}(s)=0\},$
$V_{ab}(U)$ $=\{s\in\mathcal{E}(U):\varphi_{a}(s)=0,\ \varphi_{b}(s)=0\},$
$V_{ba}(U)$ $=\{s\in\mathcal{E}(U):\varphi_{b}(s)=0,\ \varphi_{a}(s)=0\}.$
Thus $V=V_{ab}=V_{ba}$ as subsheaves of $\mathcal{E}$, proving commutativity.
We now prove (2). Consider the given filtration
$\mathcal{E}=V_{0}\supset V_{1}\supset\cdots\supset V_{h},\qquad V_{k+1}=\ker\bigl(V_{k}\xrightarrow{\varphi_{k}}\mathcal{Q}_{k}\bigr).$
Label each index $k$ as type $A$ or type $B$ according to whether $\mathrm{Supp}(\mathcal{Q}_{k})\subset D_{a}\in\mathcal{A}$ or $\mathrm{Supp}(\mathcal{Q}_{k})\subset D_{b}\in\mathcal{B}$. If a step of type $A$ is immediately followed by a step of type $B$, say
$V_{k}\supset V_{k+1}\supset V_{k+2},$
with $V_{k+1}=\ker(V_{k}\to\mathcal{Q}_{a})$ and $V_{k+2}=\ker(V_{k+1}\to\mathcal{Q}_{b})$, then by the commutativity just proved, the composite kernel
$\ker\bigl(V_{k}\to\mathcal{Q}_{a}\bigr)\xrightarrow{}\mathcal{Q}_{b}$
is equal to
$\ker\bigl(V_{k}\to\mathcal{Q}_{b}\bigr)\xrightarrow{}\mathcal{Q}_{a},$
and hence the transforms may be exchanged:
$\ker\bigl(V_{k}\xrightarrow{\varphi_{a}}\mathcal{Q}_{a}\bigr)\ \xrightarrow{\ \overline{\varphi}_{b}\ }\ Q_{b}\quad=\quad\ker\bigl(V_{k}\xrightarrow{\varphi_{b}}\mathcal{Q}_{b}\bigr)\ \xrightarrow{\ \overline{\varphi}_{a}\ }\ Q_{a}.$
Thus one may swap the AA– and BB–steps without changing the resulting subsheaf of $V_{k}$. By repeatedly swapping all out–of–order adjacent pairs (type $A$ followed by type $B$), one eventually obtains a filtration
$\mathcal{E}=V^{\prime}_{0}\supset V^{\prime}_{1}\supset\cdots\supset V^{\prime}_{h}$
in which all AA–steps precede all BB–steps. The length remains $h$ because no step is removed during the swapping.
Finally, statement (3) is immediate: reordering does not change the number of transform steps, hence disjoint support never forces a reduction in hierarchical depth.
Example 3.5.
(1) Let $X=\mathbb{P}^{1}\times\mathbb{P}^{1}$. Write $F:=\pi_{1}^{-1}(pt)$. Distinct fibers $F_{x}:=\pi_{1}^{-1}(x)$ for $x\in\mathbb{P}^{1}$ are pairwise disjoint effective Cartier divisors. Consider the split bundle
$\mathcal{E}=\bigoplus_{j=1}^{s}\mathcal{O}_{X}(a_{j}F),\qquad a_{j}\geq 0,\qquad d:=\sum_{j=1}^{s}a_{j}.$
For each summand $\mathcal{O}_{X}(a_{j}F)$ choose $a_{j}$ distinct points $x_{j,1},\dots,x_{j,a_{j}}\in\mathbb{P}^{1}$ such that all points $x_{j,t}$ are distinct as $(j,t)$ vary, and write
$a_{j}F=F_{x_{j,1}}+\cdots+F_{x_{j,a_{j}}}.$
Each unit inclusion
$\mathcal{O}_{X}((k{-}1)F)\subset\mathcal{O}_{X}(kF)$
is realized by an elementary transform along the divisor $F_{x_{j,k}}$. Since all fibers $F_{x_{j,t}}$ have pairwise disjoint supports, Lemma 3.2 implies that the corresponding elementary transforms commute, and hence Corollary 3.4 applies.
Thus the $d$ unit increments coming from all summands may be interleaved into a single saturated hierarchical filtration of $\mathcal{E}$. Therefore
$h_{\mathcal{O}_{X}}(\mathcal{E})=d.$
Actually, the class $F$ admits arbitrarily many pairwise disjoint effective representatives, so all elementary transforms occur along disjoint supports and therefore commute.
(2) Let $X=\mathbb{P}^{2}$ and let $H$ be the hyperplane class. Every effective divisor in $|H|$ is a line, and any two distinct lines meet in exactly one point. Hence no two nonzero divisors in $|H|$ have disjoint support.
For example, consider
$\mathcal{E}=\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(aH),\qquad a\geq 2.$
Although $\det(\mathcal{E})\simeq\mathcal{O}_{\mathbb{P}^{2}}(aH)$, the natural chain
$\mathcal{O}_{\mathbb{P}^{2}}\subset\mathcal{O}_{\mathbb{P}^{2}}(H)\subset\cdots\subset\mathcal{O}_{\mathbb{P}^{2}}(aH)$
cannot be realized by $a$ successive saturated elementary transforms inside $\mathcal{E}$. Each such step must occur along a divisor of type $H$, but every pair of such divisors intersects. Thus the elementary transforms fail to commute, and repeated kernels interact by saturation.
By Lemma 3.2, only transforms along disjoint divisors commute; here all supports intersect, so Corollary 3.4 does not apply. Consequently the unit–increment strategy collapses, and one typically obtains only a short saturated same–rank filtration such as
$\mathcal{O}_{\mathbb{P}^{2}}^{\oplus 2}\subset\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(aH)\subset\mathcal{E},$
of length22, with no possibility of refining it to length $a$.
Here the class $H$ has no pairwise disjoint effective representatives, so elementary transforms must occur along intersecting divisors and hence do not commute; saturation then forces the filtration length to collapse.
Consequently, these two surfaces illustrate the dichotomy between the numerical determinant bound $h(\mathcal{E})\leq d$ and the geometric possibility of achieving equality. If the linear system $|A|$ contains sufficiently many pairwise disjoint effective representatives (as on $\mathbb{P}^{1}\times\mathbb{P}^{1}$), one may realize each unit increment separately and obtain $h(\mathcal{E})=d$. If $|A|$ contains no such disjoint representatives (as on $\mathbb{P}^{2}$ with $A=H$), then geometric obstructions force noncommuting transforms and the maximal hierarchical depth may be strictly smaller than $d$.
4. Hierarchical depth and determinant bounds
In this section we collect basic numerical properties of hierarchical depth and establish general bounds in terms of the determinant. These results provide effective control of the invariant in simple geometric situations and will be used throughout the paper.
Proposition 4.1.
Let $C$ be a smooth projective curve and let
$\mathcal{E}=\bigoplus_{i=1}^{r}\mathcal{O}_{C}(d_{i})$
be a split vector bundle of rank $r$, with integers $d_{i}\in\mathbb{Z}$. Fix a normalization line bundle $\Lambda_{0}\in\mathrm{Pic}(C)$. Set
$M:=\deg\det(\mathcal{E})-\deg(\Lambda_{0})=\sum_{i=1}^{r}d_{i}-\deg(\Lambda_{0}).$
  1. (a)
    If $M<0$ then there exists no hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$, and by convention $h_{\Lambda_{0}}(\mathcal{E})=$-$\infty$.
  2. (b)
    If $M\geq 0$ then the hierarchical depth of $\mathcal{E}$ with respect to $\Lambda_{0}$ equals
    $h_{\Lambda_{0}}(\mathcal{E})=M.$
    In particular, when $\Lambda_{0}=\mathcal{O}_{C}$ one has
    $h(\mathcal{E})=\sum_{i=1}^{r}d_{i}.$
Proof.
We assume $M\geq 0$; the negative case (no filtration exists) is immediate from the determinant relation and the normalization requirement.
Let
$\mathfrak{F}:\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
be any hierarchical filtration normalized by $\Lambda_{0}$, i.e. $\det(\mathcal{E}_{0})\simeq\Lambda_{0}$ and for each $i$
$\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{C}(p_{i})$
for some effective divisor $p_{i}$ (a point of $C$, counted with multiplicity). Taking determinants and degrees yields
$\deg\det(\mathcal{E})-\deg(\Lambda_{0})=\sum_{i=1}^{h}\deg(p_{i})=h,$
since each $p_{i}$ is a single point (degree $1$). Hence $h\leq M$.
We now produce a hierarchical filtration of length exactly $M$.
Choose $M$ distinct points $q_{1},\dots,q_{M}\in C$. (If the base field is finite and $C(\mathbb{F}_{q})$ has fewer than $M$ rational points, work over a finite extension or choose geometric points; the statement and proof are unchanged up to base change.) For each $j$ define inductively a subsheaf by performing an elementary transform at $q_{j}$:
$\mathcal{E}^{(M)}:=\mathcal{E},\quad\mathcal{E}^{(M-1)}:=\ker\bigl(\mathcal{E}^{(M)}\xrightarrow{\ \rho_{M}\ }k(q_{M})\bigr),\quad\mathcal{E}^{(M-2)}:=\ker\bigl(\mathcal{E}^{(M-1)}\xrightarrow{\ \rho_{M-1}\ }k(q_{M-1})\bigr),$
and so on, where each $\rho_{j}$ is a chosen surjection of the fibre at $q_{j}$ onto the one-dimensional skyscraper $k(q_{j})$.
On a curve the kernel of a surjection from a vector bundle to a skyscraper is again locally free (torsion-free on a curve implies locally free), so each $\mathcal{E}^{(t)}$ is a vector bundle of rank $r$ and sits as a saturated subsheaf of $\mathcal{E}^{(t+1)}$. Moreover the determinant relation at each step is
$\det(\mathcal{E}^{(t+1)})\simeq\det(\mathcal{E}^{(t)})\otimes\mathcal{O}_{C}(q_{t+1}).$
After performing the $M$ transforms we obtain
$\det(\mathcal{E}^{(M)})\simeq\det(\mathcal{E}^{(0)})\otimes\mathcal{O}_{C}\Big(\sum_{j=1}^{M}q_{j}\Big).$
Rearranging,
$\det(\mathcal{E}^{(0)})\simeq\det(\mathcal{E})\otimes\mathcal{O}_{C}\Big(-\sum_{j=1}^{M}q_{j}\Big).$
By construction the right hand side is a line bundle of degree $\deg\det(\mathcal{E})-M=\deg\Lambda_{0}$; hence $\det(\mathcal{E}^{(0)})\simeq\Lambda_{0}$. Thus the ascending chain
$\mathcal{E}^{(0)}\subset\mathcal{E}^{(1)}\subset\cdots\subset\mathcal{E}^{(M)}=\mathcal{E}$
is a hierarchical filtration normalized by $\Lambda_{0}$ of length $M$.
It remains to check that the $M$ steps are distinct (no collapse via saturation). Because we chose the points $q_{j}$ to be pairwise distinct, the successive elementary transforms are supported on distinct points and therefore commute and produce distinct saturated kernels; equivalently, each step increases the determinant by a nonzero effective divisor, so no further identification occurs. Hence the length of the constructed filtration is exactly $M$.
Combining the upper bound and the explicit construction shows $h_{\Lambda_{0}}(\mathcal{E})=M$, as claimed. ∎
Remark 4.2.
Proposition 4.1 is the one–dimensional specialization of Proposition 2.4. When $\dim X=1$, every divisor linearly equivalent to $A$ is a single point and elementary transforms at distinct points commute. Hence the geometric existence condition in the higher–dimensional proposition is automatic for split bundles, and the determinant bound specializes to an exact formula for curves.
Proposition 4.3.
Let $X$ be a smooth projective variety, fix a reference line bundle $\Lambda_{0}\in\mathrm{Pic}(X)$, and let $\mathcal{E}$ be a vector bundle of rank $r$ on $X$.
  1. (A)
    Suppose $\mathcal{E}^{\prime}\subset\mathcal{E}$ is a saturated inclusion of locally free sheaves of the same rank $r$, so that $\mathcal{E}/\mathcal{E}^{\prime}$ is a nonzero torsion sheaf. Assume $\det(\mathcal{E}^{\prime})$ is measured relative to the same normalization $\Lambda_{0}$ as $\det(\mathcal{E})$. Then
    $h_{\Lambda_{0}}(\mathcal{E}^{\prime})\leq h_{\Lambda_{0}}(\mathcal{E}).$
  2. (B)
    If $\mathcal{E}^{\prime}\subset\mathcal{E}$ is a subbundle of strictly smaller rank $\mathrm{rank}(\mathcal{E}^{\prime})<\mathrm{rank}(\mathcal{E})$, no inequality between $h_{\Lambda_{0}}(\mathcal{E}^{\prime})$ and $h_{\Lambda_{0}}(\mathcal{E})$ holds in general. In particular, one may have $h_{\Lambda_{0}}(\mathcal{E}^{\prime})>h_{\Lambda_{0}}(\mathcal{E})$.
  3. (C)
    Let
    $0\longrightarrow\mathcal{E}^{\prime}\longrightarrow\mathcal{E}\longrightarrow\mathcal{E}^{\prime\prime}\longrightarrow 0$
    be a short exact sequence of vector bundles. Assume that:
    1. (i)
      the hierarchical filtration of $\mathcal{E}^{\prime}$ is normalized by $\Lambda_{0}$;
    2. (ii)
      the hierarchical filtration of $\mathcal{E}^{\prime\prime}$ is normalized by $\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})$; and
    3. (iii)
      every determinant increment divisor of the $\mathcal{E}^{\prime\prime}$–filtration lifts to a saturated same–rank elementary transform of $\mathcal{E}$ (equivalently: the effective divisors involved in the filtration of $\mathcal{E}^{\prime\prime}$ are disjoint from the support of the torsion in successive pushforwards during the reconstruction of $\mathcal{E}$).
    Under these hypotheses the concatenation of hierarchical filtrations yields a hierarchical filtration of $\mathcal{E}$, and therefore
    $h_{\Lambda_{0}}(\mathcal{E})\geq h_{\Lambda_{0}}(\mathcal{E}^{\prime})+h_{\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})}(\mathcal{E}^{\prime\prime}),$
    and in particular
    (3) $h_{\Lambda_{0}}(\mathcal{E})\geq\max\Big\{h_{\Lambda_{0}}(\mathcal{E}^{\prime}),h_{\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})}(\mathcal{E}^{\prime\prime})\Big\}.$
  4. (D)
    If $\mathcal{E}=\mathcal{E}_{1}\oplus\mathcal{E}_{2}$, and the determinant increments appearing in maximal hierarchical filtrations of $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ admit representatives whose divisors have disjoint support (so that the corresponding elementary transforms commute), then their interleaving produces a hierarchical filtration of $\mathcal{E}$ and
    $h_{\Lambda_{0}}(\mathcal{E})\geq h_{\Lambda_{0}}(\mathcal{E}_{1})+h_{\Lambda_{0}}(\mathcal{E}_{2}).$
    Without the disjoint–support (or commutativity) hypothesis the inequality may be strict or may fail in the naive additive form.
Proof.
(A) A hierarchical filtration of $\mathcal{E}^{\prime}$ consists of a sequence of saturated same–rank inclusions
$\mathcal{E}^{\prime}_{0}\subset\mathcal{E}^{\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime}_{h^{\prime}}=\mathcal{E}^{\prime},$
where each quotient is torsion supported on an effective divisor. Since $\mathcal{E}^{\prime}\subset\mathcal{E}$ is saturated of the same rank, each inclusion $\mathcal{E}^{\prime}_{i}\subset\mathcal{E}$ remains saturated in $\mathcal{E}$ and preserves the torsion quotients. Hence the same chain is a valid hierarchical filtration of $\mathcal{E}$. This proves $h_{\Lambda_{0}}(\mathcal{E}^{\prime})\leq h_{\Lambda_{0}}(\mathcal{E})$.
(B) Let $X=C$ be a smooth projective curve and normalize by $\Lambda_{0}=\mathcal{O}_{C}$. Monotonicity of hierarchical depth need not hold when passing to a proper subbundle of strictly smaller rank: one can have a subbundle $\mathcal{E}^{\prime}\subset\mathcal{E}$ with $\mathrm{rank}(\mathcal{E}^{\prime})<\mathrm{rank}(\mathcal{E})$ yet
$h(\mathcal{E}^{\prime})>h(\mathcal{E}).$
A minimal explicit example on $\mathbb{P}^{1}$ is as follows.
$\mathcal{E}^{\prime}=\mathcal{O}_{\mathbb{P}^{1}}(1),\qquad\mathcal{E}=\mathcal{O}_{\mathbb{P}^{1}}(1)\oplus\mathcal{O}_{\mathbb{P}^{1}}(-1),\qquad\mathcal{E}^{\prime}\hookrightarrow\mathcal{E}\text{ the summand inclusion.}$
Then
$\det(\mathcal{E})=\mathcal{O}_{\mathbb{P}^{1}}(1)\otimes\mathcal{O}_{\mathbb{P}^{1}}(-1)\cong\mathcal{O}_{\mathbb{P}^{1}},$
so by Proposition 4.1,
$h(\mathcal{E})=\deg\det(\mathcal{E})=0,$
whereas
$h(\mathcal{E}^{\prime})=\deg\mathcal{O}_{\mathbb{P}^{1}}(1)=1.$
Therefore $h(\mathcal{E}^{\prime})=1>0=h(\mathcal{E})$, showing that hierarchical depth is not monotone under subbundles of smaller rank.
(C) Let
$0\to\mathcal{E}^{\prime}\xrightarrow{\iota}\mathcal{E}\xrightarrow{\pi}\mathcal{E}^{\prime\prime}\to 0$
be exact. By assumption (i) there exists a normalized hierarchical filtration of $\mathcal{E}^{\prime}$,
$\mathcal{E}^{\prime}_{0}\subset\mathcal{E}^{\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime}_{a}=\mathcal{E}^{\prime},$
with $\det(\mathcal{E}^{\prime}_{0})\simeq\Lambda_{0}$ and determinant increments $\det(\mathcal{E}^{\prime}_{i})\simeq\det(\mathcal{E}^{\prime}_{i-1})\otimes\mathcal{O}_{X}(D^{\prime}_{i})$. By assumption (ii) there exists a hierarchical filtration of $\mathcal{E}^{\prime\prime}$ normalized by $\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})$,
$\mathcal{E}^{\prime\prime}_{0}\subset\mathcal{E}^{\prime\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime\prime}_{b}=\mathcal{E}^{\prime\prime},$
with determinant increments $\mathcal{O}_{X}(D^{\prime\prime}_{j})$.
Hypothesis (iii) ensures that each divisor $D^{\prime\prime}_{j}$ yields an elementary transform of $\mathcal{E}$ lifting the elementary transform of $\mathcal{E}^{\prime\prime}$: the reason is that, under disjoint support from the torsion and the universal property of elementary transforms (cf. Huybrechts–Lehn, Prop. 1.1.8), there exist saturated inclusions
$\mathcal{E}^{(j-1)}\subset\mathcal{E}^{(j)}\subset\mathcal{E}^{(j-1)}(D^{\prime\prime}_{j})$
with torsion quotient supported on $D^{\prime\prime}_{j}$, and whose pushforward to $\mathcal{E}^{\prime\prime}$ reproduces the corresponding step of the $\mathcal{E}^{\prime\prime}$–filtration.
Therefore, after exhausting the $\mathcal{E}^{\prime}$–filtration inside $\mathcal{E}$ using (A), we may lift each step of the $\mathcal{E}^{\prime\prime}$–filtration to $\mathcal{E}$, producing a chain
$\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{a+b}=\mathcal{E}$
of saturated same–rank inclusions. This is a legitimate hierarchical filtration of $\mathcal{E}$ of length $a+b$. Thus
$h_{\Lambda_{0}}(\mathcal{E})\geq a+b=h_{\Lambda_{0}}(\mathcal{E}^{\prime})+h_{\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})}(\mathcal{E}^{\prime\prime}),$
and (3) follows immediately.
(D) If $\mathcal{E}=\mathcal{E}_{1}\oplus\mathcal{E}_{2}$ and both bundles admit maximal hierarchical filtrations whose determinant increments correspond to divisors with pairwise disjoint support, then elementary transforms with disjoint support commute, so the two filtrations can be interleaved to produce a filtration of $\mathcal{E}$ whose length is the sum of the two lengths. The claimed inequality follows. If supports intersect, such commutativity need not hold, and ranks may drop under saturation, so the simple additive formula may fail. ∎
5. Effects of the Minimal Model Program on hierarchical filtrations
In this section we study the behavior of hierarchical filtrations of vector bundles under birational transformations arising in the Minimal Model Program (MMP) when the base variety has dimension at most $2$. Since the MMP for smooth varieties of dimension $\leq 2$ preserves smoothness, we may restrict throughout to smooth projective curves and surfaces.
5.1. Setup and motivation
Let $X$ be a smooth projective variety of dimension $n\leq 2$ over an algebraically closed field $\mathbb{F}$. The MMP for $X$ consists of a finite sequence of birational morphisms
$X=X_{0}\dashrightarrow X_{1}\dashrightarrow\cdots\dashrightarrow X_{t}=X_{\min},$
where each step is either an isomorphism in codimension $1$ or a contraction of a (1)(-1)–curve (in the surface case), and the final model $X_{\min}$ is minimal, i.e. $K_{X_{\min}}$ is nef. In dimension $1$ the sequence is trivial since every smooth curve is already minimal (see [10]).
Given a vector bundle $\mathcal{E}$ on $X$, we are interested in how its hierarchical filtrations behave under the birational maps of such an MMP sequence, and in particular how the hierarchical depth changes after blowing up or contracting divisors. This analysis provides a bridge between the birational geometry of $X$ and the algebraic structure of $\mathcal{E}$, and will later underlie the construction of evaluation codes associated to successive divisorial increments.
5.2. Birational pullback of filtrations
Let $f:Y\to X$ be a birational morphism between smooth projective varieties of dimension $\leq 2$, for instance the blow–up of $X$ at a smooth point with exceptional divisor $E$. Given a vector bundle $\mathcal{E}$ on $X$ and a reference line bundle $\Lambda_{0}$, consider its pullback $f^{*}\mathcal{E}$ on $Y$. The determinant relation satisfies
$\det(f^{*}\mathcal{E})\simeq f^{*}(\det\mathcal{E}),\qquad c_{1}(\det(f^{*}\mathcal{E}))=f^{*}c_{1}(\det\mathcal{E}).$
Hence a hierarchical filtration
$\mathfrak{F}:\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
on $X$ normalized by $\Lambda_{0}$ pulls back naturally to a filtration
$f^{*}\mathfrak{F}:\ f^{*}\mathcal{E}_{0}\subset f^{*}\mathcal{E}_{1}\subset\cdots\subset f^{*}\mathcal{E}_{h}=f^{*}\mathcal{E}$
on $Y$, normalized by $f^{*}\Lambda_{0}$. At each step one has
$\det(f^{*}\mathcal{E}_{i})\simeq\det(f^{*}\mathcal{E}_{i-1})\otimes\mathcal{O}_{Y}(f^{*}D_{i}),$
so that the divisorial increments are transformed by pullback: $D_{i}$ is replaced by $f^{*}D_{i}$. Consequently,
$h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})\geq h_{\Lambda_{0}}(\mathcal{E}).$
In particular, the hierarchical depth is invariant under birational pullback of smooth varieties in dimension $\leq 2$. Consequently, we have the following statement:
Proposition 5.1.
Let $f:Y\to X$ be a birational morphism between smooth projective varieties of dimension $\leq 2$, and let $\mathcal{E}$ be a vector bundle on $X$. Then
$h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})\geq h_{\Lambda_{0}}(\mathcal{E}).$
5.3. Effect of blow–ups on new divisorial increments
Let $f:Y\to X$ be the blow–up of $X$ at a smooth point $p\in X$ with exceptional divisor $E\simeq\mathbb{P}^{n-1}$. If $\mathcal{E}$ is a vector bundle on $X$, the pullback $f^{*}\mathcal{E}$ remains locally free. However, the Picard group changes:
$\mathrm{Pic}(Y)\cong f^{*}\mathrm{Pic}(X)\oplus\mathbb{Z}\langle E\rangle.$
Therefore, new effective divisors supported on the exceptional locus can appear in hierarchical filtrations of $f^{*}\mathcal{E}$ that have no counterpart on $X$.
The statements of two following lemmas can be found in [5, Ch. V].
Lemma 5.2.
Let $X$ be a smooth projective variety over an algebraically closed field, let $p\in X$ be a smooth (closed) point, and let $f:\,Y=\operatorname{Bl}_{p}X\to X$ be the blow–up of $X$ at $p$. Let $E\subset Y$ denote the exceptional divisor (which is isomorphic to $\mathbb{P}^{\,\dim X-1}$). Then the pullback map $f^{*}:\mathrm{Pic}(X)\to\mathrm{Pic}(Y)$ is injective, and every line bundle $\mathcal{L}\in\mathrm{Pic}(Y)$ can be written uniquely in the form
$\mathcal{L}\simeq f^{*}M\otimes\mathcal{O}_{Y}(mE)$
for a unique $M\in\mathrm{Pic}(X)$ and a unique integer $m\in\mathbb{Z}$. Equivalently, there is a direct sum decomposition of abelian groups
$\mathrm{Pic}(Y)\cong f^{*}\mathrm{Pic}(X)\oplus\mathbb{Z}\langle E\rangle.$
Proof.
Injectivity of $f^{*}$. Suppose $M\in\mathrm{Pic}(X)$ satisfies $f^{*}M\cong\mathcal{O}_{Y}$. Apply the pushforward functor $f_{*}$ and use the projection formula together with $f_{*}\mathcal{O}_{Y}\cong\mathcal{O}_{X}$ (since $f$ is a birational morphism contracting the exceptional divisor to a point):
$f_{*}f^{*}M\cong M\otimes f_{*}\mathcal{O}_{Y}\cong M.$
But $f_{*}f^{*}M\cong f_{*}\mathcal{O}_{Y}\cong\mathcal{O}_{X}$ because $f^{*}M\cong\mathcal{O}_{Y}$, hence $M\cong\mathcal{O}_{X}$. Thus $f^{*}$ is injective.
Existence of the decomposition. Let $\mathcal{L}\in\mathrm{Pic}(Y)$. Restrict $\mathcal{L}$ to the exceptional divisor $E\cong\mathbb{P}^{\,\dim X-1}$. Since $\mathrm{Pic}(E)\cong\mathbb{Z}$ and is generated by $\mathcal{O}_{E}(1)$, there is a unique integer $m$ with
$\mathcal{L}|_{E}\cong\mathcal{O}_{E}(m).$
Consider the twist $\mathcal{L}(-mE):=\mathcal{L}\otimes\mathcal{O}_{Y}(-mE)$. By construction $\mathcal{L}(-mE)|_{E}\cong\mathcal{O}_{E}$ is trivial on $E$. The line bundle $\mathcal{L}(-mE)$ is therefore trivial on the exceptional locus, hence it descends to $X$ in the following sense: set
$M:=f_{*}\big(\mathcal{L}(-mE)\big).$
The sheaf $M$ is reflexive of rank $1$ on the smooth variety $X$, hence locally free of rank $1$, i.e. a line bundle (see e.g. [5, II, Ex. 5.15]). Moreover $f^{*}M$ and $\mathcal{L}(-mE)$ are isomorphic away from $E$. Both are line bundles on the normal variety $Y$ and agree on the complement of a divisor, hence they are isomorphic globally:
$f^{*}M\cong\mathcal{L}(-mE).$
Therefore $\mathcal{L}\cong f^{*}M\otimes\mathcal{O}_{Y}(mE)$.
Uniqueness. If $f^{*}M\otimes\mathcal{O}_{Y}(mE)\cong f^{*}M^{\prime}\otimes\mathcal{O}_{Y}(m^{\prime}E)$, restrict to $E$ to compare integers: the left side restricts to $\mathcal{O}_{E}(m)$ and the right side to $\mathcal{O}_{E}(m^{\prime})$, hence $m=m^{\prime}$. Cancelling $\mathcal{O}_{Y}(mE)$ and using injectivity of $f^{*}$ gives $M\cong M^{\prime}$. This proves uniqueness of the pair $(M,m)$.
Group decomposition. The map
$f^{*}\mathrm{Pic}(X)\oplus\mathbb{Z}\langle E\rangle\longrightarrow\mathrm{Pic}(Y),\qquad(M,m)\longmapsto f^{*}M\otimes\mathcal{O}_{Y}(mE)$
is therefore bijective and respects the group law, yielding the claimed direct sum decomposition. ∎
Lemma 5.3.
Let $D$ be an effective Cartier divisor on $Y$. Then $D$ decomposes uniquely as
$D=f^{*}D_{X}+mE$
with $D_{X}$ an effective Cartier divisor on $X$ and $m\in\mathbb{Z}_{\geq 0}$. Moreover $D$ is supported on $E$ if and only if $D_{X}=0$.
Proof.
Write the line bundle $\mathcal{O}_{Y}(D)$ as $f^{*}\mathcal{O}_{X}(D_{X})\otimes\mathcal{O}_{Y}(mE)$ using Lemma 5.2. If $\mathcal{O}_{Y}(D)$ is effective then pulling down by $f_{*}$ shows $\mathcal{O}_{X}(D_{X})$ is effective and $m\geq 0$ (otherwise the section would vanish identically along the exceptional divisor in contradiction with effectivity). Uniqueness follows from the uniqueness in Lemma 5.2. ∎
Recall the concept of saturation from [8]. Let $\mathcal{F}$ be a subsheaf of a coherent sheaf $\mathcal{E}$ on $X$. The saturation of $\mathcal{F}$ inside $\mathcal{E}$ is defined as the minimal subsheaf $\mathcal{F}^{\prime}$ containing $\mathcal{F}$ such that $\mathcal{E}/\mathcal{F}^{\prime}$ is pure of dimension $d=\dim(\mathcal{E})$ or zero. Equivalently, the saturation of $\mathcal{F}$ is
$\mathcal{F}^{\mathrm{sat}}:=\ker\bigl(\mathcal{E}\longrightarrow\mathcal{E}/\mathcal{F}\longrightarrow(\mathcal{E}/\mathcal{F})/T(\mathcal{E}/\mathcal{F})\bigr),$
where $T(\mathcal{E}/\mathcal{F})$ denotes the maximal subsheaf of $\mathcal{E}/\mathcal{F}$ of dimension $\leq d-1$. Equivalently, $\mathcal{F}^{\mathrm{sat}}$ is the largest subsheaf of $\mathcal{E}$ that coincides with $\mathcal{F}$ on a dense open set and is torsion–free. When $\mathcal{E}=\mathcal{O}_{X}$, saturation simply removes 0–dimensional torsion.
As an example, suppose the sequence of line bundles
$\mathcal{O}_{Y}(-2E)\subset\mathcal{O}_{Y}(-E)\subset\mathcal{O}_{Y}.$
Applying $f_{*}$ gives
$f_{*}\mathcal{O}_{Y}(-2E)=\mathcal{I}_{p/X}^{2},\qquad f_{*}\mathcal{O}_{Y}(-E)=\mathcal{I}_{p/X},\qquad f_{*}\mathcal{O}_{Y}=\mathcal{O}_{X}.$
Each of these is already torsion-free on $X$ (they are ideal sheaves of a smooth point), hence saturation does nothing:
$(\mathcal{I}_{p/X}^{2})^{\mathrm{sat}}=\mathcal{I}_{p/X}^{2},\qquad(\mathcal{I}_{p/X})^{\mathrm{sat}}=\mathcal{I}_{p/X},\qquad(\mathcal{O}_{X})^{\mathrm{sat}}=\mathcal{O}_{X}.$
Thus the saturated chain is simply
$\mathcal{I}_{p/X}^{2}\subset\mathcal{I}_{p/X}\subset\mathcal{O}_{X}.$
Proposition 5.4.
Let $f:Y\to X$ be the blowup of a smooth projective variety $X$ at a smooth closed point $p\in X$, and denote by $E\subset Y$ the exceptional divisor. Assume $\dim X\leq 2$. Let $\mathcal{E}$ be a vector bundle on $X$ and fix a normalization $\Lambda_{0}\in\mathrm{Pic}(X)$.
Then every hierarchical filtration of $f^{*}\mathcal{E}$ normalized by $f^{*}\Lambda_{0}$ has length
$h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})=h_{\Lambda_{0}}(\mathcal{E})+\delta,\qquad\delta\in\mathbb{Z}_{\geq 0},$
where $\delta$ equals the number of additional determinant increments in a maximal filtration of $f^{*}\mathcal{E}$ whose divisors carry a positive coefficient of the exceptional divisor $E$. In particular $\delta=0$ unless the filtration involves elementary transforms along $E$.
Proof.
Step 1. If
$\mathfrak{F}_{X}:\ \mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
is a hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$, then pulling back termwise yields
$f^{*}\mathfrak{F}_{X}:\ f^{*}\mathcal{E}_{0}\subset f^{*}\mathcal{E}_{1}\subset\cdots\subset f^{*}\mathcal{E}_{h}=f^{*}\mathcal{E},$
a hierarchical filtration of $f^{*}\mathcal{E}$ normalized by $f^{*}\Lambda_{0}$. It follows from Proposition 5.1 that
$h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})\ \geq\ h_{\Lambda_{0}}(\mathcal{E}).$
Thus any extra length in $h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})$ must come from steps whose torsion quotient is supported on $E$.
Step 2. Let
$\mathfrak{F}_{Y}:\ \mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\cdots\subset\mathcal{F}_{h^{\prime}}=f^{*}\mathcal{E}$
be a maximal hierarchical filtration of $f^{*}\mathcal{E}$ normalized by $f^{*}\Lambda_{0}$. For each step write the determinant increment as an effective Cartier divisor $D_{i}\subset Y$. By the decomposition of effective divisors on a blowup (Lemma 5.3), each $D_{i}$ splits uniquely as
$D_{i}=f^{*}D_{i,X}+m_{i}E,\qquad D_{i,X}\geq 0,\ m_{i}\in\mathbb{Z}_{\geq 0}.$
Collect the indices into two disjoint sets:
$I_{0}:=\{i:\ m_{i}=0\},\qquad I_{E}:=\{i:\ m_{i}>0\},$
and write $h_{0}:=|I_{0}|$ and $\delta:=|I_{E}|$. Then $h^{\prime}=h_{0}+\delta$.
Step 3. For each $i\in I_{0}$ the divisor $D_{i}$ is the pullback of an effective Cartier divisor $D_{i,X}$ on $X$. The corresponding step $\mathcal{F}_{i-1}\subset\mathcal{F}_{i}$ on $Y$ therefore descends (objectwise) to an inclusion of torsion–free sheaves on $X$ after applying pushforward and saturation:
$\mathcal{G}_{i}:=\big(f_{*}\mathcal{F}_{i}\big)^{\mathrm{sat}}.$
Because $\dim X\leq 2$ the saturation (reflexive hull) of a torsion-free rank-$r$ sheaf is locally free (see for instance [5, II, Ex. 5.15] or standard references on reflexive sheaves). Hence each $\mathcal{G}_{i}$ is a vector bundle on $X$, and the inclusions $\mathcal{G}_{i-1}\subset\mathcal{G}_{i}$ are saturated with torsion quotients supported on $D_{i,X}$. Collecting these descended steps yields a hierarchical filtration on $X$ of length $h_{0}$, normalized by $\Lambda_{0}$. Consequently
$h_{0}\leq h_{\Lambda_{0}}(\mathcal{E}).$
Conversely, pulling back any maximal filtration on $X$ reproduces a filtration on $Y$ whose non-exceptional increments give at least $h_{\Lambda_{0}}(\mathcal{E})$ steps with $m_{i}=0$. Therefore $h_{0}\geq h_{\Lambda_{0}}(\mathcal{E})$, and combining the two inequalities gives
$h_{0}=h_{\Lambda_{0}}(\mathcal{E}).$
Step 4. By Step 2 we have $h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})=h_{0}+\delta$, and by Step 3, $h_{0}=h_{\Lambda_{0}}(\mathcal{E})$. Hence
$h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})=h_{\Lambda_{0}}(\mathcal{E})+\delta,$
with $\delta\geq 0$ equal to the number of determinant increments whose divisors carry a positive coefficient of $E$.
Step 5. Finally, each elementary transform along $E$ produces a torsion quotient supported on $E$ and increases the exceptional coefficient of the determinant by precisely the corresponding multiplicity; conversely any positive coefficient of $E$ in some increment $D_{i}$ arises from at least one such elementary transform (after saturation). Therefore the number $\delta$ equals the number of additional elementary transforms supported on $E$ in the maximal filtration, i.e. the number of extra steps contributed by exceptional transforms.
This completes the proof. ∎
5.4. Contraction and preservation of depth
Proposition 5.5.
Let $f:Y\to X$ be the contraction of a (1)(-1)–curve $E\subset Y$ onto a smooth point $p=f(E)\in X$, and let $\mathcal{E}_{Y}$ be a vector bundle on $Y$ admitting a hierarchical filtration
$\mathfrak{F}_{Y}:\quad 0=\mathcal{E}_{0,Y}\subset\mathcal{E}_{1,Y}\subset\cdots\subset\mathcal{E}_{h,Y}=\mathcal{E}_{Y}$
normalized by $\Lambda_{0,Y}\in\mathrm{Pic}(Y)$. Assume that for every step of the filtration, the determinant increment divisor $D_{i,Y}$ satisfies
$E\not\subset\mathrm{Supp}D_{i,Y}.$
Then the direct image $\mathcal{E}:=(f_{*}\mathcal{E}_{Y})^{\mathrm{sat}}$ carries a natural hierarchical filtration
$\mathfrak{F}_{X}:\quad 0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E},$
normalized by $\Lambda_{0}:=\big(f_{*}\Lambda_{0,Y}\big)^{\mathrm{sat}}\in\mathrm{Pic}(X)$, and the hierarchical depths coincide:
$h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,Y}}(\mathcal{E}_{Y}).$
Proof.
We proceed in several steps.
Step 1. The morphism $f:Y\to X$ is an isomorphism over $U:=X\setminus\{p\}$, and $f^{-1}(U)=Y\setminus E\simeq U$. For any locally free sheaf $\mathcal{G}$ on $Y$, the pushforward $f_{*}\mathcal{G}$ is torsion–free on $X$, and since $f$ is birational between smooth surfaces, $f_{*}\mathcal{G}$ is reflexive (see, e.g., [6, Prop. 1.6]). Because reflexive sheaves on smooth surfaces are locally free, $(f_{*}\mathcal{G})^{\vee\vee}\cong f_{*}\mathcal{G}$ defines a vector bundle on $X$ (after saturation at the point $p$). For each $i$ we set
$\mathcal{E}_{i}:=(f_{*}\mathcal{E}_{i,Y})^{\mathrm{sat}}.$
Then each $\mathcal{E}_{i}$ is a vector bundle on $X$, and $f^{*}\mathcal{E}_{i}\simeq\mathcal{E}_{i,Y}$ away from $E$.
Step 2. Each inclusion $\mathcal{E}_{i-1,Y}\hookrightarrow\mathcal{E}_{i,Y}$ is saturated on $Y$, with torsion quotient $\mathcal{T}_{i,Y}$ supported on the effective Cartier divisor $D_{i,Y}$. Since $E\not\subset\mathrm{Supp}D_{i,Y}$, the support of $\mathcal{T}_{i,Y}$ avoids $E$. Applying $f_{*}$ yields an exact sequence on $U$:
$0\to f_{*}\mathcal{E}_{i-1,Y}|_{U}\to f_{*}\mathcal{E}_{i,Y}|_{U}\to f_{*}\mathcal{T}_{i,Y}|_{U}\to 0,$
where $f_{*}\mathcal{T}_{i,Y}$ remains a torsion sheaf supported on $f(D_{i,Y})\subset X\setminus\{p\}$. Passing to saturations gives a global inclusion of vector bundles
$\mathcal{E}_{i-1}\hookrightarrow\mathcal{E}_{i}$
on $X$.
Step 3. On $Y$ one has
$\det(\mathcal{E}_{i,Y})\simeq\det(\mathcal{E}_{i-1,Y})\otimes\mathcal{O}_{Y}(D_{i,Y}).$
Because $E\not\subset\mathrm{Supp}D_{i,Y}$, each $D_{i,Y}$ is the strict transform of a unique effective divisor $D_{i,X}\subset X$, i.e. $D_{i,Y}=f^{*}D_{i,X}$. Over $U$ the projection formula gives
$\det(f_{*}\mathcal{E}_{i,Y})|_{U}\simeq\det(f_{*}\mathcal{E}_{i-1,Y})|_{U}\otimes\mathcal{O}_{U}(D_{i,X}),$
and since taking reflexive hulls does not alter determinant line bundles on smooth surfaces, this equality extends globally to $X$:
$\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i,X}).$
Thus each step $\mathcal{E}_{i-1}\subset\mathcal{E}_{i}$ on $X$ is again an elementary transform with determinant increment $D_{i,X}$.
Step 4. The chain
$\mathfrak{F}_{X}:\quad 0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
is normalized by $\Lambda_{0}:=(f_{*}\Lambda_{0,Y})^{\mathrm{sat}}$, because $\det(\mathcal{E}_{0,Y})\simeq\Lambda_{0,Y}$ implies $\det(\mathcal{E}_{0})\simeq\Lambda_{0}$ by construction. All determinant increments descend faithfully, so the number of filtration steps is preserved:
$h_{\Lambda_{0}}(\mathcal{E})\geq h_{\Lambda_{0,Y}}(\mathcal{E}_{Y}).$
Conversely, pulling back $\mathfrak{F}_{X}$ to $Y$ via $f^{*}$ reproduces $\mathfrak{F}_{Y}$, hence $h_{\Lambda_{0,Y}}(\mathcal{E}_{Y})\geq h_{\Lambda_{0}}(\mathcal{E})$. Therefore
$h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,Y}}(\mathcal{E}_{Y}),$
as required. ∎
5.5. Hierarchical depth under MMP
Let $X$ be a smooth projective surface. Consider a sequence of birational maps
$X=X_{0}\dashrightarrow X_{1}\dashrightarrow\cdots\dashrightarrow X_{t}=X_{\min},$
obtained by running the minimal model program (MMP), where $X_{\min}$ is a minimal model of $X$. Each step of the MMP can be realized as a birational morphism
$f_{i}:X_{i}\longrightarrow X_{i+1},\qquad i=0,\dots,t-1,$
where $f_{i}$ is either
  • a contraction of a (1)(-1)-curve $E_{i}\subset X_{i}$ to a smooth point of $X_{i+1}$, or
  • the inverse operation (if one considers a blowup) would be $X_{i+1}\longrightarrow X_{i}$, introducing an exceptional (1)(-1)-curve on the blowup.
By composing these morphisms one obtains a birational map
$\phi:X=X_{0}\dashrightarrow X_{t}=X_{\min},$
which is an isomorphism outside the union of the exceptional curves of the contractions.
Let $\mathcal{E}$ be a vector bundle on $X$ and $\Lambda_{0}\in\mathrm{Pic}(X)$ a reference line bundle. For each contraction $f_{i}:X_{i}\to X_{i+1}$, the pullback $f_{i}^{*}\mathcal{E}_{i+1}$ to $X_{i}$ satisfies
$h_{f_{i}^{*}\Lambda_{i+1}}(f_{i}^{*}\mathcal{E}_{i+1})\geq h_{\Lambda_{i+1}}(\mathcal{E}_{i+1}),$
with equality unless the hierarchical filtration involves additional elementary transforms along the exceptional curve $E_{i}$. Iterating this through the entire MMP sequence gives
(4) $h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})\leq h_{\Lambda_{0}}(\mathcal{E}),$
where $\mathcal{E}_{\min}$ is the pushforward (or the vector bundle descended via saturation) on $X_{\min}$ and $\Lambda_{0,\min}$ is the corresponding reference line bundle. Hence, the maximal hierarchical depth can only decrease (or remain the same) when passing from $X$ to its minimal model.
Theorem 5.6.
Let $f:X\to X_{\min}$ be the birational contraction of a smooth projective surface $X$ onto its minimal model $X_{\min}$ obtained by contracting finitely many (1)(-1)–curves $E_{1},\dots,E_{m}\subset X$. Let $\mathcal{E}$ be a vector bundle on $X$ and let $\mathcal{E}_{\min}$ denote the (saturated) direct image of $\mathcal{E}$ on $X_{\min}$; fix reference line bundles
$\Lambda_{0}\in\mathrm{Pic}(X),\qquad\Lambda_{0,\min}\in\mathrm{Pic}(X_{\min})$
related by
$\Lambda_{0}\simeq f^{*}\Lambda_{0,\min}\otimes\mathcal{O}_{X}\Big(\sum_{j=1}^{m}\beta_{j}E_{j}\Big)$
for unique integers $\beta_{j}\in\mathbb{Z}$. Assume the determinants satisfy
$\det(\mathcal{E})\ \simeq\ f^{*}\det(\mathcal{E}_{\min})\otimes\mathcal{O}_{X}\Big(\sum_{j=1}^{m}\alpha_{j}E_{j}\Big)$
for unique integers $\alpha_{j}\in\mathbb{Z}$ (so that the class $c_{1}(\det\mathcal{E})-f^{*}c_{1}(\det\mathcal{E}_{\min})=\sum_{j}\alpha_{j}[E_{j}]$ in $\mathrm{NS}(X)$). Suppose moreover that the difference divisors appearing in the normalizations and determinants are effective on $X$, i.e. $\alpha_{j}-\beta_{j}\geq 0$ for every $j$.
Then hierarchical filtrations exist for the normalized bundles and the hierarchical depths are related by the exact formula
(5) $h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})+\sum_{j=1}^{m}(\alpha_{j}-\beta_{j}).$
In particular the increase in depth from the minimal model to $X$ is exactly the total multiplicity of the exceptional divisors appearing in the determinant of $\mathcal{E}$ after accounting for the chosen normalizations.
Proof.
By hypothesis
$\det(\mathcal{E})\otimes\Lambda_{0}^{-1}\ \simeq\ f^{*}\big(\det(\mathcal{E}_{\min})\otimes\Lambda_{0,\min}^{-1}\big)\otimes\mathcal{O}_{X}\Big(\sum_{j=1}^{m}(\alpha_{j}-\beta_{j})E_{j}\Big).$
Denote
$\Delta_{\min}:=\det(\mathcal{E}_{\min})\otimes\Lambda_{0,\min}^{-1}\in\mathrm{Pic}(X_{\min}),\qquad\Delta:=\det(\mathcal{E})\otimes\Lambda_{0}^{-1}\in\mathrm{Pic}(X).$
By the usual correspondence between hierarchical filtrations and decompositions of the determinant into effective Cartier increments (Definitions 2.1 and 2.2), the maximal length $h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})$ equals the maximal number of nonzero effective Cartier summands into which any effective representative of $\Delta_{\min}$ may be decomposed; likewise $h_{\Lambda_{0}}(\mathcal{E})$ equals the maximal number of nonzero effective Cartier summands in any effective representative of $\Delta$.
Because pullback preserves effectiveness and the exceptional curves $E_{j}$ are pairwise distinct irreducible Cartier divisors on $X$, any effective decomposition of $\Delta_{\min}$ pulls back to an effective decomposition of $f^{*}\Delta_{\min}$ with the same number of summands, and the factor $\mathcal{O}_{X}\big(\sum_{j}(\alpha_{j}-\beta_{j})E_{j}\big)$ contributes exactly $\sum_{j}(\alpha_{j}-\beta_{j})$ further irreducible effective summands (all supported on the exceptional locus) when one decomposes $\Delta$ into irreducible Cartier divisors. Conversely, every effective decomposition of $\Delta$ decomposes uniquely as the pullback of an effective decomposition of $\Delta_{\min}$ together with the exceptional part (since the $E_{j}$ generate the kernel of $f^{*}:\mathrm{Pic}(X_{\min})\to\mathrm{Pic}(X)$).
Therefore the maximal possible number of irreducible effective summands in $\Delta$ equals the maximal number for $\Delta_{\min}$ plus the sum of the coefficients of the exceptional part, i.e.
$h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})+\sum_{j=1}^{m}(\alpha_{j}-\beta_{j}).$
This proves (5). ∎
Example 5.7.
Let $X_{\min}=\mathbb{P}^{2}$ with hyperplane class $H$, and let $f:X\to X_{\min}$ be the blowup of $\mathbb{P}^{2}$ at two distinct points $p_{1},p_{2}$. Denote by $E_{1},E_{2}\subset X$ the corresponding exceptional divisors. Then $\mathrm{Pic}(X)=\mathbb{Z}\langle f^{*}H\rangle\oplus\mathbb{Z}\langle E_{1}\rangle\oplus\mathbb{Z}\langle E_{2}\rangle$.
Fix the following data:
$\mathcal{E}_{\min}=\mathcal{O}_{\mathbb{P}^{2}}(3)\oplus\mathcal{O}_{\mathbb{P}^{2}}(2)\qquad(\text{rank }2),\qquad\Lambda_{0,\min}=\mathcal{O}_{\mathbb{P}^{2}}.$
Then $\det(\mathcal{E}_{\min})\cong\mathcal{O}_{\mathbb{P}^{2}}(5)$, and by Proposition 4.1 we have
$h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})=5.$
On $X$ consider the vector bundle
$\mathcal{E}=f^{*}\mathcal{E}_{\min}\otimes\mathcal{O}_{X}(E_{1}+2E_{2}).$
Since $\mathrm{rank}(\mathcal{E})=2$ we compute the determinant on $X$ as
$\det(\mathcal{E})\simeq f^{*}\det(\mathcal{E}_{\min})\otimes\mathcal{O}_{X}\big(2(E_{1}+2E_{2})\big)=f^{*}\mathcal{O}_{\mathbb{P}^{2}}(5)\otimes\mathcal{O}_{X}(2E_{1}+4E_{2}).$
Take the normalization on $X$ to be $\Lambda_{0}=\mathcal{O}_{X}(E_{2})$; equivalently $\Lambda_{0}\simeq f^{*}\Lambda_{0,\min}\otimes\mathcal{O}_{X}(0\cdot E_{1}+1\cdot E_{2})$. Thus in the notation of Theorem 5.6 we have the integer coefficients
$\alpha_{1}=2,\ \alpha_{2}=4,\qquad\beta_{1}=0,\ \beta_{2}=1,$
and indeed $\alpha_{j}-\beta_{j}\geq 0$ for $j=1,2$.
We now compute both sides of formula (5). On the minimal model we already have $h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})=5$. On $X$ the class
$\Delta=\det(\mathcal{E})\otimes\Lambda_{0}^{-1}\simeq f^{*}\mathcal{O}_{\mathbb{P}^{2}}(5)\otimes\mathcal{O}_{X}(2E_{1}+3E_{2})$
is represented by an effective divisor of class $5f^{*}H+2E_{1}+3E_{2}$. One may decompose a representative of this class as the sum of
  • five irreducible strict transforms of lines in $\mathbb{P}^{2}$ (each in the class $f^{*}H$), and
  • two copies of $E_{1}$ and three copies of $E_{2}$.
Thus the maximal possible number of nonzero irreducible effective summands in $\Delta$ equals
$5+2+3=10.$
By the definition of hierarchical depth (maximal number of effective divisorial increments for $\det(\mathcal{E})\otimes\Lambda_{0}^{-1}$) this shows $h_{\Lambda_{0}}(\mathcal{E})=10$.
Comparing with Theorem 5.6 we have
$h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})+\sum_{j=1}^{2}(\alpha_{j}-\beta_{j})=5+(2-0)+(4-1)=10,$
in perfect agreement with the direct decomposition above.
6. Interaction with Stability and Moduli Theory
The hierarchical filtration formalism interacts naturally with classical and modern stability theories for vector bundles. Since hierarchical increments are governed by effective divisors, they admit a slope interpretation with respect to a fixed polarization, yielding a discrete refinement of the usual Harder–Narasimhan filtration. Moreover, the birational invariance results of Section 5 ensure that such stability behavior is preserved throughout the steps of the Minimal Model Program.
6.1. Slope behavior under hierarchical increments
Let $X$ be a smooth projective variety of dimension $n\geq 1$ with ample divisor $H$. For a torsion–free sheaf $\mathcal{E}$ on $X$, recall that its slope with respect to $H$ is
$\mu_{H}(\mathcal{E})=\frac{c_{1}(\mathcal{E})\cdot H^{n-1}}{\mathrm{rank}(\mathcal{E})}.$
Given a hierarchical filtration
$\mathfrak{F}:\quad 0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E},\qquad\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i}),$
we may associate to each increment the numerical slope difference
(6) $\Delta_{i}(\mathcal{E},H)=\mu_{H}(\mathcal{E}_{i})-\mu_{H}(\mathcal{E}_{i-1}).$
Theorem 6.1.
Let $X$ be a smooth projective variety of dimension $n\leq 2$ over an algebraically closed field, and let $H$ be an ample divisor on $X$. Let
$\mathfrak{F}:\quad 0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
be a hierarchical filtration of a vector bundle $\mathcal{E}$ of rank $r$, with determinant increments
$\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i}),\qquad D_{i}\geq 0.$
Then
$\mu_{H}(\mathcal{E}_{0})\leq\mu_{H}(\mathcal{E}_{1})\leq\cdots\leq\mu_{H}(\mathcal{E}_{h}).$
and for each index $i$ the inequality $\mu_{H}(\mathcal{E}_{i})>\mu_{H}(\mathcal{E}_{i-1})$ holds precisely when the corresponding determinant increment $D_{i}$ is nonzero ($D_{i}\not\sim 0$).
Proof.
From the determinant relation $\det(\mathcal{E}_{i})=\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i})$ we have $c_{1}(\mathcal{E}_{i})=c_{1}(\mathcal{E}_{i-1})+[D_{i}]$. Since the rank $r$ is constant,
(7) $\mu_{H}(\mathcal{E}_{i})-\mu_{H}(\mathcal{E}_{i-1})=\frac{(c_{1}(\mathcal{E}_{i})-c_{1}(\mathcal{E}_{i-1}))\cdot H^{n-1}}{r}=\frac{c_{1}(D_{i})\cdot H^{n-1}}{r}.$
Each $D_{i}$ is effective, so the intersection number $c_{1}(D_{i})\cdot H^{n-1}$ is nonnegative, with equality if and only if $D_{i}\sim 0$. The stated monotonicity of the slopes follows immediately. ∎
6.2. Birational stability under the Minimal Model Program
The birational invariance results of Section 5 imply that hierarchical depth behaves predictably under the birational operations of the surface MMP. In particular, slope–compatibility is preserved under pullbacks, and semistability is unaffected outside exceptional loci.
Proposition 6.2.
Let $f:Y\to X$ be a birational morphism between smooth projective surfaces, and let $H$ be an ample divisor on $X$. Then for every vector bundle $\mathcal{E}$ on $X$ and its pullback $f^{*}\mathcal{E}$ on $Y$, the divisorial slope differences satisfy
$\Delta_{i}(f^{*}\mathcal{E},f^{*}H)=\Delta_{i}(\mathcal{E},H),$
and the divisorial slope filtration of $\mathcal{E}$ pulls back to that of $f^{*}\mathcal{E}$.
Proof.
Since $f^{*}H^{n-1}=H^{n-1}$ numerically and $f^{*}c_{1}(D_{i})=c_{1}(f^{*}D_{i})$, formula (7) gives $\Delta_{i}(f^{*}\mathcal{E},f^{*}H)=\Delta_{i}(\mathcal{E},H)$. The filtration structure is preserved by Proposition 5.1. ∎
Therefore, the stability behavior of hierarchical filtrations remains invariant throughout the birational steps of the MMP, except when new effective increments supported on exceptional curves appear. In that case, the newly introduced increments correspond to additional positive slopes localized on the exceptional locus, and their total contribution is governed by Theorem 5.6.
Corollary 6.3.
Let $f:Y\to X$ be a birational morphism between smooth projective surfaces obtained by contracting a finite collection of (1)(-1)–curves, and let $H$ be an ample divisor on $X$. Put $H_{Y}:=f^{*}H$.
Let $\mathcal{E}$ be a vector bundle on $X$. If $\mathcal{E}$ is divisorially slope-semistable (respectively, divisorially slope-stable) with respect to $H$, then the pullback bundle $f^{*}\mathcal{E}$ is divisorially slope–semistable (respectively, divisorially slope–stable) with respect to $H_{Y}$.
Proof.
We treat the semistable case; the stable case is identical with strict inequalities.
Let $r=\mathrm{rank}(\mathcal{E})$. Suppose $\mathcal{E}$ is divisorially slope–semistable on $X$ in meaning of [7], i.e. for every saturated nonzero proper subsheaf $F\subset\mathcal{E}$ one has
$\mu_{H}(F)\leq\mu_{H}(\mathcal{E}).$
We must show that every saturated nonzero proper subsheaf $G\subset f^{*}\mathcal{E}$ satisfies
$\mu_{H_{Y}}(G)\leq\mu_{H_{Y}}(f^{*}\mathcal{E})=\mu_{H}(\mathcal{E}).$
Consider any saturated subsheaf $G\subset f^{*}\mathcal{E}$. Push it forward and take saturation on $X$:
$F:=\big(f_{*}G\big)^{\mathrm{sat}}\subset\mathcal{E}.$
(Here saturation means reflexive hull / saturation in the sense of subsheaves.) By construction $F$ is a torsion–free subsheaf of $\mathcal{E}$ of rank $s:=\mathrm{rank}(F)$ with $0<s\leq r$.
There is an equality of determinant classes on $Y$ of the form
$\det(G)\simeq(f^{*}\det F)\otimes\mathcal{O}_{Y}\Big(\sum_{j}a_{j}E_{j}\Big),\qquad a_{j}\in\mathbb{Z}_{\geq 0},$
where the $E_{j}$ are the exceptional curves of $f$ and the integers $a_{j}$ record the possible exceptional components appearing in $\det(G)$. (Indeed every line bundle on $Y$ decomposes uniquely as the pullback of a line bundle on $X$ tensor a line bundle supported on exceptional curves.)
Now compute slopes with respect to the pulled back polarization $H_{Y}=f^{*}H$. Since intersection is functorial under pullback and every exceptional curve is contracted to a point, we have for every $j$
$E_{j}\cdot H_{Y}=E_{j}\cdot f^{*}H=0.$
Consequently the exceptional part does not contribute to the slope:
$\mu_{H_{Y}}(G)=\frac{c_{1}(\det G)\cdot H_{Y}}{\mathrm{rank}(G)}=\frac{\big(c_{1}(f^{*}\det F)+\sum_{j}a_{j}[E_{j}]\big)\cdot f^{*}H}{\mathrm{rank}(G)}=\frac{c_{1}(\det F)\cdot H}{\mathrm{rank}(G)}=\mu_{H}(F).$
But the right-hand side equals $\mu_{H}(F)$ (by definition). Hence
$\mu_{H_{Y}}(G)=\mu_{H}(F).$
Since $\mathcal{E}$ is divisorially slope–semistable, $\mu_{H}(F)\leq\mu_{H}(\mathcal{E})$. Moreover $\mu_{H}(\mathcal{E})=\mu_{H_{Y}}(f^{*}\mathcal{E})$ because $c_{1}(f^{*}\mathcal{E})=f^{*}c_{1}(\mathcal{E})$ and intersection with $f^{*}H$ is compatible with pullback. Therefore
$\mu_{H_{Y}}(G)=\mu_{H}(F)\leq\mu_{H}(\mathcal{E})=\mu_{H_{Y}}(f^{*}\mathcal{E}),$
as required. This proves that $f^{*}\mathcal{E}$ is divisorially slope–semistable with respect to $H_{Y}$.
The same argument with strict inequalities shows that divisorial slope–stability is preserved: if $\mathcal{E}$ is stable and $G\subset f^{*}\mathcal{E}$ is a proper saturated subsheaf, then the strict inequality $\mu_{H}(F)<\mu_{H}(\mathcal{E})$ holds, and hence $\mu_{H_{Y}}(G)<\mu_{H_{Y}}(f^{*}\mathcal{E})$. ∎
6.3. Comparison with Harder–Narasimhan filtrations
Let $\mathcal{E}$ be a torsion–free sheaf on $(X,H)$ with Harder–Narasimhan filtration
$\mathcal{E}^{(0)}\subset\mathcal{E}^{(1)}\subset\cdots\subset\mathcal{E}^{(s)}=\mathcal{E},$
whose successive quotients $\mathrm{Gr}_{i}^{\mathrm{HN}}(\mathcal{E})=\mathcal{E}^{(i)}/\mathcal{E}^{(i-1)}$ are semistable with decreasing slopes $\mu_{1}>\mu_{2}>\cdots>\mu_{s}$. Then the determinant relation gives
$c_{1}(\det\mathcal{E})=\sum_{i=1}^{s}c_{1}\big(\mathrm{Gr}_{i}^{\mathrm{HN}}(\mathcal{E})\big),$
so that each Harder–Narasimhan quotient contributes an effective class whenever $\mathcal{E}$ is globally generated or arises as a positive extension.
Split bundle on a curve
Let $C$ be a smooth projective curve over a field and set
$\mathcal{E}=\bigoplus_{i=1}^{r}\mathcal{O}_{C}(d_{i}),\qquad d_{i}\in\mathbb{Z}.$
Write $D:=\sum_{i=1}^{r}d_{i}=\deg(\det\mathcal{E})$. We consider two invariants:
With notation as above we obtain the hierarchical depth and the length of Harder-Narasimhan filtration of $\mathcal{E}$.
(1) Hierarchical depth.
It was shown in Proposition 4.1 that
$h_{C}(\mathcal{E})=\sum_{i=1}^{r}\max\{d_{i},0\}.$ In particular, if all $d_{i}\geq 0$ then $h_{\mathcal{O}_{C}}(\mathcal{E})=\sum_{i}d_{i}=D$. If some $d_{i}<0$ the corresponding summand contributes 0 to the depth.
(2) Harder–Narasimhan length.
The Harder-Narasimhan length $s$ equals the number of distinct values among the degrees $\{d_{1},\dots,d_{r}\}$. Equivalently, if we arrange the degrees in nonincreasing order and group equal values,
$\mathcal{E}=\bigoplus_{j=1}^{s}\Big(\bigoplus_{k=1}^{r_{j}}\mathcal{O}_{C}(d^{(j)})\Big),$
where $d^{(1)}>\cdots>d^{(s)}$ are the distinct degrees and $r_{j}$ the multiplicities, then the HN filtration has length $s$ with graded pieces $\mathrm{Gr}_{j}^{\mathrm{HN}}\cong\mathcal{O}_{C}(d^{(j)})^{\oplus r_{j}}$. In the following we imply to this fact.
Let $C$ be a smooth projective curve over an algebraically closed field, and let
$\mathcal{E}=\bigoplus_{i=1}^{r}\mathcal{O}_{C}(d_{i})$
be a direct sum of line bundles of degrees $d_{1},\dots,d_{r}\in\mathbb{Z}$. Then the Harder–Narasimhan filtration of $\mathcal{E}$ (with respect to any polarization) is obtained by ordering the summands in nonincreasing degree:
$\mathcal{E}^{(0)}=0\subset\mathcal{E}^{(1)}=\bigoplus_{d_{i}=d_{(1)}}\mathcal{O}_{C}(d_{i})\subset\mathcal{E}^{(2)}=\bigoplus_{d_{i}\geq d_{(2)}}\mathcal{O}_{C}(d_{i})\subset\cdots\subset\mathcal{E}^{(s)}=\mathcal{E},$
where $d_{(1)}>d_{(2)}>\cdots>d_{(s)}$ are the distinct degrees among the $d_{i}$. In particular,
[…]
The existence and uniqueness of the Harder–Narasimhan filtration for torsion-free sheaves on smooth projective varieties was established by Harder and Narasimhan [7].
On a smooth projective curve $C$, every line bundle $\mathcal{O}_{C}(d)$ is stable with respect to any polarization, and its slope is simply
$\mu(\mathcal{O}_{C}(d))=\frac{\deg\mathcal{O}_{C}(d)}{\mathrm{rank}\mathcal{O}_{C}(d)}=d.$
Hence each direct summand $\mathcal{O}_{C}(d_{i})$ is a semistable bundle of slope $d_{i}$. Let $d_{(1)}>\cdots>d_{(s)}$ be the distinct degrees among the $d_{i}$, and define
$\mathcal{E}^{(k)}:=\bigoplus_{d_{i}\geq d_{(k)}}\mathcal{O}_{C}(d_{i}).$
Then the successive quotients
$\mathrm{Gr}_{k}^{\mathrm{HN}}(\mathcal{E}):=\mathcal{E}^{(k)}/\mathcal{E}^{(k-1)}=\bigoplus_{d_{i}=d_{(k)}}\mathcal{O}_{C}(d_{i})$
are semistable bundles of slope $\mu_{k}=d_{(k)}$, and the sequence $\mu_{1}>\mu_{2}>\cdots>\mu_{s}$ is strictly decreasing. By the uniqueness of the Harder-Narasimhan filtration this is precisely the HN filtration of $\mathcal{E}$.
Example 6.4.
  1. (i)
    Take $r=3$ and $\mathcal{E}=\mathcal{O}_{C}(3)\oplus\mathcal{O}_{C}(1)\oplus\mathcal{O}_{C}(0)$. Then $h_{\mathcal{O}_{C}}(\mathcal{E})=3+1+0=4$. The HN degrees are $3>1>0$, so $s=3$.
  2. (ii)
    Take $r=3$ and $\mathcal{E}=\mathcal{O}_{C}(1)\oplus\mathcal{O}_{C}(1)\oplus\mathcal{O}_{C}(0)$. Then $h_{\mathcal{O}_{C}}(\mathcal{E})=1+1+0=2$. The HN degrees are $1>0$, so $s=2$.
  3. (iii)
    Take $r=3$ and $\mathcal{E}=\mathcal{O}_{C}(1)\oplus\mathcal{O}_{C}(0)\oplus\mathcal{O}_{C}(0)$. Then $h_{\mathcal{O}_{C}}(\mathcal{E})=1$ and the HN filtration has length $s=2$.
It follows that there is no relation between hierarchical depth of a vector bundle and the length of its HN filtration.
7. Applications to algebraic–geometric codes
In the study of constructing algebraic geometry (AG) codes on varieties, it is standard to assume that the underlying variety is smooth and irreducible, since the definition of such codes relies on the well-behaved structure of the function field and divisor theory on smooth varieties. However, within the framework of the Minimal Model Program (MMP), the situation becomes more delicate: the birational operations used in the MMP—such as divisorial contractions and flips—may introduce singularities, even if one starts with a smooth variety. In particular, in dimensions three and higher, the MMP involves both divisorial contractions (which contract subvarieties of codimension one) and flips (which replace certain singular loci by others), and these operations generally lead to varieties with terminal or canonical singularities. Consequently, to maintain smoothness throughout the process, one must restrict attention to varieties of dimension at most two. For smooth projective surfaces, the MMP simplifies considerably: there are no flips, and the only possible operations are contractions of (1)(-1)-curves. Each such contraction is the inverse of a blowup of a smooth point and therefore preserves smoothness. Hence, in the case of smooth surfaces, the MMP coincides with the classical Castelnuovo contraction process, ensuring that the minimal model obtained remains smooth. This restriction to smooth varieties (and thus to surfaces) is essential for the consistent application of AG code constructions on varieties governed by the MMP framework.
7.1. AG codes from vector bundles
Let $X$ be a smooth projective variety over $\mathbb{F}_{q}$ and let $\mathcal{E}$ be a vector bundle of rank $r$ on $X$. Let
$D=\{P_{1},\dots,P_{N}\}\subset X(\mathbb{F}_{q})$
be a finite ordered set of $N$ distinct $\mathbb{F}_{q}$–rational points such that evaluation of sections of $\mathcal{E}$ is well defined at every $P_{i}$. Fix, for each $P_{i}$, an $\mathbb{F}_{q}$–linear identification (trivialization) of the fiber $\mathcal{E}|_{P_{i}}\cong\mathbb{F}_{q}^{\,r}$.
The algebraic–geometric code associated to $(\mathcal{E},D)$ is defined by
$C_{X}(\mathcal{E},D):=\operatorname{ev}_{D}\big(H^{0}(X,\mathcal{E})\big)=\{(s(P_{1}),\dots,s(P_{N})):s\in H^{0}(X,\mathcal{E})\}\subseteq\mathbb{F}_{q}^{\,rN}.$
For details of AG codes from vector bundles see [11, 16].
7.2. Minimal models and coding–theoretic optimization
Let $X$ be a smooth projective surface defined over $\mathbb{F}_{q}$, and let $f:Y\to X$ denote a blow–up of $X$ at a smooth $\mathbb{F}_{q}$–rational point, with exceptional divisor $E\subset Y$. As discussed in Proposition 5.4, such a birational modification increases the hierarchical depth of vector bundles by an integer $\delta\geq 0$, reflecting the presence of additional effective increments supported on $E$.
From the coding–theoretic viewpoint, these exceptional divisors have unfavourable effects on the parameters of algebraic–geometric codes constructed from vector bundles. In order to optimize the geometric data used for encoding, it is therefore natural to pass to the minimal model of $X$, obtained by contracting all (1)(-1)–curves.
Proposition 7.1.
Let $f:Y\to X$ be the contraction of a (1)(-1)–curve $E\subset Y$ onto a smooth $\mathbb{F}_{q}$–rational point $P_{0}\in X$. Let $\mathcal{E}_{Y}$ be a vector bundle on $Y$ whose hierarchical filtration has no effective increment supported on $E$, and set $\mathcal{E}:=f_{*}\mathcal{E}_{Y}$ (after torsion saturation). Then there is a canonical identification
$H^{0}(Y,\mathcal{E}_{Y})\cong H^{0}(X,\mathcal{E}),$
and for any evaluation set $D\subset X(\mathbb{F}_{q})\setminus\{P_{0}\}$ one has a commutative diagram
[…]
Consequently,
$C_{X}(\mathcal{E},D)=C_{Y}(\mathcal{E}_{Y},f^{-1}(D)).$
Proof.
Since $f$ is a birational morphism contracting a smooth rational curve with normal bundle $\mathcal{O}_{E}(-1)$, we have $R^{i}f_{*}\mathcal{O}_{Y}=0$ for $i>0$. If no increment of the hierarchical filtration of $\mathcal{E}_{Y}$ is supported on $E$, then $\mathcal{E}_{Y}$ is numerically trivial along $E$, and thus $R^{i}f_{*}\mathcal{E}_{Y}=0$ for $i>0$. By the projection formula,
$H^{0}(Y,\mathcal{E}_{Y})\cong H^{0}(X,f_{*}\mathcal{E}_{Y})=H^{0}(X,\mathcal{E}),$
which yields the claimed isomorphism of global sections. For any evaluation divisor $D\subset X(\mathbb{F}_{q})\setminus\{P_{0}\}$, the points $f^{-1}(D)$ correspond bijectively to $D$, and the fibers $\mathcal{E}|_{P_{i}}\cong\mathcal{E}_{Y}|_{f^{-1}(P_{i})}$ are canonically identified. Hence the evaluation maps coincide, giving the commutative diagram and equality of codes. ∎
Corollary 7.2.
Let $X_{\min}$ be the minimal model of a smooth projective surface $X$ over $\mathbb{F}_{q}$, and let $f:X\to X_{\min}$ denote the composition of its (1)(-1)–curve contractions. For every vector bundle $\mathcal{E}$ on $X$ whose hierarchical filtration has no effective increment supported on the exceptional locus of $f$, one has
$C_{X}(\mathcal{E},D)=C_{X_{\min}}(f_{*}\mathcal{E},\,f(D)).$
In particular, all essential code parameters (length, dimension, and minimum distance) are preserved under the MMP reduction to the minimal model.
In the following we show that MMP can improve code parameters.
Proposition 7.3.
Let $f:Y\to X$ be the contraction of a set of (1)(-1)-curves $E_{1},\dots,E_{m}$ on a smooth projective surface $Y$, and let $\mathcal{E}_{Y}$ be a vector bundle of rank $r$ on $Y$ with $H^{0}(Y,\mathcal{E}_{Y})$ of dimension $k$. Set $\mathcal{E}:=f_{*}\mathcal{E}_{Y}$.
Let $D_{Y}\subset Y(\mathbb{F}_{q})$ be a set of $N_{Y}$ evaluation points, and let $D_{E}=D_{Y}\cap\bigcup_{i=1}^{m}E_{i}$ be the subset of points lying on the exceptional divisors, with $D_{E}=N_{E}$. Let $D=f(D_{Y}\setminus D_{E})\subset X$.
Define the AG codes:
$C_{Y}:=\operatorname{ev}_{D_{Y}}(H^{0}(Y,\mathcal{E}_{Y}))\subset(\mathbb{F}_{q}^{r})^{N_{Y}},\quad C_{X}:=\operatorname{ev}_{D}(H^{0}(X,\mathcal{E}))\subset(\mathbb{F}_{q}^{r})^{N_{Y}-N_{E}}.$
Then:
  1. (i)
    The dimension is preserved under contraction:
    $\dim C_{X}=\dim C_{Y}=k.$
  2. (ii)
    The absolute minimum distance is also preserved:
    $d_{\min}(C_{X})=d_{\min}(C_{Y})=:d_{\min}.$
  3. (iii)
    The normalized minimum distance improves according to
    $\delta_{Y}:=\frac{d_{\min}}{r\,N_{Y}}\leq\delta_{X}:=\frac{d_{\min}}{r\,(N_{Y}-N_{E})}.$
  4. (iv)
    In particular, the improvement factor of relative distance is
    $\frac{\delta_{X}}{\delta_{Y}}=\frac{N_{Y}}{N_{Y}-N_{E}}>1,$
    which grows with the proportion of evaluation points lying on exceptional divisors.
Proof.
We adopt the notation of the proposition. Let
$f:Y\to X$
be the contraction of a set of (1)(-1)-curves $E_{1},\dots,E_{m}\subset Y$, and let $\mathcal{E}_{Y}$ be a vector bundle of rank $r$ on $Y$. Let $D_{Y}\subset Y(\mathbb{F}_{q})$ be the evaluation set, with $D_{E}=D_{Y}\cap\bigcup_{i}E_{i}$ the points on exceptional divisors, $D_{E}=N_{E}$, and $D=f(D_{Y}\setminus D_{E})$ the evaluation set on the minimal model $X$, $D=N_{Y}-N_{E}$.
Step 1:
By Proposition 5.5, the global sections of $\mathcal{E}_{Y}$ descend via $f_{*}$ to global sections of the bundle $\mathcal{E}=f_{*}\mathcal{E}_{Y}$ on $X$:
$H^{0}(X,\mathcal{E})\cong H^{0}(Y,\mathcal{E}_{Y}).$
Hence the code dimensions are identical:
$\dim C_{X}=\dim\operatorname{ev}_{D}(H^{0}(X,\mathcal{E}))=\dim H^{0}(X,\mathcal{E})=\dim H^{0}(Y,\mathcal{E}_{Y})=\dim C_{Y}=k.$
Step 2:
Let $d_{\min}(C_{Y})$ denote the minimal Hamming weight of a nonzero codeword of $C_{Y}$. Every section of $\mathcal{E}_{Y}$ vanishes along the exceptional divisors $E_{i}$ by the twisting or by the kernel of the elementary transform along $E_{i}$. Consequently, for all $s\in H^{0}(Y,\mathcal{E}_{Y})$, the coordinates corresponding to $D_{E}$ are identically zero.
The restriction of $s$ to $D^{\prime}:=D_{Y}\setminus D_{E}$ defines a codeword in $C_{X}$. Conversely, any nonzero codeword in $C_{X}$ lifts to a codeword in $C_{Y}$ with zeros in the $D_{E}$ coordinates. Hence the minimum weight satisfies
$d_{\min}(C_{X})=d_{\min}(C_{Y})=:d_{\min}.$
Step 3:
By definition, the normalized minimum distance is the ratio of absolute minimum distance to total number of coordinates:
$\delta_{Y}:=\frac{d_{\min}}{r\,N_{Y}},\qquad\delta_{X}:=\frac{d_{\min}}{r\,(N_{Y}-N_{E})}.$
Since $N_{E}\geq 0$, we have $N_{Y}-N_{E}\leq N_{Y}$, so
$\delta_{X}=\frac{d_{\min}}{r(N_{Y}-N_{E})}\geq\frac{d_{\min}}{rN_{Y}}=\delta_{Y}.$
Step 4:
The relative improvement in normalized distance is
$\frac{\delta_{X}}{\delta_{Y}}=\frac{d_{\min}/[r(N_{Y}-N_{E})]}{d_{\min}/(rN_{Y})}=\frac{N_{Y}}{N_{Y}-N_{E}}>1,$
as long as $N_{E}>0$, and it grows with the proportion of evaluation points lying on exceptional divisors.
All statements (i)–(iv) of the proposition follow directly from the descent of global sections via $f_{*}$, the structure of sections vanishing on exceptional divisors, and the definitions of absolute and normalized minimum distance. This completes the proof. ∎
Example 7.4.
Let $X=\mathbb{P}^{2}_{\mathbb{F}_{5}}$ be the projective plane over $\mathbb{F}_{5}$, and let $f:Y\to X$ be the blow–up of $X$ at the $\mathbb{F}_{5}$–rational point $p=[1:0:0]$, with exceptional divisor $E\subset Y$.
Consider the line bundle on $Y$
$\mathcal{E}_{Y}:=f^{*}\mathcal{O}_{X}(3)\otimes\mathcal{O}_{Y}(-E)=\mathcal{O}_{Y}(3H-E),$
where $H=f^{*}(\text{line})$ is the pullback of a line in $\mathbb{P}^{2}$. Then $H^{0}(Y,\mathcal{E}_{Y})$ has dimension
$h^{0}(Y,\mathcal{E}_{Y})=h^{0}(\mathbb{P}^{2},\mathcal{O}(3))-1=10-1=9,$
because every cubic passing through $p$ defines a section, and subtracting $E$ forces vanishing at the blown-up point.
Let $D_{Y}$ be a set of $N_{Y}=12$ $\mathbb{F}_{5}$–rational points on $Y$, decomposed as
$D_{Y}=D^{\prime}\cup D_{E},\qquad\#D^{\prime}=10,\ \#D_{E}=2,\quad D_{E}\subset E(\mathbb{F}_{5}),$
so that two evaluation points lie on the exceptional curve $E$.
Code on the blown-up surface.
The AG code on $Y$ is
$C_{Y}(\mathcal{E}_{Y},D_{Y})=\operatorname{ev}_{D_{Y}}(H^{0}(Y,\mathcal{E}_{Y}))\subset\mathbb{F}_{5}^{\,rN_{Y}},\quad r=1.$
By construction, all sections vanish on $E$, so the coordinates corresponding to $D_{E}$ are identically zero for every codeword. Hence
$C_{Y}(\mathcal{E}_{Y},D_{Y})\cong\{0\}^{\,2}\times C_{Y}(\mathcal{E}_{Y},D^{\prime}),$
with $\dim C_{Y}(\mathcal{E}_{Y},D_{Y})=\dim C_{Y}(\mathcal{E}_{Y},D^{\prime})=9$ and minimum distance
$d_{\min}(C_{Y}(\mathcal{E}_{Y},D_{Y}))=d_{\min}(C_{Y}(\mathcal{E}_{Y},D^{\prime}))=4.$
The normalized (relative) minimum distance is
$\delta_{Y}=\frac{d_{\min}}{N_{Y}}=\frac{4}{12}=\frac{1}{3}.$
Code on the minimal model.
On the contracted surface $X=\mathbb{P}^{2}$, the direct image bundle $\mathcal{E}:=f_{*}\mathcal{E}_{Y}=\mathcal{O}_{X}(3)$ has
[…]
Let $D:=f(D^{\prime})\subset X(\mathbb{F}_{5})$, with #D=10\#D=10. The corresponding code is
$C_{X}(\mathcal{E},D)=\operatorname{ev}_{D}(H^{0}(X,\mathcal{E}))\subset\mathbb{F}_{5}^{10}.$
The minimum distance remains $d_{\min}(C_{X}(\mathcal{E},D))=4$, and the dimension is still $9$. However, the normalized minimum distance is
$\delta_{X}=\frac{d_{\min}}{N}=\frac{4}{10}=0.40>0.33=\delta_{Y}.$
Conclusion.
Contracting the exceptional divisor removes the degenerate coordinates introduced by evaluation points on $E$. While the absolute dimension and absolute minimum distance remain unchanged, the relative minimum distance (distance per code length) improves from $0.33$ to $0.40$.
Example 7.5.
Let $X=\mathbb{P}^{2}_{\mathbb{F}_{7}}$ and let $f:Y\to X$ be the blowup of $X$ at the $\mathbb{F}_{7}$-rational point $p=[1:0:0]$. Let $E\subset Y$ denote the exceptional (1)(-1)-curve, with $E(\mathbb{F}_{7})$ containing 8 rational points.
Consider the rank $r=3$ vector bundle on $Y$:
$\mathcal{E}_{Y}:=\mathcal{O}_{Y}(3H-E)\oplus\mathcal{O}_{Y}(2H-E)\oplus\mathcal{O}_{Y}(H-E),$
where $H=f^{*}(\text{line})$ is the pullback of a line in $\mathbb{P}^{2}$.
Step 1: Global sections.
The global sections of each summand are:
[…]
$H^{0}(Y,\mathcal{O}_{Y}(2H-E))$ $\cong H^{0}(X,\mathcal{O}_{X}(2))-1=6-1=5,$
$H^{0}(Y,\mathcal{O}_{Y}(H-E))$ $\cong H^{0}(X,\mathcal{O}_{X}(1))-1=3-1=2.$
Hence
$\dim_{\mathbb{F}_{7}}H^{0}(Y,\mathcal{E}_{Y})=9+5+2=16.$
Step 2: Evaluation points.
Choose $N_{Y}=18$ $\mathbb{F}_{7}$-rational points on $Y$ for evaluation, decomposed as
$D_{Y}=D^{\prime}\cup D_{E},\qquad\#D_{E}=5,\ \#D^{\prime}=13,\quad D_{E}\subset E(\mathbb{F}_{7}).$
Step 3: AG code on $Y$.
The evaluation map
$\operatorname{ev}_{D_{Y}}:H^{0}(Y,\mathcal{E}_{Y})\longrightarrow(\mathbb{F}_{7}^{3})^{18}$
defines the AG code $C_{Y}(\mathcal{E}_{Y},D_{Y})$.
Every section of $\mathcal{E}_{Y}$ vanishes along $E$ (due to the E-E twist) therefore the coordinates corresponding to $D_{E}$ are all zeros:
$C_{Y}(\mathcal{E}_{Y},D_{Y})\cong\{0\}^{3\cdot 5}\times C_{Y}(\mathcal{E}_{Y},D^{\prime}),\quad\dim C_{Y}(\mathcal{E}_{Y},D_{Y})=16.$
The absolute minimum distance (number of nonzero coordinates) for $C_{Y}(\mathcal{E}_{Y},D_{Y})$ is
$d_{\min}(C_{Y}(\mathcal{E}_{Y},D_{Y}))=d_{\min}(C_{Y}(\mathcal{E}_{Y},D^{\prime}))=4\quad(\text{computed on }D^{\prime}).$
The normalized (relative) minimum distance is
$\delta_{Y}=\frac{d_{\min}}{3\cdot 18}=\frac{4}{54}\approx 0.074.$
Step 4: AG code on the minimal model.
On the contracted surface $X$, the direct image bundle
$\mathcal{E}:=f_{*}\mathcal{E}_{Y}\cong\mathcal{O}_{X}(3)\oplus\mathcal{O}_{X}(2)\oplus\mathcal{O}_{X}(1)$
by Proposition 5.5. Let
$D=f(D^{\prime})\subset X(\mathbb{F}_{7}),\quad\#D=13.$
The corresponding code $C_{X}(\mathcal{E},D)$ is
$C_{X}(\mathcal{E},D)=\operatorname{ev}_{D}(H^{0}(X,\mathcal{E}))\subset(\mathbb{F}_{7}^{3})^{13}.$
Dimension remains $16$, absolute minimum distance remains $d_{\min}(C_{X}(\mathcal{E},D))=4$ and normalized minimum distance improves:
$\delta_{X}=\frac{4}{3\cdot 13}=\frac{4}{39}\approx 0.103>\delta_{Y}\approx 0.074.$
Consequently, contracting the exceptional divisor $E$ removes the 15 degenerate coordinates ($5$ points $\times$ 33 ranks), while keeping the dimension and absolute minimum distance unchanged. As a result, the relative minimum distance increases from $\sim 0.074$ to $\sim 0.103$, demonstrating quantitatively the effect described in Proposition 7.3.
This shows that the MMP operation (contraction of (1)(-1)-curves) can strictly improve the coding-theoretic quality of AG codes when degenerate coordinates are introduced by exceptional divisors.