### v3

# Hierarchical filtrations of vector bundles and birational geometry

Abstract.
We introduce and systematically study hierarchical filtrations of vector bundles on smooth projective varieties. These are filtrations by saturated subsheaves of equal rank whose successive quotients are torsion sheaves supported in codimension one. The associated numerical invariant, called hierarchical depth, measures the maximal length of such filtrations.
We establish general bounds for hierarchical depth in terms of the determinant class and provide exact formulas for smooth curves and varieties of Picard rank one. A key technical result concerns the commutativity of elementary transforms along disjoint divisors and their role in constructing filtrations.
For surfaces, we analyze the behavior of hierarchical depth under birational morphisms and prove that it transforms additively along the steps of the minimal model program. In particular, we obtain an explicit formula relating the depth on a surface to that on its minimal model via exceptional divisor contributions.
As an application, we connect hierarchical depth to degeneracies in algebraic–geometric codes and show that birational simplification via the MMP leads to effective improvements of code parameters. This establishes hierarchical depth as a new bridge between birational geometry, vector bundle theory, and coding theory.
Key words and phrases:
Vector bundles, Hierarchical filtration, MMP, AG codes
2020 Mathematics Subject Classification:
primary; 14G50, Secondary; 14C20, 14H52
Introduction
Filtrations of vector bundles are fundamental tools in algebraic geometry, from Grothendieck’s foundational work on coherent sheaves [4] to the Harder–Narasimhan filtration in stability theory [7]. While stability filtrations depend on a choice of polarization, there remains a need for intrinsic, polarization–free filtrations that reflect divisorial and birational properties.
This paper introduces a new filtration–theoretic framework for torsion–free sheaves adapted precisely to birational geometry. We define hierarchical filtrations—chains of saturated subsheaves of equal rank with torsion quotients supported in codimension one—and their associated numerical invariant, the hierarchical depth. Unlike stability–theoretic invariants, hierarchical depth is independent of any polarization and governed instead by the determinant and the divisor theory of the underlying variety.
Our work builds on and substantially extends the hierarchical filtration theory for line bundles developed in [14], where applications to algebraic–geometric codes were first explored. The present paper develops a comprehensive theory for arbitrary rank, investigates its interaction with birational geometry and the minimal model program, and establishes new connections to coding theory.
Main results
  1. 1.
    We introduce hierarchical filtrations and hierarchical depth (Definition 2.1), prove their finiteness (Proposition 2.4), and give explicit formulas in basic geometric situations (Propositions 4.1, 2.4).
  2. 2.
    We establish commutativity properties for elementary transforms along disjoint divisors (Lemma 3.2, Corollary 3.4), providing a concrete mechanism for constructing filtrations.
  3. 3.
    For surfaces, we prove that hierarchical depth transforms additively under birational morphisms arising in the MMP (Theorem 5.6). In particular, it admits an explicit description on minimal models with variation controlled by exceptional divisors.
  4. 4.
    We relate hierarchical depth to slope stability (Theorem 6.1) and show it is preserved under birational morphisms (Corollary 6.3).
  5. 5.
    As applications to coding theory, we demonstrate that hierarchical depth detects divisorial degeneracies affecting algebraic–geometric code parameters (Proposition 7.1). Birational simplification via the MMP leads to effective improvements in both absolute and normalized minimum distances (Corollary 7.5)
These results situate hierarchical depth as a genuinely birational invariant in the surface case and provide a new computational tool linking the geometry of vector bundles to arithmetic applications.
Acknowledgments
The author is grateful to Professor Pierre Deligne for valuable comments that led to a clarification and correction of the definition of hierarchical filtration.
1. Preliminaries
Conventions. Throughout, $\mathbb{F}$ denotes an algebraically closed field of arbitrary characteristic. All schemes are separated and of finite type over $\mathbb{F}$. A variety is an integral normal projective $\mathbb{F}$-scheme. Unless otherwise stated, varieties are assumed smooth; the restriction to dimension at most $2$ applies only when invoking the minimal model program in Sections 5 and onward.
Divisors and linear equivalence. For a variety $X$, denote by $\mathrm{Div}(X)$ the group of Weil divisors and by $\mathrm{Pic}(X)$ the Picard group. When $X$ is smooth, Cartier and Weil divisors coincide. An effective divisor is written $D\geq 0$. Linear equivalence is denoted by $\sim$.
Sheaves and determinants. All sheaves are coherent $\mathcal{O}_{X}$-modules. A coherent sheaf $\mathcal{E}$ is torsion-free if it has no nonzero subsheaves supported in codimension $\geq 1$. A subsheaf $\mathcal{F}\subseteq\mathcal{E}$ is saturated if $\mathcal{E}/\mathcal{F}$ is torsion-free.
For a torsion-free sheaf $\mathcal{E}$ of rank $r$, its determinant is
$\det(\mathcal{E}):=\bigl(\wedge^{r}\mathcal{E}\bigr)^{\vee\vee},$
identified with the corresponding divisor class. If $\mathcal{F}\hookrightarrow\mathcal{E}$ is an inclusion of torsion-free sheaves of equal rank with torsion quotient, there exists an effective divisor $D$ (the determinantal divisor) such that
$\det(\mathcal{F})\simeq\det(\mathcal{E})(-D).$
Elementary transforms. Let $\mathcal{E}$ be a torsion-free sheaf on $X$ and $D\subset X$ an effective divisor. An elementary transform of $\mathcal{E}$ along $D$ is a sheaf $\mathcal{E}^{\prime}$ fitting into
$0\longrightarrow\mathcal{E}^{\prime}\longrightarrow\mathcal{E}\longrightarrow\mathcal{T}\longrightarrow 0,$
where $\mathcal{T}$ is a torsion sheaf supported on $D$. Such transforms preserve rank and satisfy $\det(\mathcal{E}^{\prime})\simeq\det(\mathcal{E})(-D)$.
Birational morphisms. For a proper birational morphism $f:Y\to X$ between normal varieties:
  • The pullback $f^{*}\mathcal{E}$ of a torsion-free sheaf $\mathcal{E}$ on $X$ may contain torsion; in practice we work with its torsion-free quotient.
  • The pushforward $f_{*}\mathcal{F}$ of a torsion-free sheaf $\mathcal{F}$ on $Y$ is understood to be replaced by its saturation when viewed as a subsheaf on $X$.
These conventions ensure that all sheaves considered remain torsion-free and all inclusions saturated.
Cohomology and slopes. We write $h^{i}(X,\mathcal{F})=\dim_{\mathbb{F}}H^{i}(X,\mathcal{F})$. For a smooth projective variety $X$ with ample divisor $H$, the slope of a torsion-free sheaf $\mathcal{E}$ is
$\mu_{H}(\mathcal{E})=\frac{c_{1}(\det\mathcal{E})\cdot H^{\dim X-1}}{\mathrm{rank}\mathcal{E}}.$
2. Hierarchical filtration of vector bundles
Definition 2.1.
Let $X$ be a smooth projective variety over a field, $\Lambda_{0}\in\mathrm{Pic}(X)$ a reference line bundle (typically $\Lambda_{0}=\mathcal{O}_{X}$), and $\mathcal{E}$ a vector bundle of rank $r$ on $X$. A hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$ is a finite chain
$\mathfrak{F}:\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E},$
where for each $i=1,\dots,h$:
  1. (i)
    $\mathcal{E}_{i}$ is a vector bundle of rank $r$, and $\mathcal{E}_{i-1}\hookrightarrow\mathcal{E}_{i}$ is saturated;
  2. (ii)
    There exists a nonzero effective Cartier divisor $D_{i}$ with $\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i})$;
  3. (iii)
    $\det(\mathcal{E}_{0})\simeq\Lambda_{0}$.
Note that each one of two following statements are equivalent to condition (ii) of above definition:
(ii-1) the torsion sheaf $\mathcal{Q}_{i}$ is supported (set-theoretically) on $D_{i}$, and
$c_{1}(\mathcal{E}_{i})=c_{1}(\mathcal{E}_{i-1})+[D_{i}]\ \in\mathrm{NS}(X).$
(ii-2) locally on $X$ the inclusion $\mathcal{E}_{i-1}\hookrightarrow\mathcal{E}_{i}$ is obtained from $\mathcal{E}_{i-1}$ by an elementary transform (Hecke modification) along the Cartier divisor $D_{i}$; equivalently there is an exact sequence
$0\longrightarrow\mathcal{E}_{i-1}\longrightarrow\mathcal{E}_{i}\longrightarrow\mathcal{T}_{i}\longrightarrow 0$
with $\mathcal{T}_{i}$ a torsion sheaf supported on $D_{i}$ whose determinant (in the Knudsen-Mumford sense, or via the Fitting ideal) is $\mathcal{O}_{X}(D_{i})$.
Definition 2.2.
Let $\mathcal{E}$ and $\Lambda_{0}$ be as above. The hierarchical depth of $\mathcal{E}$ with respect to $\Lambda_{0}$, denoted $h_{\Lambda_{0}}(\mathcal{E})$, is the maximal integer $h\geq 0$ for which there exists a hierarchical filtration of length $h$ normalized by $\Lambda_{0}$. If no such finite filtration exists we set $h_{\Lambda_{0}}(\mathcal{E})=$-$\infty$. We adopt the convention $h_{\Lambda_{0}}(\mathcal{E}_{0})=0$ when $\mathcal{E}_{0}$ already equals the chosen starting bundle (equivalently when the chain is trivial).
Remark 2.3.
  1. (i)
    Normalization. If one does not normalize the filtration (for instance by fixing $\det(\mathcal{E}_{0})$), the notion of hierarchical filtration (and the numerical data $(D_{i})$) is only defined up to simultaneous twist of every term by a line bundle. In particular one can “renormalize” a filtration by tensoring the entire chain with $\mathcal{O}_{X}(-D)$ and thereby create non-canonical shifts. Requiring $\det(\mathcal{E}_{0})\simeq\Lambda_{0}$ (for example $\Lambda_{0}=\mathcal{O}_{X}$) pins down this ambiguity and prevents the filtration from sliding to -$\infty$.
  2. (ii)
    Twist–invariance property. The hierarchical depth is stable under tensor twists, provided the normalization is adjusted accordingly: for any $M\in\mathrm{Pic}(X)$ one has
    (1) $h_{\Lambda_{0}\otimes M^{\otimes r}}(\mathcal{E}\otimes M)=h_{\Lambda_{0}}(\mathcal{E}),$
    where $r=\mathrm{rank}(\mathcal{E})$. Indeed, tensoring a normalized filtration of $\mathcal{E}$ by $M$ yields a normalized filtration of $\mathcal{E}\otimes M$ with respect to $\Lambda_{0}\otimes M^{\otimes r}$, since $\det(\mathcal{E}_{i}\otimes M)\simeq\det(\mathcal{E}_{i})\otimes M^{\otimes r}$ at every step. Hence the numerical invariant $h_{\Lambda_{0}}(\mathcal{E})$ depends only on the pair $(\det(\mathcal{E}),\Lambda_{0})$ modulo this canonical renormalization.
  3. (iii)
    Intrinsic interpretation. Equation (1) shows that the hierarchical depth is an intrinsic invariant of the bundle $\mathcal{E}$ up to tensor equivalence, analogous to how the slope $\mu(\mathcal{E})=\deg(\mathcal{E})/\mathrm{rank}(\mathcal{E})$ is invariant under twists by degree-zero line bundles on curves. The normalization by $\Lambda_{0}$ thus provides a canonical reference frame for measuring divisorial increments in $\mathrm{Pic}(X)$.
  4. (iv)
    Rank one case. When $r=1$ this definition recovers Definition 1.1 of [14] (with the same normalization choice $\Lambda_{0}=\mathcal{O}_{X}$ if one so desires): then $\mathcal{E}_{i}$ are line bundles and the determinant condition is the same as $\mathcal{E}_{i}\cong\mathcal{E}_{i-1}\otimes\mathcal{O}_{X}(D_{i})$.
Proposition 2.4.
Let $X$ be a smooth projective variety of dimension $n$ such that $\mathrm{Pic}(X)\cong\mathbb{Z}\cdot A$ for some ample effective divisor $A$. Fix a reference line bundle $\Lambda_{0}\in\mathrm{Pic}(X)$ and an ample polarization $\mathcal{O}_{X}(1)$. Let $\mathcal{E}$ be a vector bundle on $X$, and write
$\det(\mathcal{E})\simeq\mathcal{O}_{X}(dA)$
for a unique integer $d\in\mathbb{Z}$. Write $\Lambda_{0}\simeq\mathcal{O}_{X}(d_{0}A)$.
Then any hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$ has length bounded by
$h\leq d-d_{0}.$
In particular, when $\Lambda_{0}\cong\mathcal{O}_{X}$ (so $d_{0}=0$) one has the uniform upper bound
$h_{\Lambda_{0}}(\mathcal{E})\leq d.$
Moreover, equality $h_{\Lambda_{0}}(\mathcal{E})=d-d_{0}$ occurs whenever there exists a hierarchical filtration of $\mathcal{E}$ whose successive determinant increments are all linearly equivalent to $A$, i.e. there exist effective divisors $D_{1},\dots,D_{d-d_{0}}\sim A$ and a chain
$\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{d-d_{0}}=\mathcal{E}$
(normalized by $\det(\mathcal{E}_{0})\simeq\Lambda_{0}$) with $\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i})$ for every $i$.
Proof.
Let
$\mathfrak{F}:\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
be a hierarchical filtration normalized by $\Lambda_{0}$, so $\det(\mathcal{E}_{0})\simeq\Lambda_{0}$. For each $i$ there is a nonzero effective Cartier divisor $D_{i}$ with
$\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i}).$
Thus
$\det(\mathcal{E})\simeq\det(\mathcal{E}_{0})\otimes\mathcal{O}_{X}\Big(\sum_{i=1}^{h}D_{i}\Big),$
and in $\mathrm{NS}(X)$
$c_{1}(\det\mathcal{E})-c_{1}(\Lambda_{0})=\big[D_{1}+\cdots+D_{h}\big].$
Under the hypothesis $\mathrm{Pic}(X)\cong\mathbb{Z}\cdot A$, each effective Cartier divisor $D_{i}$ is linearly equivalent to a positive integer multiple of $A$. Hence there exist integers $d_{i}\geq 1$ such that
$D_{i}\sim d_{i}A\qquad(i=1,\dots,h).$
Therefore in $\mathrm{Pic}(X)$ we have
$\mathcal{O}_{X}(dA)\simeq\det(\mathcal{E})\simeq\Lambda_{0}\otimes\mathcal{O}_{X}\Big(\sum_{i=1}^{h}D_{i}\Big)\simeq\mathcal{O}_{X}\Big(d_{0}A+\sum_{i=1}^{h}d_{i}A\Big),$
so equality of divisor classes yields the integer identity
$d=d_{0}+\sum_{i=1}^{h}d_{i}.$
Since each $d_{i}\geq 1$ we obtain
$d-d_{0}=\sum_{i=1}^{h}d_{i}\geq h,$
which is exactly the desired inequality $h\leq d-d_{0}$.
The claim about equality is immediate: equality $h=d-d_{0}$ holds precisely when every $d_{i}=1$, i.e. when each determinant increment $D_{i}$ is linearly equivalent to $A$. This completes the proof. ∎
Remark 2.5.
(1) The upper bound $h_{\Lambda_{0}}(\mathcal{E})\leq d-d_{0}$ follows directly from the fact that, under the hypothesis $\mathrm{Pic}(X)\cong\mathbb{Z}\cdot A$, every effective increment $D_{i}$ satisfies $D_{i}\sim d_{i}A$ with $d_{i}\geq 1$. Thus
$d-d_{0}=\sum_{i=1}^{h}d_{i}\geq h,$
showing that the maximal possible hierarchical depth depends only on the difference of determinant classes $\det(\mathcal{E})$ and $\Lambda_{0}$, but not on the internal geometry of $\mathcal{E}$. This is the exact analogue of the line–bundle case.
(2) Unlike the line–bundle case, equality $h_{\Lambda_{0}}(\mathcal{E})=d-d_{0}$ is generally not automatic for vector bundles. Equality would require that each increment divisor satisfy $D_{i}\sim A$, i.e. that $d_{i}=1$ for all $i$ in the decomposition
$D_{i}\sim d_{i}A,\qquad d_{i}\geq 1.$
However, for a given bundle $\mathcal{E}$ there may be no hierarchical filtration whose determinant increments are all linearly equivalent to $A$. Geometrically, this means that although $\mathrm{Pic}(X)$ has rank one, the structure of $\mathcal{E}$ (e.g. its possible elementary transforms, base–locus behavior, or torsion constraints) may force certain increments to have degree $d_{i}>1$, reducing the achievable depth below the theoretical maximum $d-d_{0}$.
3. Elementary transforms and commutativity
Definition 3.1 (Elementary transformation).
[8, Definition 5.2.1] Let $X$ be a smooth projective surface and $C$ be an effective divisor on $X$. Let $i\colon C\hookrightarrow X$ denote the embedding $C\subset X$. Let $\mathcal{F}$ be a vector bundle on $X$ and let $\mathcal{G}$ be a vector bundle on $C$. A coherent sheaf $\mathcal{E}$ on $X$ is called an elementary transform of $\mathcal{F}$ along $\mathcal{G}$ if there exists a short exact sequence
(2) $0\longrightarrow\mathcal{E}\longrightarrow\mathcal{F}\longrightarrow i_{*}\mathcal{G}\longrightarrow 0.$
It follows from Proposition 5.2.2 from [8] that $\mathcal{E}$ is locally free and, moreover, it is a vector bundle since $X$ is smooth.
Lemma 3.2.
Let $X$ be a smooth projective variety and let $\mathcal{E}$ be a vector bundle on $X$. Let $D_{1},D_{2}\subset X$ be effective Cartier divisors with
$\mathrm{Supp}(D_{1})\cap\mathrm{Supp}(D_{2})=\varnothing.$
For $i=1,2$ let $\mathcal{Q}_{i}$ be a nonzero coherent torsion sheaf supported on $D_{i}$ and let $\varphi_{i}:\mathcal{E}\to\mathcal{Q}_{i}$ be surjective $\mathcal{O}_{X}$–linear maps. Define
$V_{1}:=\ker(\varphi_{1}),\qquad V_{2}:=\ker(\varphi_{2}),$
and the successive kernels
$V_{12}:=\ker\big(V_{1}\xrightarrow{\ \overline{\varphi}_{2}\ }\mathcal{Q}_{2}\big),\qquad V_{21}:=\ker\big(V_{2}\xrightarrow{\ \overline{\varphi}_{1}\ }\mathcal{Q}_{1}\big),$
where $\overline{\varphi}_{j}$ denotes the map induced by $\varphi_{j}$ on the relevant kernel. Finally set
$V:=\ker\big(\mathcal{E}\xrightarrow{\ (\varphi_{1},\varphi_{2})\ }\mathcal{Q}_{1}\oplus\mathcal{Q}_{2}\big).$
Then the following hold.
  1. (1)
    As subsheaves of $\mathcal{E}$ we have
    $V_{12}=V_{21}=V.$
    In particular, the order of taking kernels (elementary transforms) with respect to $\varphi_{1}$ and $\varphi_{2}$ does not affect the resulting subsheaf of $\mathcal{E}$.
  2. (2)
    If, in addition, the three kernels $V_{1},V_{2},V$ are locally free , then the equalities in (1) are equalities of subbundles of $\mathcal{E}$. In this case we say that the corresponding elementary transforms commute and the order of the transforms is immaterial in the category of vector bundles.
Proof.
(1) Since $\mathrm{Supp}(\mathcal{Q}_{1})$ and $\mathrm{Supp}(\mathcal{Q}_{2})$ are disjoint, the two surjective maps $\varphi_{1},\varphi_{2}$ impose independent conditions on local sections of $\mathcal{E}$. Concretely, for any open set $U\subset X$ we have
$V(U)$ $=\{s\in\mathcal{E}(U):\varphi_{1}(s)=0\text{ in }\mathcal{Q}_{1}(U),\ \varphi_{2}(s)=0\text{ in }\mathcal{Q}_{2}(U)\},$
$V_{12}(U)$ $=\{s\in\mathcal{E}(U):\varphi_{1}(s)=0\text{ and }\varphi_{2}(s)=0\},$
$V_{21}(U)$ $=\{s\in\mathcal{E}(U):\varphi_{2}(s)=0\text{ and }\varphi_{1}(s)=0\}.$
The equalities of the three sets above are immediate; hence the three kernels coincide as subsheaves of $\mathcal{E}$.
(2) If the kernels $V_{1},V_{2},V$ are locally free, then the equalities of subsheaves are equalities of locally free subsheaves (subbundles). In particular all constructions may be carried out inside the category of vector bundles and the elementary transforms commute as vector bundles. ∎
Remark 3.3.
(1) The disjointness hypothesis in Lemma 3.2 is essential. If $\mathrm{Supp}(D_{1})\cap\mathrm{Supp}(D_{2})\neq\varnothing$ then the two quotient maps may interact on the overlap and the iterated kernels need not coincide (indeed saturation may collapse intermediate steps). (2) The lemma is stated for arbitrary torsion quotients $\mathcal{Q}_{i}$ supported on the Cartier divisors. The additional local-freeness hypothesis in (2) is exactly the bridge to the Huybrechts–Lehn notion of elementary transform ([8]) where one typically takes $\mathcal{Q}_{i}=i_{*}(\mathcal{G}_{i})$ for a vector bundle $\mathcal{G}_{i}$ on $D_{i}$; in that case the kernel is usually a vector bundle.
Corollary 3.4.
Let $X$ be a smooth projective variety and let $\mathcal{E}$ be a vector bundle on $X$. Suppose a hierarchical filtration of $\mathcal{E}$ is given by a finite sequence of elementary transforms
$\mathcal{E}=V_{0}\supset V_{1}\supset\cdots\supset V_{h},\qquad V_{k+1}=\ker\bigl(V_{k}\xrightarrow{\ \varphi_{k}\ }\mathcal{Q}_{k}\bigr),$
where for each $k$ the sheaf $\mathcal{Q}_{k}$ is a nonzero torsion sheaf supported on an effective Cartier divisor $D_{k}\subset X$.
Assume that the set of divisors $\{D_{k}\}$ is partitioned into two disjoint collections
$\mathcal{A}=\{D_{a_{1}},\dots,D_{a_{r}}\},\qquad\mathcal{B}=\{D_{b_{1}},\dots,D_{b_{s}}\},$
such that
$\mathrm{Supp}(D_{a})\cap\mathrm{Supp}(D_{b})=\varnothing\qquad\text{for all }D_{a}\in\mathcal{A},\ D_{b}\in\mathcal{B}.$
Then:
  1. (1)
    For every pair $D_{a}\in\mathcal{A}$ and $D_{b}\in\mathcal{B}$, the elementary transforms along $D_{a}$ and $D_{b}$ commute. More precisely, for any surjective morphisms $\varphi_{a}:\mathcal{E}\to\mathcal{Q}_{a}$ and $\varphi_{b}:\mathcal{E}\to\mathcal{Q}_{b}$ with $\mathrm{Supp}(\mathcal{Q}_{a})\subset D_{a}$ and $\mathrm{Supp}(\mathcal{Q}_{b})\subset D_{b}$, one has isomorphisms of subsheaves of $\mathcal{E}$:
    $\ker\Bigl(\ker(\mathcal{E}\xrightarrow{\varphi_{a}}\mathcal{Q}_{a})\xrightarrow{\overline{\varphi}_{b}}\mathcal{Q}_{b}\Bigr)=\ker\Bigl(\ker(\mathcal{E}\xrightarrow{\varphi_{b}}\mathcal{Q}_{b})\xrightarrow{\overline{\varphi}_{a}}\mathcal{Q}_{a}\Bigr)=\ker\Bigl(\mathcal{E}\xrightarrow{\ (\varphi_{a},\varphi_{b})\ }\mathcal{Q}_{a}\oplus\mathcal{Q}_{b}\Bigr),$
    where $\overline{\varphi}_{j}$ denotes the map induced on the kernel of $\varphi_{i}$.
  2. (2)
    By repeatedly interchanging adjacent transforms involving one divisor from $\mathcal{A}$ and one from $\mathcal{B}$, the filtration can be reordered into another hierarchical filtration of the same length $h$, in which all transforms supported on divisors in $\mathcal{A}$ occur before all transforms supported on divisors in $\mathcal{B}$.
  3. (3)
    In particular, increments supported on disjoint divisors impose no obstruction to achieving the maximal possible hierarchical depth: their order can be rearranged arbitrarily without changing the total number of steps.
Proof.
We first prove (1). Let $D_{a}\in\mathcal{A}$, $D_{b}\in\mathcal{B}$ and suppose $\varphi_{a}:\mathcal{E}\to\mathcal{Q}_{a}$, $\varphi_{b}:\mathcal{E}\to\mathcal{Q}_{b}$ are surjective with $\mathrm{Supp}(\mathcal{Q}_{a})\subset D_{a}$, $\mathrm{Supp}(\mathcal{Q}_{b})\subset D_{b}$. Set
$V_{a}:=\ker(\varphi_{a}),\qquad V_{b}:=\ker(\varphi_{b}).$
Define the successive transforms
$V_{ab}:=\ker\bigl(V_{a}\xrightarrow{\overline{\varphi}_{b}}\mathcal{Q}_{b}\bigr),\qquad V_{ba}:=\ker\bigl(V_{b}\xrightarrow{\overline{\varphi}_{a}}\mathcal{Q}_{a}\bigr),$
and the simultaneous kernel
$V:=\ker\bigl(\mathcal{E}\xrightarrow{(\varphi_{a},\varphi_{b})}\mathcal{Q}_{a}\oplus\mathcal{Q}_{b}\bigr).$
Since $\mathrm{Supp}(D_{a})\cap\mathrm{Supp}(D_{b})=\varnothing$, the torsion sheaves $\mathcal{Q}_{a}$ and $\mathcal{Q}_{b}$ are supported on disjoint closed subsets. Therefore, for any section $s\in\mathcal{E}(U)$ on an open set $U\subset X$,
$\varphi_{a}(s)=0\quad\text{and}\quad\varphi_{b}(s)=0$
are independent conditions. Consequently,
$V(U)$ $=\{s\in\mathcal{E}(U):\varphi_{a}(s)=0,\ \varphi_{b}(s)=0\},$
$V_{ab}(U)$ $=\{s\in\mathcal{E}(U):\varphi_{a}(s)=0,\ \varphi_{b}(s)=0\},$
$V_{ba}(U)$ $=\{s\in\mathcal{E}(U):\varphi_{b}(s)=0,\ \varphi_{a}(s)=0\}.$
Thus $V=V_{ab}=V_{ba}$ as subsheaves of $\mathcal{E}$, proving commutativity.
We now prove (2). Consider the given filtration
$\mathcal{E}=V_{0}\supset V_{1}\supset\cdots\supset V_{h},\qquad V_{k+1}=\ker\bigl(V_{k}\xrightarrow{\varphi_{k}}\mathcal{Q}_{k}\bigr).$
Label each index $k$ as type $A$ or type $B$ according to whether $\mathrm{Supp}(\mathcal{Q}_{k})\subset D_{a}\in\mathcal{A}$ or $\mathrm{Supp}(\mathcal{Q}_{k})\subset D_{b}\in\mathcal{B}$. If a step of type $A$ is immediately followed by a step of type $B$, say
$V_{k}\supset V_{k+1}\supset V_{k+2},$
with $V_{k+1}=\ker(V_{k}\to\mathcal{Q}_{a})$ and $V_{k+2}=\ker(V_{k+1}\to\mathcal{Q}_{b})$, then by the commutativity just proved, the composite kernel
$\ker\bigl(V_{k}\to\mathcal{Q}_{a}\bigr)\xrightarrow{}\mathcal{Q}_{b}$
is equal to
$\ker\bigl(V_{k}\to\mathcal{Q}_{b}\bigr)\xrightarrow{}\mathcal{Q}_{a},$
and hence the transforms may be exchanged:
$\ker\bigl(V_{k}\xrightarrow{\varphi_{a}}\mathcal{Q}_{a}\bigr)\ \xrightarrow{\ \overline{\varphi}_{b}\ }\ Q_{b}\quad=\quad\ker\bigl(V_{k}\xrightarrow{\varphi_{b}}\mathcal{Q}_{b}\bigr)\ \xrightarrow{\ \overline{\varphi}_{a}\ }\ Q_{a}.$
Thus one may swap the AA– and BB–steps without changing the resulting subsheaf of $V_{k}$. By repeatedly swapping all out–of–order adjacent pairs (type $A$ followed by type $B$), one eventually obtains a filtration
$\mathcal{E}=V^{\prime}_{0}\supset V^{\prime}_{1}\supset\cdots\supset V^{\prime}_{h}$
in which all AA–steps precede all BB–steps. The length remains $h$ because no step is removed during the swapping.
Finally, statement (3) is immediate: reordering does not change the number of transform steps, hence disjoint support never forces a reduction in hierarchical depth.
Example 3.5.
(1) Let $X=\mathbb{P}^{1}\times\mathbb{P}^{1}$. Write $F:=\pi_{1}^{-1}(pt)$. Distinct fibers $F_{x}:=\pi_{1}^{-1}(x)$ for $x\in\mathbb{P}^{1}$ are pairwise disjoint effective Cartier divisors. Consider the split bundle
$\mathcal{E}=\bigoplus_{j=1}^{s}\mathcal{O}_{X}(a_{j}F),\qquad a_{j}\geq 0,\qquad d:=\sum_{j=1}^{s}a_{j}.$
For each summand $\mathcal{O}_{X}(a_{j}F)$ choose $a_{j}$ distinct points $x_{j,1},\dots,x_{j,a_{j}}\in\mathbb{P}^{1}$ such that all points $x_{j,t}$ are distinct as $(j,t)$ vary, and write
$a_{j}F=F_{x_{j,1}}+\cdots+F_{x_{j,a_{j}}}.$
Each unit inclusion
$\mathcal{O}_{X}((k{-}1)F)\subset\mathcal{O}_{X}(kF)$
is realized by an elementary transform along the divisor $F_{x_{j,k}}$. Since all fibers $F_{x_{j,t}}$ have pairwise disjoint supports, Lemma 3.2 implies that the corresponding elementary transforms commute, and hence Corollary 3.4 applies.
Thus the $d$ unit increments coming from all summands may be interleaved into a single saturated hierarchical filtration of $\mathcal{E}$. Therefore
$h_{\mathcal{O}_{X}}(\mathcal{E})=d.$
Actually, the class $F$ admits arbitrarily many pairwise disjoint effective representatives, so all elementary transforms occur along disjoint supports and therefore commute.
(2) Let $X=\mathbb{P}^{2}$ and let $H$ be the hyperplane class. Every effective divisor in $|H|$ is a line, and any two distinct lines meet in exactly one point. Hence no two nonzero divisors in $|H|$ have disjoint support.
For example, consider
$\mathcal{E}=\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(aH),\qquad a\geq 2.$
Although $\det(\mathcal{E})\simeq\mathcal{O}_{\mathbb{P}^{2}}(aH)$, the natural chain
$\mathcal{O}_{\mathbb{P}^{2}}\subset\mathcal{O}_{\mathbb{P}^{2}}(H)\subset\cdots\subset\mathcal{O}_{\mathbb{P}^{2}}(aH)$
cannot be realized by $a$ successive saturated elementary transforms inside $\mathcal{E}$. Each such step must occur along a divisor of type $H$, but every pair of such divisors intersects. Thus the elementary transforms fail to commute, and repeated kernels interact by saturation.
By Lemma 3.2, only transforms along disjoint divisors commute; here all supports intersect, so Corollary 3.4 does not apply. Consequently the unit–increment strategy collapses, and one typically obtains only a short saturated same–rank filtration such as
$\mathcal{O}_{\mathbb{P}^{2}}^{\oplus 2}\subset\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(aH)\subset\mathcal{E},$
of length22, with no possibility of refining it to length $a$.
Here the class $H$ has no pairwise disjoint effective representatives, so elementary transforms must occur along intersecting divisors and hence do not commute; saturation then forces the filtration length to collapse.
Consequently, these two surfaces illustrate the dichotomy between the numerical determinant bound $h(\mathcal{E})\leq d$ and the geometric possibility of achieving equality. If the linear system $|A|$ contains sufficiently many pairwise disjoint effective representatives (as on $\mathbb{P}^{1}\times\mathbb{P}^{1}$), one may realize each unit increment separately and obtain $h(\mathcal{E})=d$. If $|A|$ contains no such disjoint representatives (as on $\mathbb{P}^{2}$ with $A=H$), then geometric obstructions force noncommuting transforms and the maximal hierarchical depth may be strictly smaller than $d$.
4. Hierarchical depth and determinant bounds
In this section we collect basic numerical properties of hierarchical depth and establish general bounds in terms of the determinant. These results provide effective control of the invariant in simple geometric situations and will be used throughout the paper.
Proposition 4.1.
Let $C$ be a smooth projective curve and let
$\mathcal{E}=\bigoplus_{i=1}^{r}\mathcal{O}_{C}(d_{i})$
be a split vector bundle of rank $r$, with integers $d_{i}\in\mathbb{Z}$. Fix a normalization line bundle $\Lambda_{0}\in\mathrm{Pic}(C)$. Set
$M:=\deg\det(\mathcal{E})-\deg(\Lambda_{0})=\sum_{i=1}^{r}d_{i}-\deg(\Lambda_{0}).$
  1. (a)
    If $M<0$ then there exists no hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$, and by convention $h_{\Lambda_{0}}(\mathcal{E})=$-$\infty$.
  2. (b)
    If $M\geq 0$ then the hierarchical depth of $\mathcal{E}$ with respect to $\Lambda_{0}$ equals
    $h_{\Lambda_{0}}(\mathcal{E})=M.$
    In particular, when $\Lambda_{0}=\mathcal{O}_{C}$ one has
    $h(\mathcal{E})=\sum_{i=1}^{r}d_{i}.$
Proof.
We assume $M\geq 0$; the negative case (no filtration exists) is immediate from the determinant relation and the normalization requirement.
Let
$\mathfrak{F}:\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
be any hierarchical filtration normalized by $\Lambda_{0}$, i.e. $\det(\mathcal{E}_{0})\simeq\Lambda_{0}$ and for each $i$
$\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{C}(p_{i})$
for some effective divisor $p_{i}$ (a point of $C$, counted with multiplicity). Taking determinants and degrees yields
$\deg\det(\mathcal{E})-\deg(\Lambda_{0})=\sum_{i=1}^{h}\deg(p_{i})=h,$
since each $p_{i}$ is a single point (degree $1$). Hence $h\leq M$.
We now produce a hierarchical filtration of length exactly $M$.
Choose $M$ distinct points $q_{1},\dots,q_{M}\in C$. (If the base field is finite and $C(\mathbb{F}_{q})$ has fewer than $M$ rational points, work over a finite extension or choose geometric points; the statement and proof are unchanged up to base change.) For each $j$ define inductively a subsheaf by performing an elementary transform at $q_{j}$:
$\mathcal{E}^{(M)}:=\mathcal{E},\quad\mathcal{E}^{(M-1)}:=\ker\bigl(\mathcal{E}^{(M)}\xrightarrow{\ \rho_{M}\ }k(q_{M})\bigr),\quad\mathcal{E}^{(M-2)}:=\ker\bigl(\mathcal{E}^{(M-1)}\xrightarrow{\ \rho_{M-1}\ }k(q_{M-1})\bigr),$
and so on, where each $\rho_{j}$ is a chosen surjection of the fibre at $q_{j}$ onto the one-dimensional skyscraper $k(q_{j})$.
On a curve the kernel of a surjection from a vector bundle to a skyscraper is again locally free (torsion-free on a curve implies locally free), so each $\mathcal{E}^{(t)}$ is a vector bundle of rank $r$ and sits as a saturated subsheaf of $\mathcal{E}^{(t+1)}$. Moreover the determinant relation at each step is
$\det(\mathcal{E}^{(t+1)})\simeq\det(\mathcal{E}^{(t)})\otimes\mathcal{O}_{C}(q_{t+1}).$
After performing the $M$ transforms we obtain
$\det(\mathcal{E}^{(M)})\simeq\det(\mathcal{E}^{(0)})\otimes\mathcal{O}_{C}\Big(\sum_{j=1}^{M}q_{j}\Big).$
Rearranging,
$\det(\mathcal{E}^{(0)})\simeq\det(\mathcal{E})\otimes\mathcal{O}_{C}\Big(-\sum_{j=1}^{M}q_{j}\Big).$
By construction the right hand side is a line bundle of degree $\deg\det(\mathcal{E})-M=\deg\Lambda_{0}$; hence $\det(\mathcal{E}^{(0)})\simeq\Lambda_{0}$. Thus the ascending chain
$\mathcal{E}^{(0)}\subset\mathcal{E}^{(1)}\subset\cdots\subset\mathcal{E}^{(M)}=\mathcal{E}$
is a hierarchical filtration normalized by $\Lambda_{0}$ of length $M$.
It remains to check that the $M$ steps are distinct (no collapse via saturation). Because we chose the points $q_{j}$ to be pairwise distinct, the successive elementary transforms are supported on distinct points and therefore commute and produce distinct saturated kernels; equivalently, each step increases the determinant by a nonzero effective divisor, so no further identification occurs. Hence the length of the constructed filtration is exactly $M$.
Combining the upper bound and the explicit construction shows $h_{\Lambda_{0}}(\mathcal{E})=M$, as claimed. ∎
Remark 4.2.
Proposition 4.1 is the one–dimensional specialization of Proposition 2.4. When $\dim X=1$, every divisor linearly equivalent to $A$ is a single point and elementary transforms at distinct points commute. Hence the geometric existence condition in the higher–dimensional proposition is automatic for split bundles, and the determinant bound specializes to an exact formula for curves.
Proposition 4.3 (Basic properties of hierarchical depth).
Let $X$ be a smooth projective variety, $\Lambda_{0}\in\mathrm{Pic}(X)$ a reference line bundle, and $\mathcal{E}$ a vector bundle of rank $r$ on $X$.
  1. (A)
    (Monotonicity) If $\mathcal{E}^{\prime}\subset\mathcal{E}$ is a saturated inclusion of vector bundles of the same rank $r$, then
    $h_{\Lambda_{0}}(\mathcal{E}^{\prime})\leq h_{\Lambda_{0}}(\mathcal{E}).$
  2. (B)
    (Extension inequality) Let $0\to\mathcal{E}^{\prime}\to\mathcal{E}\to\mathcal{E}^{\prime\prime}\to 0$ be a short exact sequence of vector bundles on $X$.
    Suppose there exist hierarchical filtrations
    $\mathfrak{F}^{\prime}:\mathcal{E}^{\prime}_{0}\subset\mathcal{E}^{\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime}_{a}=\mathcal{E}^{\prime},\qquad\mathfrak{F}^{\prime\prime}:\mathcal{E}^{\prime\prime}_{0}\subset\mathcal{E}^{\prime\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime\prime}_{b}=\mathcal{E}^{\prime\prime},$
    normalized by $\Lambda_{0}$ and $\Lambda_{0}^{\prime}:=\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})^{-1}$ respectively. Assume that for each $j=1,\dots,b$, the elementary transform $\mathcal{E}^{\prime\prime}_{j-1}\subset\mathcal{E}^{\prime\prime}_{j}$ along an effective divisor $D_{j}$ lifts to an elementary transform of $\mathcal{E}$ in the following sense: there exists a commutative diagram with exact rows
    ${0}{\mathcal{E}^{\prime}}$$\mathcal{F}_{j-1}$$\mathcal{E}^{\prime\prime}_{j-1}$$0{0}{\mathcal{E}^{\prime}}$$\mathcal{F}_{j}$$\mathcal{E}^{\prime\prime}_{j}$}0{0}
    where:
    • $\mathcal{F}_{j-1},\mathcal{F}_{j}\subset\mathcal{E}$ are saturated subbundles of rank $\mathrm{rank}(\mathcal{E})$,
    • the vertical inclusions $\mathcal{F}_{j-1}\hookrightarrow\mathcal{F}_{j}$ and $\mathcal{E}^{\prime\prime}_{j-1}\hookrightarrow\mathcal{E}^{\prime\prime}_{j}$ are elementary transforms along $D_{j}$,
    • the middle vertical map induces the given inclusion $\mathcal{E}^{\prime\prime}_{j-1}\hookrightarrow\mathcal{E}^{\prime\prime}_{j}$ on the quotients.
    Then
    $h_{\Lambda_{0}}(\mathcal{E})\geq h_{\Lambda_{0}}(\mathcal{E}^{\prime})+h_{\Lambda^{\prime}_{0}}(\mathcal{E}^{\prime\prime}).$
  3. (C)
    (Direct sums) If $\mathcal{E}=\mathcal{E}_{1}\oplus\mathcal{E}_{2}$ and there exist maximal hierarchical filtrations of $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ whose determinant increments have pairwise disjoint support, then
    $h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0}}(\mathcal{E}_{1})+h_{\Lambda_{0}}(\mathcal{E}_{2}).$
    Without disjoint supports, only the inequality $\geq$ holds in general.
Proof.
(A) Let $\mathcal{E}^{\prime}_{0}\subset\cdots\subset\mathcal{E}^{\prime}_{h^{\prime}}=\mathcal{E}^{\prime}$ be a hierarchical filtration of $\mathcal{E}^{\prime}$ normalized by $\Lambda_{0}$. Since $\mathcal{E}^{\prime}\subset\mathcal{E}$ is saturated of the same rank, each $\mathcal{E}^{\prime}_{i}\subset\mathcal{E}$ remains saturated and the quotients $\mathcal{E}/\mathcal{E}^{\prime}_{i}$ are torsion with the same support as $\mathcal{E}^{\prime}/\mathcal{E}^{\prime}_{i}$. Thus the same chain is a hierarchical filtration of $\mathcal{E}$, yielding $h_{\Lambda_{0}}(\mathcal{E})\geq h^{\prime}$. Taking suprema gives $h_{\Lambda_{0}}(\mathcal{E}^{\prime})\leq h_{\Lambda_{0}}(\mathcal{E})$.
[…]
Step 1: Embedding the filtration of $\mathcal{E}^{\prime}$.
By (A) (monotonicity), the hierarchical filtration $\mathfrak{F}^{\prime}$ of $\mathcal{E}^{\prime}$ embeds into $\mathcal{E}$: each $\mathcal{E}^{\prime}_{i}\subset\mathcal{E}^{\prime}\subset\mathcal{E}$ is saturated in $\mathcal{E}$ and the quotients $\mathcal{E}/\mathcal{E}^{\prime}_{i}$ are torsion with the same support as $\mathcal{E}^{\prime}/\mathcal{E}^{\prime}_{i}$. Thus we obtain a chain of saturated subbundles of $\mathcal{E}$:
$\mathcal{E}^{\prime}_{0}\subset\mathcal{E}^{\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime}_{a}=\mathcal{E}^{\prime}\subset\mathcal{E}.$
Each inclusion $\mathcal{E}^{\prime}_{i-1}\subset\mathcal{E}^{\prime}_{i}$ remains an elementary transform along some effective divisor $C_{i}$, and $\det(\mathcal{E}^{\prime}_{0})\simeq\Lambda_{0}$ by hypothesis.
Step 2: Lifting the filtration of $\mathcal{E}^{\prime\prime}$ step by step.
Set $\mathcal{F}_{0}:=\mathcal{E}^{\prime}$ (the last term of (2)). By assumption, for $j=1$ we have a diagram ($\star_{1}$) with $\mathcal{F}_{0}=\mathcal{E}^{\prime}$ and $\mathcal{F}_{1}\subset\mathcal{E}$ a saturated subbundle fitting into
$0\to\mathcal{E}^{\prime}\to\mathcal{F}_{1}\to\mathcal{E}^{\prime\prime}_{1}\to 0,$
and $\mathcal{F}_{0}\subset\mathcal{F}_{1}$ is an elementary transform along $D_{1}$. The determinant satisfies
$\det(\mathcal{F}_{1})\simeq\det(\mathcal{F}_{0})\otimes\mathcal{O}_{X}(D_{1})\simeq\det(\mathcal{E}^{\prime})\otimes\mathcal{O}_{X}(D_{1}).$
Proceed inductively: having constructed $\mathcal{F}_{j-1}\subset\mathcal{E}$ with $0\to\mathcal{E}^{\prime}\to\mathcal{F}_{j-1}\to\mathcal{E}^{\prime\prime}_{j-1}\to 0$, diagram ($\star_{j}$) provides $\mathcal{F}_{j}\subset\mathcal{E}$ containing $\mathcal{F}_{j-1}$ as an elementary transform along $D_{j}$, and $0\to\mathcal{E}^{\prime}\to\mathcal{F}_{j}\to\mathcal{E}^{\prime\prime}_{j}\to 0$. The determinant relation is
$\det(\mathcal{F}_{j})\simeq\det(\mathcal{F}_{j-1})\otimes\mathcal{O}_{X}(D_{j}).$
Step 3: Concatenation yields a hierarchical filtration of $\mathcal{E}$.
Chaining (2) with the lifts of the $\mathcal{E}^{\prime\prime}$-filtration gives a sequence
$\mathcal{E}^{\prime}_{0}\subset\cdots\subset\mathcal{E}^{\prime}_{a}=\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\cdots\subset\mathcal{F}_{b}=:\widetilde{\mathcal{E}}$.
We claim $\widetilde{\mathcal{E}}$$\mathcal{E}$. Indeed, $\mathcal{F}_{b}$ fits into $0\to\mathcal{E}^{\prime}\to\mathcal{F}_{b}\to\mathcal{E}^{\prime\prime}_{b}=\mathcal{E}^{\prime\prime}\to 0$, and the inclusion $\mathcal{F}_{b}\subset\mathcal{E}$ induces the identity on $\mathcal{E}^{\prime}$ and the given surjection $\mathcal{E}\to\mathcal{E}^{\prime\prime}$. Hence $\mathcal{F}_{b}=\mathcal{E}$ by the five‑lemma (or by comparing ranks and the fact that both are subbundles of $\mathcal{E}$).
Each inclusion in (4) is saturated of rank $r$ and the successive quotients are torsion sheaves supported on effective divisors ($C_{i}$ for the $\mathcal{E}^{\prime}$-part, $D_{j}$ for the $\mathcal{E}^{\prime\prime}$-part). Moreover, from (3) and the normalization of $\mathfrak{F}^{\prime\prime}$ we have
$\det(\mathcal{F}_{j})\simeq\det(\mathcal{E}^{\prime})\otimes\det(\mathcal{E}^{\prime\prime}_{j})\simeq\det(\mathcal{E}^{\prime})\otimes\bigl(\Lambda_{0}^{\prime}\otimes\mathcal{O}_{X}(D_{1}+\cdots+D_{j})\bigr)\simeq\Lambda_{0}\otimes\mathcal{O}_{X}(D_{1}+\cdots+D_{j}),$
so the determinant increments match those of $\mathfrak{F}^{\prime\prime}$. Consequently, (4) is a hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$, of total length $a+b$.
Step 4: Conclusion.
Since $a=h_{\Lambda_{0}}(\mathcal{E}^{\prime})$ and $b=h_{\Lambda_{0}^{\prime}}(\mathcal{E}^{\prime\prime})$ by definition of hierarchical depth, the existence of the filtration (4) shows
$h_{\Lambda_{0}}(\mathcal{E})\geq a+b=h_{\Lambda_{0}}(\mathcal{E}^{\prime})+h_{\Lambda^{\prime}_{0}}(\mathcal{E}^{\prime\prime}),$
which is precisely inequality (1).
(C) Under the disjoint support hypothesis, Lemma 3.2 ensures the elementary transforms from the two filtrations commute. Interleaving them produces a hierarchical filtration of $\mathcal{E}$ of length $h_{\Lambda_{0}}(\mathcal{E}_{1})+h_{\Lambda_{0}}(\mathcal{E}_{2})$, giving equality. If supports intersect, commutativity may fail and saturation can reduce the achievable length, yielding only the inequality. ∎
Remark 4.4.
(Non-monotonicity for smaller rank) If $\mathcal{E}^{\prime}\subset\mathcal{E}$ is a subbundle with $\mathrm{rank}(\mathcal{E}^{\prime})<\mathrm{rank}(\mathcal{E})$, the inequality $h_{\Lambda_{0}}(\mathcal{E}^{\prime})\leq h_{\Lambda_{0}}(\mathcal{E})$ may fail. For example, on $X=\mathbb{P}^{1}$ with $\Lambda_{0}=\mathcal{O}_{\mathbb{P}^{1}}$, take $\mathcal{E}^{\prime}=\mathcal{O}(1)\subset\mathcal{E}=\mathcal{O}(1)\oplus\mathcal{O}(-1)$. Then $h(\mathcal{E}^{\prime})=\deg\mathcal{O}(1)=1$ while $h(\mathcal{E})=\deg\det(\mathcal{E})=0$.
5. Effects of the Minimal Model Program on hierarchical filtrations
In this section we study the behavior of hierarchical filtrations of vector bundles under birational transformations arising in the Minimal Model Program (MMP) when the base variety has dimension at most $2$. For smooth projective surfaces, the MMP consists solely of contractions of (1)(-1)-curves, which preserve smoothness.
5.1. Setup and motivation
Let $X$ be a smooth projective variety of dimension $n\leq 2$ over an algebraically closed field $\mathbb{F}$. The MMP for $X$ consists of a finite sequence of birational morphisms
$X=X_{0}\dashrightarrow X_{1}\dashrightarrow\cdots\dashrightarrow X_{t}=X_{\min},$
where each step is either an isomorphism in codimension $1$ or a contraction of a (1)(-1)–curve (in the surface case), and the final model $X_{\min}$ is minimal, i.e. $K_{X_{\min}}$ is nef. In dimension $1$ the sequence is trivial since every smooth curve is already minimal (see [10]).
Given a vector bundle $\mathcal{E}$ on $X$, we are interested in how its hierarchical filtrations behave under the birational maps of such an MMP sequence, and in particular how the hierarchical depth changes after blowing up or contracting divisors. This analysis provides a bridge between the birational geometry of $X$ and the algebraic structure of $\mathcal{E}$, and will later underlie the construction of evaluation codes associated to successive divisorial increments.
5.2. Birational pullback of filtrations
Let $f:Y\to X$ be a birational morphism between smooth projective varieties of dimension $\leq 2$, for instance the blow–up of $X$ at a smooth point with exceptional divisor $E$. Given a vector bundle $\mathcal{E}$ on $X$ and a reference line bundle $\Lambda_{0}$, consider its pullback $f^{*}\mathcal{E}$ on $Y$. The determinant relation satisfies
$\det(f^{*}\mathcal{E})\simeq f^{*}(\det\mathcal{E}),\qquad c_{1}(\det(f^{*}\mathcal{E}))=f^{*}c_{1}(\det\mathcal{E}).$
Hence a hierarchical filtration
$\mathfrak{F}:\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
on $X$ normalized by $\Lambda_{0}$ pulls back naturally to a filtration
$f^{*}\mathfrak{F}:\ f^{*}\mathcal{E}_{0}\subset f^{*}\mathcal{E}_{1}\subset\cdots\subset f^{*}\mathcal{E}_{h}=f^{*}\mathcal{E}$
on $Y$, normalized by $f^{*}\Lambda_{0}$. At each step one has
$\det(f^{*}\mathcal{E}_{i})\simeq\det(f^{*}\mathcal{E}_{i-1})\otimes\mathcal{O}_{Y}(f^{*}D_{i}),$
so that the divisorial increments are transformed by pullback: $D_{i}$ is replaced by $f^{*}D_{i}$. Consequently,
$h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})\geq h_{\Lambda_{0}}(\mathcal{E}).$
In particular, the hierarchical depth is invariant under birational pullback of smooth varieties in dimension $\leq 2$. Consequently, we have the following statement:
Proposition 5.1.
Let $f:Y\to X$ be a birational morphism between smooth projective varieties of dimension $\leq 2$, and let $\mathcal{E}$ be a vector bundle on $X$. Then
$h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})\geq h_{\Lambda_{0}}(\mathcal{E}).$
5.3. Effect of blow–ups on new divisorial increments
Let $f:Y\to X$ be the blow–up of $X$ at a smooth point $p\in X$ with exceptional divisor $E\simeq\mathbb{P}^{n-1}$. If $\mathcal{E}$ is a vector bundle on $X$, the pullback $f^{*}\mathcal{E}$ remains locally free. However, the Picard group changes:
$\mathrm{Pic}(Y)\cong f^{*}\mathrm{Pic}(X)\oplus\mathbb{Z}\langle E\rangle.$
Therefore, new effective divisors supported on the exceptional locus can appear in hierarchical filtrations of $f^{*}\mathcal{E}$ that have no counterpart on $X$.
The statements of two following lemmas can be found in [5, Ch. V].
Lemma 5.2.
Let $X$ be a smooth projective variety over an algebraically closed field, let $p\in X$ be a smooth (closed) point, and let $f:\,Y=\operatorname{Bl}_{p}X\to X$ be the blow–up of $X$ at $p$. Let $E\subset Y$ denote the exceptional divisor (which is isomorphic to $\mathbb{P}^{\,\dim X-1}$). Then the pullback map $f^{*}:\mathrm{Pic}(X)\to\mathrm{Pic}(Y)$ is injective, and every line bundle $\mathcal{L}\in\mathrm{Pic}(Y)$ can be written uniquely in the form
$\mathcal{L}\simeq f^{*}M\otimes\mathcal{O}_{Y}(mE)$
for a unique $M\in\mathrm{Pic}(X)$ and a unique integer $m\in\mathbb{Z}$. Equivalently, there is a direct sum decomposition of abelian groups
$\mathrm{Pic}(Y)\cong f^{*}\mathrm{Pic}(X)\oplus\mathbb{Z}\langle E\rangle.$
Proof.
Injectivity of $f^{*}$. Suppose $M\in\mathrm{Pic}(X)$ satisfies $f^{*}M\cong\mathcal{O}_{Y}$. Apply the pushforward functor $f_{*}$ and use the projection formula together with $f_{*}\mathcal{O}_{Y}\cong\mathcal{O}_{X}$ (since $f$ is a birational morphism contracting the exceptional divisor to a point):
$f_{*}f^{*}M\cong M\otimes f_{*}\mathcal{O}_{Y}\cong M.$
But $f_{*}f^{*}M\cong f_{*}\mathcal{O}_{Y}\cong\mathcal{O}_{X}$ because $f^{*}M\cong\mathcal{O}_{Y}$, hence $M\cong\mathcal{O}_{X}$. Thus $f^{*}$ is injective.
Existence of the decomposition. Let $\mathcal{L}\in\mathrm{Pic}(Y)$. Restrict $\mathcal{L}$ to the exceptional divisor $E\cong\mathbb{P}^{\,\dim X-1}$. Since $\mathrm{Pic}(E)\cong\mathbb{Z}$ and is generated by $\mathcal{O}_{E}(1)$, there is a unique integer $m$ with
$\mathcal{L}|_{E}\cong\mathcal{O}_{E}(m).$
Consider the twist $\mathcal{L}(-mE):=\mathcal{L}\otimes\mathcal{O}_{Y}(-mE)$. By construction $\mathcal{L}(-mE)|_{E}\cong\mathcal{O}_{E}$ is trivial on $E$. The line bundle $\mathcal{L}(-mE)$ is therefore trivial on the exceptional locus, hence it descends to $X$ in the following sense: set
$M:=f_{*}\big(\mathcal{L}(-mE)\big).$
The sheaf $M$ is reflexive of rank $1$ on the smooth variety $X$, hence locally free of rank $1$, i.e. a line bundle (see e.g. [5, II, Ex. 5.15]). Moreover $f^{*}M$ and $\mathcal{L}(-mE)$ are isomorphic away from $E$. Both are line bundles on the normal variety $Y$ and agree on the complement of a divisor, hence they are isomorphic globally:
$f^{*}M\cong\mathcal{L}(-mE).$
Therefore $\mathcal{L}\cong f^{*}M\otimes\mathcal{O}_{Y}(mE)$.
Uniqueness. If $f^{*}M\otimes\mathcal{O}_{Y}(mE)\cong f^{*}M^{\prime}\otimes\mathcal{O}_{Y}(m^{\prime}E)$, restrict to $E$ to compare integers: the left side restricts to $\mathcal{O}_{E}(m)$ and the right side to $\mathcal{O}_{E}(m^{\prime})$, hence $m=m^{\prime}$. Cancelling $\mathcal{O}_{Y}(mE)$ and using injectivity of $f^{*}$ gives $M\cong M^{\prime}$. This proves uniqueness of the pair $(M,m)$.
Group decomposition. The map
$f^{*}\mathrm{Pic}(X)\oplus\mathbb{Z}\langle E\rangle\longrightarrow\mathrm{Pic}(Y),\qquad(M,m)\longmapsto f^{*}M\otimes\mathcal{O}_{Y}(mE)$
is therefore bijective and respects the group law, yielding the claimed direct sum decomposition. ∎
Lemma 5.3.
Let $D$ be an effective Cartier divisor on $Y$. Then $D$ decomposes uniquely as
$D=f^{*}D_{X}+mE$
with $D_{X}$ an effective Cartier divisor on $X$ and $m\in\mathbb{Z}_{\geq 0}$. Moreover $D$ is supported on $E$ if and only if $D_{X}=0$.
Proof.
Write the line bundle $\mathcal{O}_{Y}(D)$ as $f^{*}\mathcal{O}_{X}(D_{X})\otimes\mathcal{O}_{Y}(mE)$ using Lemma 5.2. If $\mathcal{O}_{Y}(D)$ is effective then pulling down by $f_{*}$ shows $\mathcal{O}_{X}(D_{X})$ is effective and $m\geq 0$ (otherwise the section would vanish identically along the exceptional divisor in contradiction with effectivity). Uniqueness follows from the uniqueness in Lemma 5.2. ∎
Recall the concept of saturation from [8]. Let $\mathcal{F}$ be a subsheaf of a coherent sheaf $\mathcal{E}$ on $X$. The saturation of $\mathcal{F}$ inside $\mathcal{E}$ is defined as the minimal subsheaf $\mathcal{F}^{\prime}$ containing $\mathcal{F}$ such that $\mathcal{E}/\mathcal{F}^{\prime}$ is pure of dimension $d=\dim(\mathcal{E})$ or zero. Equivalently, the saturation of $\mathcal{F}$ is
$\mathcal{F}^{\mathrm{sat}}:=\ker\bigl(\mathcal{E}\longrightarrow\mathcal{E}/\mathcal{F}\longrightarrow(\mathcal{E}/\mathcal{F})/T(\mathcal{E}/\mathcal{F})\bigr),$
where $T(\mathcal{E}/\mathcal{F})$ denotes the maximal subsheaf of $\mathcal{E}/\mathcal{F}$ of dimension $\leq d-1$. Equivalently, $\mathcal{F}^{\mathrm{sat}}$ is the largest subsheaf of $\mathcal{E}$ that coincides with $\mathcal{F}$ on a dense open set and is torsion–free. When $\mathcal{E}=\mathcal{O}_{X}$, saturation simply removes 0–dimensional torsion.
As an example, suppose the sequence of line bundles
$\mathcal{O}_{Y}(-2E)\subset\mathcal{O}_{Y}(-E)\subset\mathcal{O}_{Y}.$
Applying $f_{*}$ gives
$f_{*}\mathcal{O}_{Y}(-2E)=\mathcal{I}_{p/X}^{2},\qquad f_{*}\mathcal{O}_{Y}(-E)=\mathcal{I}_{p/X},\qquad f_{*}\mathcal{O}_{Y}=\mathcal{O}_{X}.$
Each of these is already torsion-free on $X$ (they are ideal sheaves of a smooth point), hence saturation does nothing:
$(\mathcal{I}_{p/X}^{2})^{\mathrm{sat}}=\mathcal{I}_{p/X}^{2},\qquad(\mathcal{I}_{p/X})^{\mathrm{sat}}=\mathcal{I}_{p/X},\qquad(\mathcal{O}_{X})^{\mathrm{sat}}=\mathcal{O}_{X}.$
Thus the saturated chain is simply
$\mathcal{I}_{p/X}^{2}\subset\mathcal{I}_{p/X}\subset\mathcal{O}_{X}.$
Proposition 5.4.
Let $f:Y\to X$ be the blowup of a smooth projective variety $X$ at a smooth closed point $p\in X$, and denote by $E\subset Y$ the exceptional divisor. Assume $\dim X\leq 2$. Let $\mathcal{E}$ be a vector bundle on $X$ and fix a normalization $\Lambda_{0}\in\mathrm{Pic}(X)$.
Then every hierarchical filtration of $f^{*}\mathcal{E}$ normalized by $f^{*}\Lambda_{0}$ has length
$h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})=h_{\Lambda_{0}}(\mathcal{E})+\delta,\qquad\delta\in\mathbb{Z}_{\geq 0},$
where $\delta$ equals the number of additional determinant increments in a maximal filtration of $f^{*}\mathcal{E}$ whose divisors carry a positive coefficient of the exceptional divisor $E$. In particular $\delta=0$ unless the filtration involves elementary transforms along $E$.
Proof.
Step 1. If
$\mathfrak{F}_{X}:\ \mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
is a hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$, then pulling back termwise yields
$f^{*}\mathfrak{F}_{X}:\ f^{*}\mathcal{E}_{0}\subset f^{*}\mathcal{E}_{1}\subset\cdots\subset f^{*}\mathcal{E}_{h}=f^{*}\mathcal{E},$
a hierarchical filtration of $f^{*}\mathcal{E}$ normalized by $f^{*}\Lambda_{0}$. It follows from Proposition 5.1 that
$h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})\ \geq\ h_{\Lambda_{0}}(\mathcal{E}).$
Thus any extra length in $h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})$ must come from steps whose torsion quotient is supported on $E$.
Step 2. Let
$\mathfrak{F}_{Y}:\ \mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\cdots\subset\mathcal{F}_{h^{\prime}}=f^{*}\mathcal{E}$
be a maximal hierarchical filtration of $f^{*}\mathcal{E}$ normalized by $f^{*}\Lambda_{0}$. For each step write the determinant increment as an effective Cartier divisor $D_{i}\subset Y$. By the decomposition of effective divisors on a blowup (Lemma 5.3), each $D_{i}$ splits uniquely as
$D_{i}=f^{*}D_{i,X}+m_{i}E,\qquad D_{i,X}\geq 0,\ m_{i}\in\mathbb{Z}_{\geq 0}.$
Collect the indices into two disjoint sets:
$I_{0}:=\{i:\ m_{i}=0\},\qquad I_{E}:=\{i:\ m_{i}>0\},$
and write $h_{0}:=|I_{0}|$ and $\delta:=|I_{E}|$. Then $h^{\prime}=h_{0}+\delta$.
Step 3. For each $i\in I_{0}$ the divisor $D_{i}$ is the pullback of an effective Cartier divisor $D_{i,X}$ on $X$. The corresponding step $\mathcal{F}_{i-1}\subset\mathcal{F}_{i}$ on $Y$ therefore descends (objectwise) to an inclusion of torsion–free sheaves on $X$ after applying pushforward and saturation:
$\mathcal{G}_{i}:=\big(f_{*}\mathcal{F}_{i}\big)^{\mathrm{sat}}.$
Because $\dim X\leq 2$ the saturation (reflexive hull) of a torsion-free rank-$r$ sheaf is locally free (see for instance [5, II, Ex. 5.15] or standard references on reflexive sheaves). Hence each $\mathcal{G}_{i}$ is a vector bundle on $X$, and the inclusions $\mathcal{G}_{i-1}\subset\mathcal{G}_{i}$ are saturated with torsion quotients supported on $D_{i,X}$. Collecting these descended steps yields a hierarchical filtration on $X$ of length $h_{0}$, normalized by $\Lambda_{0}$. Consequently
$h_{0}\leq h_{\Lambda_{0}}(\mathcal{E}).$
Conversely, pulling back any maximal filtration on $X$ reproduces a filtration on $Y$ whose non-exceptional increments give at least $h_{\Lambda_{0}}(\mathcal{E})$ steps with $m_{i}=0$. Therefore $h_{0}\geq h_{\Lambda_{0}}(\mathcal{E})$, and combining the two inequalities gives
$h_{0}=h_{\Lambda_{0}}(\mathcal{E}).$
Step 4. By Step 2 we have $h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})=h_{0}+\delta$, and by Step 3, $h_{0}=h_{\Lambda_{0}}(\mathcal{E})$. Hence
$h_{f^{*}\Lambda_{0}}(f^{*}\mathcal{E})=h_{\Lambda_{0}}(\mathcal{E})+\delta,$
with $\delta\geq 0$ equal to the number of determinant increments whose divisors carry a positive coefficient of $E$.
Step 5. Finally, each elementary transform along $E$ produces a torsion quotient supported on $E$ and increases the exceptional coefficient of the determinant by precisely the corresponding multiplicity; conversely any positive coefficient of $E$ in some increment $D_{i}$ arises from at least one such elementary transform (after saturation). Therefore the number $\delta$ equals the number of additional elementary transforms supported on $E$ in the maximal filtration, i.e. the number of extra steps contributed by exceptional transforms.
This completes the proof. ∎
5.4. Contraction and preservation of depth
Proposition 5.5.
Let $f:Y\to X$ be the contraction of a (1)(-1)–curve $E\subset Y$ onto a smooth point $p=f(E)\in X$, and let $\mathcal{E}_{Y}$ be a vector bundle on $Y$ admitting a hierarchical filtration
$\mathfrak{F}_{Y}:\quad 0=\mathcal{E}_{0,Y}\subset\mathcal{E}_{1,Y}\subset\cdots\subset\mathcal{E}_{h,Y}=\mathcal{E}_{Y}$
normalized by $\Lambda_{0,Y}\in\mathrm{Pic}(Y)$. Assume that for every step of the filtration, the determinant increment divisor $D_{i,Y}$ satisfies
$E\not\subset\mathrm{Supp}D_{i,Y}.$
Then the direct image $\mathcal{E}:=(f_{*}\mathcal{E}_{Y})^{\mathrm{sat}}$ carries a natural hierarchical filtration
$\mathfrak{F}_{X}:\quad 0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E},$
normalized by $\Lambda_{0}:=\big(f_{*}\Lambda_{0,Y}\big)^{\mathrm{sat}}\in\mathrm{Pic}(X)$, and the hierarchical depths coincide:
$h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,Y}}(\mathcal{E}_{Y}).$
Proof.
We proceed in several steps.
Step 1. The morphism $f:Y\to X$ is an isomorphism over $U:=X\setminus\{p\}$, and $f^{-1}(U)=Y\setminus E\simeq U$. For any locally free sheaf $\mathcal{G}$ on $Y$, the pushforward $f_{*}\mathcal{G}$ is torsion–free on $X$, and since $f$ is birational between smooth surfaces, $f_{*}\mathcal{G}$ is reflexive (see, e.g., [6, Prop. 1.6]). Because reflexive sheaves on smooth surfaces are locally free, $(f_{*}\mathcal{G})^{\vee\vee}\cong f_{*}\mathcal{G}$ defines a vector bundle on $X$ (after saturation at the point $p$). For each $i$ we set
$\mathcal{E}_{i}:=(f_{*}\mathcal{E}_{i,Y})^{\mathrm{sat}}.$
Then each $\mathcal{E}_{i}$ is a vector bundle on $X$, and $f^{*}\mathcal{E}_{i}\simeq\mathcal{E}_{i,Y}$ away from $E$.
Step 2. Each inclusion $\mathcal{E}_{i-1,Y}\hookrightarrow\mathcal{E}_{i,Y}$ is saturated on $Y$, with torsion quotient $\mathcal{T}_{i,Y}$ supported on the effective Cartier divisor $D_{i,Y}$. Since $E\not\subset\mathrm{Supp}D_{i,Y}$, the support of $\mathcal{T}_{i,Y}$ avoids $E$. Applying $f_{*}$ yields an exact sequence on $U$:
$0\to f_{*}\mathcal{E}_{i-1,Y}|_{U}\to f_{*}\mathcal{E}_{i,Y}|_{U}\to f_{*}\mathcal{T}_{i,Y}|_{U}\to 0,$
where $f_{*}\mathcal{T}_{i,Y}$ remains a torsion sheaf supported on $f(D_{i,Y})\subset X\setminus\{p\}$. Passing to saturations gives a global inclusion of vector bundles
$\mathcal{E}_{i-1}\hookrightarrow\mathcal{E}_{i}$
on $X$.
Step 3. On $Y$ one has
$\det(\mathcal{E}_{i,Y})\simeq\det(\mathcal{E}_{i-1,Y})\otimes\mathcal{O}_{Y}(D_{i,Y}).$
Because $E\not\subset\mathrm{Supp}D_{i,Y}$, each $D_{i,Y}$ is the strict transform of a unique effective divisor $D_{i,X}\subset X$, i.e. $D_{i,Y}=f^{*}D_{i,X}$. Over $U$ the projection formula gives
$\det(f_{*}\mathcal{E}_{i,Y})|_{U}\simeq\det(f_{*}\mathcal{E}_{i-1,Y})|_{U}\otimes\mathcal{O}_{U}(D_{i,X}),$
and since taking reflexive hulls does not alter determinant line bundles on smooth surfaces, this equality extends globally to $X$:
$\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i,X}).$
Thus each step $\mathcal{E}_{i-1}\subset\mathcal{E}_{i}$ on $X$ is again an elementary transform with determinant increment $D_{i,X}$.
Step 4. The chain
$\mathfrak{F}_{X}:\quad 0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
is normalized by $\Lambda_{0}:=(f_{*}\Lambda_{0,Y})^{\mathrm{sat}}$, because $\det(\mathcal{E}_{0,Y})\simeq\Lambda_{0,Y}$ implies $\det(\mathcal{E}_{0})\simeq\Lambda_{0}$ by construction. All determinant increments descend faithfully, so the number of filtration steps is preserved:
$h_{\Lambda_{0}}(\mathcal{E})\geq h_{\Lambda_{0,Y}}(\mathcal{E}_{Y}).$
Conversely, pulling back $\mathfrak{F}_{X}$ to $Y$ via $f^{*}$ reproduces $\mathfrak{F}_{Y}$, hence $h_{\Lambda_{0,Y}}(\mathcal{E}_{Y})\geq h_{\Lambda_{0}}(\mathcal{E})$. Therefore
$h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,Y}}(\mathcal{E}_{Y}),$
as required. ∎
5.5. Hierarchical depth under MMP
Let $X$ be a smooth projective surface. Consider a sequence of birational maps
$X=X_{0}\dashrightarrow X_{1}\dashrightarrow\cdots\dashrightarrow X_{t}=X_{\min},$
obtained by running the minimal model program (MMP), where $X_{\min}$ is a minimal model of $X$. Each step of the MMP can be realized as a birational morphism
$f_{i}:X_{i}\longrightarrow X_{i+1},\qquad i=0,\dots,t-1,$
where $f_{i}$ is either
  • a contraction of a (1)(-1)-curve $E_{i}\subset X_{i}$ to a smooth point of $X_{i+1}$, or
  • the inverse operation (if one considers a blowup) would be $X_{i+1}\longrightarrow X_{i}$, introducing an exceptional (1)(-1)-curve on the blowup.
By composing these morphisms one obtains a birational map
$\phi:X=X_{0}\dashrightarrow X_{t}=X_{\min},$
which is an isomorphism outside the union of the exceptional curves of the contractions.
Let $\mathcal{E}$ be a vector bundle on $X$ and $\Lambda_{0}\in\mathrm{Pic}(X)$ a reference line bundle. For each contraction $f_{i}:X_{i}\to X_{i+1}$, the pullback $f_{i}^{*}\mathcal{E}_{i+1}$ to $X_{i}$ satisfies
$h_{f_{i}^{*}\Lambda_{i+1}}(f_{i}^{*}\mathcal{E}_{i+1})\geq h_{\Lambda_{i+1}}(\mathcal{E}_{i+1}),$
with equality unless the hierarchical filtration involves additional elementary transforms along the exceptional curve $E_{i}$. Iterating this through the entire MMP sequence gives
(2) $h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})\leq h_{\Lambda_{0}}(\mathcal{E}),$
where $\mathcal{E}_{\min}$ is the pushforward (or the vector bundle descended via saturation) on $X_{\min}$ and $\Lambda_{0,\min}$ is the corresponding reference line bundle. Hence, the maximal hierarchical depth can only decrease (or remain the same) when passing from $X$ to its minimal model.
Theorem 5.6.
Let $f:X\to X_{\min}$ be the birational contraction of a smooth projective surface $X$ onto its minimal model $X_{\min}$ obtained by contracting finitely many (1)(-1)–curves $E_{1},\dots,E_{m}\subset X$. Let $\mathcal{E}$ be a vector bundle on $X$ and let $\mathcal{E}_{\min}$ denote the (saturated) direct image of $\mathcal{E}$ on $X_{\min}$; fix reference line bundles
$\Lambda_{0}\in\mathrm{Pic}(X),\qquad\Lambda_{0,\min}\in\mathrm{Pic}(X_{\min})$
related by
$\Lambda_{0}\simeq f^{*}\Lambda_{0,\min}\otimes\mathcal{O}_{X}\Big(\sum_{j=1}^{m}\beta_{j}E_{j}\Big)$
for unique integers $\beta_{j}\in\mathbb{Z}$. Assume the determinants satisfy
$\det(\mathcal{E})\ \simeq\ f^{*}\det(\mathcal{E}_{\min})\otimes\mathcal{O}_{X}\Big(\sum_{j=1}^{m}\alpha_{j}E_{j}\Big)$
for unique integers $\alpha_{j}\in\mathbb{Z}$ (so that the class $c_{1}(\det\mathcal{E})-f^{*}c_{1}(\det\mathcal{E}_{\min})=\sum_{j}\alpha_{j}[E_{j}]$ in $\mathrm{NS}(X)$). Suppose moreover that the difference divisors appearing in the normalizations and determinants are effective on $X$, i.e. $\alpha_{j}-\beta_{j}\geq 0$ for every $j$.
Then hierarchical filtrations exist for the normalized bundles and the hierarchical depths are related by the exact formula
(3) $h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})+\sum_{j=1}^{m}(\alpha_{j}-\beta_{j}).$
In particular the increase in depth from the minimal model to $X$ is exactly the total multiplicity of the exceptional divisors appearing in the determinant of $\mathcal{E}$ after accounting for the chosen normalizations.
Proof.
By hypothesis
$\det(\mathcal{E})\otimes\Lambda_{0}^{-1}\ \simeq\ f^{*}\big(\det(\mathcal{E}_{\min})\otimes\Lambda_{0,\min}^{-1}\big)\otimes\mathcal{O}_{X}\Big(\sum_{j=1}^{m}(\alpha_{j}-\beta_{j})E_{j}\Big).$
Denote
$\Delta_{\min}:=\det(\mathcal{E}_{\min})\otimes\Lambda_{0,\min}^{-1}\in\mathrm{Pic}(X_{\min}),\qquad\Delta:=\det(\mathcal{E})\otimes\Lambda_{0}^{-1}\in\mathrm{Pic}(X).$
By the usual correspondence between hierarchical filtrations and decompositions of the determinant into effective Cartier increments (Definitions 2.1 and 2.2), the maximal length $h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})$ equals the maximal number of nonzero effective Cartier summands into which any effective representative of $\Delta_{\min}$ may be decomposed; likewise $h_{\Lambda_{0}}(\mathcal{E})$ equals the maximal number of nonzero effective Cartier summands in any effective representative of $\Delta$.
Because pullback preserves effectiveness and the exceptional curves $E_{j}$ are pairwise distinct irreducible Cartier divisors on $X$, any effective decomposition of $\Delta_{\min}$ pulls back to an effective decomposition of $f^{*}\Delta_{\min}$ with the same number of summands, and the factor $\mathcal{O}_{X}\big(\sum_{j}(\alpha_{j}-\beta_{j})E_{j}\big)$ contributes exactly $\sum_{j}(\alpha_{j}-\beta_{j})$ further irreducible effective summands (all supported on the exceptional locus) when one decomposes $\Delta$ into irreducible Cartier divisors. Conversely, every effective decomposition of $\Delta$ decomposes uniquely as the pullback of an effective decomposition of $\Delta_{\min}$ together with the exceptional part (since the $E_{j}$ generate the kernel of $f^{*}:\mathrm{Pic}(X_{\min})\to\mathrm{Pic}(X)$).
Therefore the maximal possible number of irreducible effective summands in $\Delta$ equals the maximal number for $\Delta_{\min}$ plus the sum of the coefficients of the exceptional part, i.e.
$h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})+\sum_{j=1}^{m}(\alpha_{j}-\beta_{j}).$
This proves (3). ∎
Example 5.7.
Let $X_{\min}=\mathbb{P}^{2}$ with hyperplane class $H$, and let $f:X\to X_{\min}$ be the blowup of $\mathbb{P}^{2}$ at two distinct points $p_{1},p_{2}$. Denote by $E_{1},E_{2}\subset X$ the corresponding exceptional divisors. Then
$\mathrm{Pic}(X)=\mathbb{Z}\langle f^{*}H\rangle\oplus\mathbb{Z}\langle E_{1}\rangle\oplus\mathbb{Z}\langle E_{2}\rangle.$
Fix the following data on $X_{\min}$:
$\mathcal{E}_{\min}=\mathcal{O}_{\mathbb{P}^{2}}(3)\oplus\mathcal{O}_{\mathbb{P}^{2}}(2)\qquad(\text{rank }2),\qquad\Lambda_{0,\min}=\mathcal{O}_{\mathbb{P}^{2}}.$
Then $\det(\mathcal{E}_{\min})\cong\mathcal{O}_{\mathbb{P}^{2}}(5)$. For a split bundle on $\mathbb{P}^{2}$, Proposition 2.4 (with $A=H$, $d=5$, $d_{0}=0$) gives the bound
$h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})\leq 5.$
In fact, since $\mathbb{P}^{2}$ has Picard rank one and the class $H$ admits pairwise disjoint representatives (distinct lines), one can achieve equality:
$h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})=5.$
On $X$ consider the vector bundle
$\mathcal{E}=f^{*}\mathcal{E}_{\min}\otimes\mathcal{O}_{X}(E_{1}+2E_{2}).$
Since $\mathrm{rank}(\mathcal{E})=2$, we compute
$\det(\mathcal{E})\simeq f^{*}\det(\mathcal{E}_{\min})\otimes\mathcal{O}_{X}\big(2(E_{1}+2E_{2})\big)=f^{*}\mathcal{O}_{\mathbb{P}^{2}}(5)\otimes\mathcal{O}_{X}(2E_{1}+4E_{2}).$
Take the normalization on $X$ to be
$\Lambda_{0}=\mathcal{O}_{X}(E_{2})\simeq f^{*}\mathcal{O}_{\mathbb{P}^{2}}\otimes\mathcal{O}_{X}(E_{2}).$
Thus in the notation of Theorem 5.6 we have:
$\det(\mathcal{E})\simeq f^{*}\det(\mathcal{E}_{\min})\otimes\mathcal{O}_{X}(\alpha_{1}E_{1}+\alpha_{2}E_{2}),\quad\alpha_{1}=2,\ \alpha_{2}=4,$
$\Lambda_{0}\simeq f^{*}\Lambda_{0,\min}\otimes\mathcal{O}_{X}(\beta_{1}E_{1}+\beta_{2}E_{2}),\quad\beta_{1}=0,\ \beta_{2}=1.$
Now Theorem 5.6 predicts:
$h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})+(\alpha_{1}-\beta_{1})+(\alpha_{2}-\beta_{2})=5+(2-0)+(4-1)=10.$
To verify directly, consider the class
$\Delta:=\det(\mathcal{E})\otimes\Lambda_{0}^{-1}\simeq f^{*}\mathcal{O}_{\mathbb{P}^{2}}(5)\otimes\mathcal{O}_{X}(2E_{1}+3E_{2}).$
A representative of $\Delta$ can be decomposed as:
$\Delta\sim(f^{*}H)+\cdots+(f^{*}H)+E_{1}+E_{1}+E_{2}+E_{2}+E_{2},$
where we have 5 copies of $f^{*}H$ (strict transforms of lines not passing through $p_{1}$ or $p_{2}$), 2 copies of $E_{1}$, and 3 copies of $E_{2}$. Each summand gives a possible determinant increment in a hierarchical filtration, so the maximal length is $5+2+3=10$. Hence
$h_{\Lambda_{0}}(\mathcal{E})=10,$
confirming the theorem.
6. Interaction with Stability and Moduli Theory
The hierarchical filtration formalism interacts naturally with classical and modern stability theories for vector bundles. Since hierarchical increments are governed by effective divisors, they admit a slope interpretation with respect to a fixed polarization, yielding a discrete refinement of the usual Harder–Narasimhan filtration. Moreover, the birational invariance results of Section 5 ensure that such stability behavior is preserved throughout the steps of the Minimal Model Program.
6.1. Slope behavior under hierarchical increments
Let $X$ be a smooth projective variety of dimension $n\geq 1$ with ample divisor $H$. For a torsion–free sheaf $\mathcal{E}$ on $X$, recall that its slope with respect to $H$ is
$\mu_{H}(\mathcal{E})=\frac{c_{1}(\mathcal{E})\cdot H^{n-1}}{\mathrm{rank}(\mathcal{E})}.$
Given a hierarchical filtration
$\mathfrak{F}:\quad 0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E},\qquad\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i}),$
we may associate to each increment the numerical slope difference
(4) $\Delta_{i}(\mathcal{E},H)=\mu_{H}(\mathcal{E}_{i})-\mu_{H}(\mathcal{E}_{i-1}).$
Theorem 6.1.
Let $X$ be a smooth projective variety of dimension $n\leq 2$ over an algebraically closed field, and let $H$ be an ample divisor on $X$. Let
$\mathfrak{F}:\quad 0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{h}=\mathcal{E}$
be a hierarchical filtration of a vector bundle $\mathcal{E}$ of rank $r$, with determinant increments
$\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i}),\qquad D_{i}\geq 0.$
Then
$\mu_{H}(\mathcal{E}_{0})\leq\mu_{H}(\mathcal{E}_{1})\leq\cdots\leq\mu_{H}(\mathcal{E}_{h}).$
and for each index $i$ the inequality $\mu_{H}(\mathcal{E}_{i})>\mu_{H}(\mathcal{E}_{i-1})$ holds precisely when $D_{i}$ is not numerically equivalent to zero. ($D_{i}\not\sim 0$).
Proof.
From the determinant relation $\det(\mathcal{E}_{i})=\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i})$ we have $c_{1}(\mathcal{E}_{i})=c_{1}(\mathcal{E}_{i-1})+[D_{i}]$. Since the rank $r$ is constant,
(5) $\mu_{H}(\mathcal{E}_{i})-\mu_{H}(\mathcal{E}_{i-1})=\frac{(c_{1}(\mathcal{E}_{i})-c_{1}(\mathcal{E}_{i-1}))\cdot H^{n-1}}{r}=\frac{c_{1}(D_{i})\cdot H^{n-1}}{r}.$
Each $D_{i}$ is effective, so the intersection number $c_{1}(D_{i})\cdot H^{n-1}$ is nonnegative, with equality if and only if $D_{i}\sim 0$. The stated monotonicity of the slopes follows immediately. ∎
6.2. Birational stability under the Minimal Model Program
The birational invariance results of Section 5 imply that hierarchical depth behaves predictably under the birational operations of the surface MMP. In particular, slope–compatibility is preserved under pullbacks, and semistability is unaffected outside exceptional loci.
Proposition 6.2.
Let $f:Y\to X$ be a birational morphism between smooth projective surfaces, and let $H$ be an ample divisor on $X$. Then for every vector bundle $\mathcal{E}$ on $X$ and its pullback $f^{*}\mathcal{E}$ on $Y$, the divisorial slope differences satisfy
$\Delta_{i}(f^{*}\mathcal{E},f^{*}H)=\Delta_{i}(\mathcal{E},H),$
and the divisorial slope filtration of $\mathcal{E}$ pulls back to that of $f^{*}\mathcal{E}$.
Proof.
Since $f^{*}H^{n-1}=H^{n-1}$ numerically and $f^{*}c_{1}(D_{i})=c_{1}(f^{*}D_{i})$, formula (5) gives $\Delta_{i}(f^{*}\mathcal{E},f^{*}H)=\Delta_{i}(\mathcal{E},H)$. The filtration structure is preserved by Proposition 5.1. ∎
Therefore, the stability behavior of hierarchical filtrations remains invariant throughout the birational steps of the MMP, except when new effective increments supported on exceptional curves appear. In that case, the newly introduced increments correspond to additional positive slopes localized on the exceptional locus, and their total contribution is governed by Theorem 5.6.
Corollary 6.3.
Let $f:Y\to X$ be a birational morphism between smooth projective surfaces obtained by contracting a finite collection of (1)(-1)–curves, and let $H$ be an ample divisor on $X$. Put $H_{Y}:=f^{*}H$.
Let $\mathcal{E}$ be a vector bundle on $X$. If $\mathcal{E}$ is divisorially slope-semistable (respectively, divisorially slope-stable) with respect to $H$, then the pullback bundle $f^{*}\mathcal{E}$ is divisorially slope–semistable (respectively, divisorially slope–stable) with respect to $H_{Y}$.
Proof.
We treat the semistable case; the stable case is identical with strict inequalities.
Let $r=\mathrm{rank}(\mathcal{E})$. Suppose $\mathcal{E}$ is divisorially slope–semistable on $X$ in meaning of [7], i.e. for every saturated nonzero proper subsheaf $F\subset\mathcal{E}$ one has
$\mu_{H}(F)\leq\mu_{H}(\mathcal{E}).$
We must show that every saturated nonzero proper subsheaf $G\subset f^{*}\mathcal{E}$ satisfies
$\mu_{H_{Y}}(G)\leq\mu_{H_{Y}}(f^{*}\mathcal{E})=\mu_{H}(\mathcal{E}).$
Consider any saturated subsheaf $G\subset f^{*}\mathcal{E}$. Push it forward and take saturation on $X$:
$F:=\big(f_{*}G\big)^{\mathrm{sat}}\subset\mathcal{E}.$
(Here saturation means reflexive hull / saturation in the sense of subsheaves.) By construction $F$ is a torsion–free subsheaf of $\mathcal{E}$ of rank $s:=\mathrm{rank}(F)$ with $0<s\leq r$.
There is an equality of determinant classes on $Y$ of the form
$\det(G)\simeq(f^{*}\det F)\otimes\mathcal{O}_{Y}\Big(\sum_{j}a_{j}E_{j}\Big),\qquad a_{j}\in\mathbb{Z}_{\geq 0},$
where the $E_{j}$ are the exceptional curves of $f$ and the integers $a_{j}$ record the possible exceptional components appearing in $\det(G)$. (Indeed every line bundle on $Y$ decomposes uniquely as the pullback of a line bundle on $X$ tensor a line bundle supported on exceptional curves.)
Now compute slopes with respect to the pulled back polarization $H_{Y}=f^{*}H$. Since intersection is functorial under pullback and every exceptional curve is contracted to a point, we have for every $j$
$E_{j}\cdot H_{Y}=E_{j}\cdot f^{*}H=0.$
Consequently the exceptional part does not contribute to the slope:
$\mu_{H_{Y}}(G)=\frac{c_{1}(\det G)\cdot H_{Y}}{\mathrm{rank}(G)}=\frac{\big(c_{1}(f^{*}\det F)+\sum_{j}a_{j}[E_{j}]\big)\cdot f^{*}H}{\mathrm{rank}(G)}=\frac{c_{1}(\det F)\cdot H}{\mathrm{rank}(G)}=\mu_{H}(F).$
But the right-hand side equals $\mu_{H}(F)$ (by definition). Hence
$\mu_{H_{Y}}(G)=\mu_{H}(F).$
Since $\mathcal{E}$ is divisorially slope–semistable, $\mu_{H}(F)\leq\mu_{H}(\mathcal{E})$. Moreover $\mu_{H}(\mathcal{E})=\mu_{H_{Y}}(f^{*}\mathcal{E})$ because $c_{1}(f^{*}\mathcal{E})=f^{*}c_{1}(\mathcal{E})$ and intersection with $f^{*}H$ is compatible with pullback. Therefore
$\mu_{H_{Y}}(G)=\mu_{H}(F)\leq\mu_{H}(\mathcal{E})=\mu_{H_{Y}}(f^{*}\mathcal{E}),$
as required. This proves that $f^{*}\mathcal{E}$ is divisorially slope–semistable with respect to $H_{Y}$.
The same argument with strict inequalities shows that divisorial slope–stability is preserved: if $\mathcal{E}$ is stable and $G\subset f^{*}\mathcal{E}$ is a proper saturated subsheaf, then the strict inequality $\mu_{H}(F)<\mu_{H}(\mathcal{E})$ holds, and hence $\mu_{H_{Y}}(G)<\mu_{H_{Y}}(f^{*}\mathcal{E})$. ∎
6.3. Comparison with Harder–Narasimhan filtrations
Let $\mathcal{E}$ be a torsion–free sheaf on $(X,H)$ with Harder–Narasimhan filtration
$\mathcal{E}^{(0)}\subset\mathcal{E}^{(1)}\subset\cdots\subset\mathcal{E}^{(s)}=\mathcal{E},$
whose successive quotients $\mathrm{Gr}_{i}^{\mathrm{HN}}(\mathcal{E})=\mathcal{E}^{(i)}/\mathcal{E}^{(i-1)}$ are semistable with decreasing slopes $\mu_{1}>\mu_{2}>\cdots>\mu_{s}$. Then the determinant relation gives
$c_{1}(\det\mathcal{E})=\sum_{i=1}^{s}c_{1}\big(\mathrm{Gr}_{i}^{\mathrm{HN}}(\mathcal{E})\big),$
so that each Harder–Narasimhan quotient contributes an effective class whenever $\mathcal{E}$ is globally generated or arises as a positive extension.
Split bundle on a curve. Let $C$ be a smooth projective curve over a field and set
$\mathcal{E}=\bigoplus_{i=1}^{r}\mathcal{O}_{C}(d_{i}),\qquad d_{i}\in\mathbb{Z}.$
Write $D:=\sum_{i=1}^{r}d_{i}=\deg(\det\mathcal{E})$. We consider two invariants:
With notation as above we obtain the hierarchical depth and the length of Harder-Narasimhan filtration of $\mathcal{E}$.
(1) Hierarchical depth.
It was shown in Proposition 4.1 that
$h_{C}(\mathcal{E})=\sum_{i=1}^{r}\max\{d_{i},0\}.$ In particular, if all $d_{i}\geq 0$ then $h_{\mathcal{O}_{C}}(\mathcal{E})=\sum_{i}d_{i}=D$. If some $d_{i}<0$ the corresponding summand contributes 0 to the depth.
(2) Harder–Narasimhan length.
The Harder-Narasimhan length $s$ equals the number of distinct values among the degrees $\{d_{1},\dots,d_{r}\}$. Equivalently, if we arrange the degrees in nonincreasing order and group equal values,
$\mathcal{E}=\bigoplus_{j=1}^{s}\Big(\bigoplus_{k=1}^{r_{j}}\mathcal{O}_{C}(d^{(j)})\Big),$
where $d^{(1)}>\cdots>d^{(s)}$ are the distinct degrees and $r_{j}$ the multiplicities, then the HN filtration has length $s$ with graded pieces $\mathrm{Gr}_{j}^{\mathrm{HN}}\cong\mathcal{O}_{C}(d^{(j)})^{\oplus r_{j}}$. In the following we imply to this fact.
Let $C$ be a smooth projective curve over an algebraically closed field, and let
$\mathcal{E}=\bigoplus_{i=1}^{r}\mathcal{O}_{C}(d_{i})$
be a direct sum of line bundles of degrees $d_{1},\dots,d_{r}\in\mathbb{Z}$. Then the Harder–Narasimhan filtration of $\mathcal{E}$ (with respect to any polarization) is obtained by ordering the summands in nonincreasing degree:
$\mathcal{E}^{(0)}=0\subset\mathcal{E}^{(1)}=\bigoplus_{d_{i}=d_{(1)}}\mathcal{O}_{C}(d_{i})\subset\mathcal{E}^{(2)}=\bigoplus_{d_{i}\geq d_{(2)}}\mathcal{O}_{C}(d_{i})\subset\cdots\subset\mathcal{E}^{(s)}=\mathcal{E},$
where $d_{(1)}>d_{(2)}>\cdots>d_{(s)}$ are the distinct degrees among the $d_{i}$. In particular,
[…]
The existence and uniqueness of the Harder–Narasimhan filtration for torsion-free sheaves on smooth projective varieties was established by Harder and Narasimhan [7].
On a smooth projective curve $C$, every line bundle $\mathcal{O}_{C}(d)$ is stable with respect to any polarization, and its slope is simply
$\mu(\mathcal{O}_{C}(d))=\frac{\deg\mathcal{O}_{C}(d)}{\mathrm{rank}\mathcal{O}_{C}(d)}=d.$
Hence each direct summand $\mathcal{O}_{C}(d_{i})$ is a semistable bundle of slope $d_{i}$. Let $d_{(1)}>\cdots>d_{(s)}$ be the distinct degrees among the $d_{i}$, and define
$\mathcal{E}^{(k)}:=\bigoplus_{d_{i}\geq d_{(k)}}\mathcal{O}_{C}(d_{i}).$
Then the successive quotients
$\mathrm{Gr}_{k}^{\mathrm{HN}}(\mathcal{E}):=\mathcal{E}^{(k)}/\mathcal{E}^{(k-1)}=\bigoplus_{d_{i}=d_{(k)}}\mathcal{O}_{C}(d_{i})$
are semistable bundles of slope $\mu_{k}=d_{(k)}$, and the sequence $\mu_{1}>\mu_{2}>\cdots>\mu_{s}$ is strictly decreasing. By the uniqueness of the Harder-Narasimhan filtration this is precisely the HN filtration of $\mathcal{E}$.
Example 6.4.
  1. (i)
    Take $r=3$ and $\mathcal{E}=\mathcal{O}_{C}(3)\oplus\mathcal{O}_{C}(1)\oplus\mathcal{O}_{C}(0)$. Then $h_{\mathcal{O}_{C}}(\mathcal{E})=3+1+0=4$. The HN degrees are $3>1>0$, so $s=3$.
  2. (ii)
    Take $r=3$ and $\mathcal{E}=\mathcal{O}_{C}(1)\oplus\mathcal{O}_{C}(1)\oplus\mathcal{O}_{C}(0)$. Then $h_{\mathcal{O}_{C}}(\mathcal{E})=1+1+0=2$. The HN degrees are $1>0$, so $s=2$.
  3. (iii)
    Take $r=3$ and $\mathcal{E}=\mathcal{O}_{C}(1)\oplus\mathcal{O}_{C}(0)\oplus\mathcal{O}_{C}(0)$. Then $h_{\mathcal{O}_{C}}(\mathcal{E})=1$ and the HN filtration has length $s=2$.
It follows that there is no relation between hierarchical depth of a vector bundle and the length of its HN filtration.
7. Applications to algebraic–geometric codes
Let $X$ be a smooth projective variety over a finite field $\mathbb{F}_{q}$, $\mathcal{E}$ a vector bundle on $X$, and $D\subset X(\mathbb{F}_{q})$ a set of $N$ rational points. The associated algebraic–geometric code is $C_{X}(\mathcal{E},D)=\operatorname{ev}_{D}(H^{0}(X,\mathcal{E}))\subset\mathbb{F}_{q}^{\,rN}$ where $r=\mathrm{rank}(\mathcal{E})$.
Hierarchical filtrations of $\mathcal{E}$ expose degeneracies: if an elementary transform occurs along a divisor containing evaluation points, the corresponding code coordinates vanish. Consequently, hierarchical depth $h_{\Lambda_{0}}(\mathcal{E})$ bounds the number of such degenerate layers.
For surfaces, the minimal model program contracts (1)(-1)-curves—smooth point blowups in reverse—and Theorem 5.6 quantifies how hierarchical depth changes under these birational maps. Contracting exceptional divisors eliminates the associated degenerate coordinates, improving the code’s normalized minimum distance while preserving dimension. The next subsections provide explicit constructions demonstrating this optimization.
7.1. AG codes from vector bundles
Let $X$ be a smooth projective variety over $\mathbb{F}_{q}$ and let $\mathcal{E}$ be a vector bundle of rank $r$ on $X$. Let
$D=\{P_{1},\dots,P_{N}\}\subset X(\mathbb{F}_{q})$
be a finite ordered set of $N$ distinct $\mathbb{F}_{q}$–rational points such that evaluation of sections of $\mathcal{E}$ is well defined at every $P_{i}$. Fix, for each $P_{i}$, an $\mathbb{F}_{q}$–linear identification (trivialization) of the fiber $\mathcal{E}|_{P_{i}}\cong\mathbb{F}_{q}^{\,r}$.
The algebraic–geometric code associated to $(\mathcal{E},D)$ is defined by
$C_{X}(\mathcal{E},D):=\operatorname{ev}_{D}\big(H^{0}(X,\mathcal{E})\big)=\{(s(P_{1}),\dots,s(P_{N})):s\in H^{0}(X,\mathcal{E})\}\subseteq\mathbb{F}_{q}^{\,rN}.$
For details of AG codes from vector bundles see [11, 16].
7.2. Minimal models and coding–theoretic optimization
Let $X$ be a smooth projective surface defined over $\mathbb{F}_{q}$, and let $f:Y\to X$ denote a blow–up of $X$ at a smooth $\mathbb{F}_{q}$–rational point, with exceptional divisor $E\subset Y$. As discussed in Proposition 5.4, such a birational modification increases the hierarchical depth of vector bundles by an integer $\delta\geq 0$, reflecting the presence of additional effective increments supported on $E$.
From the coding–theoretic viewpoint, these exceptional divisors have unfavourable effects on the parameters of algebraic–geometric codes constructed from vector bundles. In order to optimize the geometric data used for encoding, it is therefore natural to pass to the minimal model of $X$, obtained by contracting all (1)(-1)–curves.
Proposition 7.1.
Let $f:Y\to X$ be the contraction of a (1)(-1)-curve $E\subset Y$ onto a smooth $\mathbb{F}_{q}$-rational point $P_{0}\in X$. Let $\mathcal{E}_{Y}$ be a vector bundle on $Y$ such that:
  1. (i)
    $\mathcal{E}_{Y}$ is $f$-trivial in the sense that $R^{1}f_{*}\mathcal{E}_{Y}=0$;
  2. (ii)
    $\det(\mathcal{E}_{Y})$ has no component supported on $E$ (equivalently, in the decomposition $\det(\mathcal{E}_{Y})\simeq f^{*}L\otimes\mathcal{O}_{Y}(mE)$, we have $m=0$).
Set $\mathcal{E}:=(f_{*}\mathcal{E}_{Y})^{\mathrm{sat}}$. Then there is a canonical identification
$H^{0}(Y,\mathcal{E}_{Y})\cong H^{0}(X,\mathcal{E}),$
and for any evaluation set $D\subset X(\mathbb{F}_{q})\setminus\{P_{0}\}$ one has
$C_{X}(\mathcal{E},D)=C_{Y}(\mathcal{E}_{Y},f^{-1}(D)).$
Proof.
Condition (i) gives $R^{i}f_{*}\mathcal{E}_{Y}=0$ for $i>0$ (since $R^{i}f_{*}\mathcal{E}_{Y}$ are supported at $P_{0}$ and vanish for $i>1$ by dimension reasons). The projection formula and the fact that $f_{*}\mathcal{O}_{Y}\cong\mathcal{O}_{X}$ yield
$H^{0}(Y,\mathcal{E}_{Y})\cong H^{0}(X,f_{*}\mathcal{E}_{Y})\cong H^{0}(X,\mathcal{E}),$
where the second isomorphism uses that saturation does not change global sections.
Condition (ii) ensures that for any $P\in D$, the natural map $(f_{*}\mathcal{E}_{Y})_{P}\to\mathcal{E}_{Y}|_{f^{-1}(P)}$ is an isomorphism (the torsion of $f_{*}\mathcal{E}_{Y}$ is concentrated at $P_{0}$). Hence evaluation at $P$ and at $f^{-1}(P)$ agree, giving the equality of codes. ∎
Remark 7.2.
Condition (i) holds, for instance, when $\mathcal{E}_{Y}|_{E}$ is a direct sum of line bundles of degree $\geq-1$, or when $\mathcal{E}_{Y}$ is pulled back from $X$ twisted by $\mathcal{O}_{Y}(-mE)$ with $0\leq m\leq 1$. Condition (ii) is exactly the requirement that no step in a hierarchical filtration of $\mathcal{E}_{Y}$ involves an elementary transform along $E$.
Remark 7.3.
While it is folklore that birational varieties can yield equivalent AG codes when evaluation points avoid exceptional loci, Proposition 7.1 provides precise cohomological and determinantal criteria in terms of hierarchical filtrations. In particular, condition (ii) is equivalent to the vanishing of the exceptional contribution $\delta$ in Theorem 5.6.
Corollary 7.4.
Let $X$ be a smooth projective surface over $\mathbb{F}_{q}$, $f:X\to X_{\min}$ the contraction of all (1)(-1)-curves to its minimal model, and $\mathcal{E}$ a vector bundle on $X$ satisfying the hypotheses of Proposition 7.1 for each contraction step. Let $D\subset X(\mathbb{F}_{q})$ be an evaluation set, and decompose $D=D_{\mathrm{ex}}\cup D_{\mathrm{reg}}$ where $D_{\mathrm{ex}}$ lies on exceptional curves and $D_{\mathrm{reg}}$ does not. Set $D_{\min}:=f(D_{\mathrm{reg}})\subset X_{\min}(\mathbb{F}_{q})$.
Then the AG codes are related by
$C_{X}(\mathcal{E},D)\cong C_{X_{\min}}(f_{*}\mathcal{E},\,D_{\min})\times\{0\}^{r\cdot|D_{\mathrm{ex}}|},$
where $\cong$ denotes isomorphism of $\mathbb{F}_{q}$-vector spaces. Consequently:
  • $\dim C_{X}(\mathcal{E},D)=\dim C_{X_{\min}}(f_{*}\mathcal{E},D_{\min})$;
  • $d_{\min}(C_{X}(\mathcal{E},D))=d_{\min}(C_{X_{\min}}(f_{*}\mathcal{E},D_{\min}))$;
  • $|D|=|D_{\min}|+|D_{\mathrm{ex}}|$, so the length decreases by $r\cdot|D_{\mathrm{ex}}|$ coordinates.
In particular, the normalized minimum distance $\delta=d_{\min}/(r\cdot|D|)$ strictly increases when $D_{\mathrm{ex}}\neq\varnothing$.
Corollary 7.5.
Under the hypotheses of Proposition 7.1, let $D_{\mathrm{ex}}\subset D$ be evaluation points on exceptional curves. Then
$\frac{d_{\min}(C_{X_{\min}}(f_{*}\mathcal{E},f(D\setminus D_{\mathrm{ex}})))}{r\cdot|D\setminus D_{\mathrm{ex}}|}>\frac{d_{\min}(C_{X}(\mathcal{E},D))}{r\cdot|D|},$
with strict inequality when $D_{\mathrm{ex}}\neq\varnothing$.
Proof.
Immediate from Proposition 7.1 since removing zero coordinates preserves $d_{\min}$ while reducing length. ∎
Example 7.6.
Let $X=\mathbb{P}^{2}_{\mathbb{F}_{5}}$ be the projective plane over $\mathbb{F}_{5}$, and let $f:Y\to X$ be the blow-up of $X$ at the $\mathbb{F}_{5}$–rational point $p=[1:0:0]$, with exceptional divisor $E\subset Y$.
Consider the line bundle on $Y$
$\mathcal{E}_{Y}:=f^{*}\mathcal{O}_{X}(3)\otimes\mathcal{O}_{Y}(-E)=\mathcal{O}_{Y}(3H-E),$
where $H=f^{*}(\text{line})$ is the pullback of a line in $\mathbb{P}^{2}$.
Global sections.
From the exact sequence
$0\to\mathcal{O}_{Y}(3H-E)\to\mathcal{O}_{Y}(3H)\to\mathcal{O}_{E}(3H|_{E})\to 0,$
where $E\cong\mathbb{P}^{1}$ and $\mathcal{O}_{E}(3H|_{E})\cong\mathcal{O}_{\mathbb{P}^{1}}(3)$, we have
$0\to H^{0}(Y,\mathcal{O}_{Y}(3H-E))\to H^{0}(Y,\mathcal{O}_{Y}(3H))\to H^{0}(E,\mathcal{O}_{\mathbb{P}^{1}}(3)).$
Since $H^{0}(Y,\mathcal{O}_{Y}(3H))\cong H^{0}(X,\mathcal{O}_{X}(3))$ has dimension $10$ and $H^{0}(E,\mathcal{O}_{\mathbb{P}^{1}}(3))$ has dimension $4$, we obtain
$h^{0}(Y,\mathcal{E}_{Y})\geq 10-4=6.$
In fact, equality holds as the map $H^{0}(Y,\mathcal{O}_{Y}(3H))\to H^{0}(E,\mathcal{O}_{\mathbb{P}^{1}}(3))$ is surjective (cubics can take arbitrary values at $p$ to order $3$). Thus
$h^{0}(Y,\mathcal{E}_{Y})=6.$
Let $D_{Y}$ be a set of $N_{Y}=12$ $\mathbb{F}_{5}$–rational points on $Y$, decomposed as
$D_{Y}=D^{\prime}\cup D_{E},\qquad\#D^{\prime}=10,\ \#D_{E}=2,\quad D_{E}\subset E(\mathbb{F}_{5}),$
so that two evaluation points lie on the exceptional curve $E$.
Code on the blown-up surface.
The AG code on $Y$ is
$C_{Y}(\mathcal{E}_{Y},D_{Y})=\operatorname{ev}_{D_{Y}}(H^{0}(Y,\mathcal{E}_{Y}))\subset\mathbb{F}_{5}^{\,rN_{Y}},\quad r=1.$
By construction, all sections vanish on $E$, so the coordinates corresponding to $D_{E}$ are identically zero for every codeword. Hence
$C_{Y}(\mathcal{E}_{Y},D_{Y})\cong\{0\}^{\,2}\times C_{Y}(\mathcal{E}_{Y},D^{\prime}),$
with $\dim C_{Y}(\mathcal{E}_{Y},D_{Y})=\dim C_{Y}(\mathcal{E}_{Y},D^{\prime})=6$.
For a concrete minimum distance, note that a cubic vanishing at $p$ can vanish at up to $7$ additional $\mathbb{F}_{5}$-points on a line (since a line has 66 $\mathbb{F}_{5}$-points, and the cubic restricted to the line is a degree $3$ polynomial). A careful choice of evaluation points yields, for instance,
$d_{\min}(C_{Y}(\mathcal{E}_{Y},D_{Y}))=d_{\min}(C_{Y}(\mathcal{E}_{Y},D^{\prime}))=4.$
The normalized (relative) minimum distance is
$\delta_{Y}=\frac{d_{\min}}{N_{Y}}=\frac{4}{12}=\frac{1}{3}\approx 0.333.$
Code on the minimal model.
On the contracted surface $X=\mathbb{P}^{2}$, the direct image bundle
$\mathcal{E}:=f_{*}\mathcal{E}_{Y}=\mathcal{O}_{X}(3)$
has $\dim H^{0}(X,\mathcal{E})=10$, but only the $6$-dimensional subspace corresponding to cubics vanishing at $p$ yields nonzero codes when evaluated away from $p$. Let
$D:=f(D^{\prime})\subset X(\mathbb{F}_{5}),\quad\#D=10.$
The corresponding code using the same $6$-dimensional subspace is
$C_{X}(\mathcal{E},D)=\operatorname{ev}_{D}(H^{0}(Y,\mathcal{E}_{Y}))\subset\mathbb{F}_{5}^{10},$
where we identify $H^{0}(Y,\mathcal{E}_{Y})$ with the subspace of $H^{0}(X,\mathcal{O}_{X}(3))$ vanishing at $p$.
The minimum distance remains $d_{\min}(C_{X}(\mathcal{E},D))=4$, and the dimension is still $6$. The normalized minimum distance is
$\delta_{X}=\frac{d_{\min}}{10}=\frac{4}{10}=0.40>0.333=\delta_{Y}.$
Example 7.7.
Let $X=\mathbb{P}^{2}_{\mathbb{F}_{7}}$ and let $f:Y\to X$ be the blowup of $X$ at the $\mathbb{F}_{7}$-rational point $p=[1:0:0]$. Let $E\subset Y$ denote the exceptional (1)(-1)-curve, with $E\cong\mathbb{P}^{1}$ having 88 $\mathbb{F}_{7}$-rational points. Denote by $\mathcal{I}_{p}\subset\mathcal{O}_{X}$ the ideal sheaf of the point $p$.
Consider the rank $r=3$ vector bundle on $Y$:
$\mathcal{E}_{Y}:=\mathcal{O}_{Y}(3H-E)\oplus\mathcal{O}_{Y}(2H-E)\oplus\mathcal{O}_{Y}(H-E),$
where $H=f^{*}(\text{line})$ is the pullback of a line in $\mathbb{P}^{2}$.
Step 1: Global sections.
For each summand, we compute dimensions using that $H^{0}(Y,\mathcal{O}_{Y}(dH-E))$ is isomorphic to the space of degree $d$ homogeneous polynomials vanishing at $p$:
$h^{0}(Y,\mathcal{O}_{Y}(H-E))$ $=h^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(1))=\dim\{\text{linear forms through }p\}=2,$
$h^{0}(Y,\mathcal{O}_{Y}(2H-E))$ $=h^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(2))=\dim\{\text{quadrics through }p\}=5,$
$h^{0}(Y,\mathcal{O}_{Y}(3H-E))$ $=h^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(3))=\dim\{\text{cubics through }p\}=9.$
The dimensions follow from: $h^{0}(\mathbb{P}^{2},\mathcal{O}(d))=\binom{d+2}{2}$ and imposing one independent linear condition reduces dimension by 1. Hence
$h^{0}(Y,\mathcal{E}_{Y})=9+5+2=16.$
Step 2: Evaluation points.
Choose $N_{Y}=18$ $\mathbb{F}_{7}$-rational points on $Y$ for evaluation, decomposed as
$D_{Y}=D^{\prime}\cup D_{E},\qquad\#D_{E}=5,\ \#D^{\prime}=13,\quad D_{E}\subset E(\mathbb{F}_{7}).$
We assume $D^{\prime}$ consists of points not on $E$ and in general position.
Step 3: AG code on $Y$.
The evaluation map
$\operatorname{ev}_{D_{Y}}:H^{0}(Y,\mathcal{E}_{Y})\longrightarrow(\mathbb{F}_{7}^{3})^{18}$
defines the AG code $C_{Y}(\mathcal{E}_{Y},D_{Y})$.
Since every section of $\mathcal{E}_{Y}$ vanishes along $E$ (due to the E-E twist), the coordinates corresponding to $D_{E}$ are identically zero:
$C_{Y}(\mathcal{E}_{Y},D_{Y})\cong\{0\}^{3\cdot 5}\times C_{Y}(\mathcal{E}_{Y},D^{\prime}),\quad\dim C_{Y}(\mathcal{E}_{Y},D_{Y})=16.$
To estimate the minimum distance, note that a nonzero section $s=(s_{3},s_{2},s_{1})\in H^{0}(Y,\mathcal{E}_{Y})$ has components $s_{d}\in H^{0}(Y,\mathcal{O}_{Y}(dH-E))$. Each $s_{d}$ vanishes at $p$ and can vanish at additional points. For generic $D^{\prime}$, a section can vanish at at most $(\deg(s_{3})+\deg(s_{2})+\deg(s_{1}))-1=(3+2+1)-1=5$ points of $D^{\prime}$ (by a parameter count). Thus we may have
$d_{\min}(C_{Y}(\mathcal{E}_{Y},D_{Y}))=d_{\min}(C_{Y}(\mathcal{E}_{Y},D^{\prime}))=13-5=8.$
The normalized minimum distance is
$\delta_{Y}=\frac{d_{\min}}{3\cdot 18}=\frac{8}{54}\approx 0.148.$
Step 4: AG code on the minimal model.
On $X=\mathbb{P}^{2}$, consider the subspace
$V:=H^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(3))\oplus H^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(2))\oplus H^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(1))$
of $H^{0}(X,\mathcal{O}_{X}(3)\oplus\mathcal{O}_{X}(2)\oplus\mathcal{O}_{X}(1))$, which corresponds to $H^{0}(Y,\mathcal{E}_{Y})$ under the natural isomorphism $H^{0}(Y,\mathcal{E}_{Y})\cong V$. Let
$D=f(D^{\prime})\subset X(\mathbb{F}_{7}),\quad\#D=13.$
The code using this subspace is
$C_{X}(V,D)=\operatorname{ev}_{D}(V)\subset(\mathbb{F}_{7}^{3})^{13}.$
The dimension remains $16$, and the minimum distance analysis is similar: a nonzero section $(s_{3},s_{2},s_{1})\in V$ vanishes at $p$ and can vanish at up to 5 additional points of $D$ (since $s_{3}$ is a cubic, $s_{2}$ a quadric, $s_{1}$ a linear form, all vanishing at $p$). Thus
$d_{\min}(C_{X}(V,D))=13-5=8.$
The normalized minimum distance improves to
$\delta_{X}=\frac{8}{3\cdot 13}=\frac{8}{39}\approx 0.205>0.148\approx\delta_{Y}.$
Remark 7.8.
For a contraction $f:Y\to X$ of (1)(-1)-curves $E_{1},\dots,E_{m}$ and a bundle $\mathcal{E}_{Y}$ whose sections vanish along $\bigcup_{i}E_{i}$, let $N_{E}$ be the number of evaluation points on the exceptional locus. Then
$\frac{\delta_{X}}{\delta_{Y}}=\frac{N_{Y}}{N_{Y}-N_{E}}>1,$
provided $N_{E}<N_{Y}$. This shows the normalized distance improvement is purely combinatorial, depending only on the proportion of “wasted” evaluation points.